A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1965–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.14

Title word cross-reference

(0p, 0d, 1s) [1221]. (d + s)^{n}p [1264]. (l, s) [31]. (λ, μ) ⊕ (3, 0) [200]. (λ, μ) ⊗ (1, 1) [90]. (N, p, T) [543]. (N, V, T) [543]. (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 [819, 1167]. 12 [823]. 2 × 2 [1297]. 2V [891]. 3 [344, 823, 1139, 1013, 1014, 1051, 816]. 3 + 2 [569]. 35 ⊗ 70 [1286]. 4 [481, 362]. 6 [823]. 9 [908, 823]. ^{−4} [264]. ^{4} [881]. _{3} [634, 600, 257]. _{5} [1422]. _{n} [636]. A_{2}(SU(3)) [489]. α [749]. B^{∗} [730]. B_{2}(SO(5)) [489]. D [1212, 1285]. d^{n−1}sp [945]. – D [817]. δ [788]. E(2) [1148]. E(3, 1) [1370]. ℓ^{n−1}ℓ′ℓ′′ [1095]. F [1287, 455]. f(n + 1) + f(n − 1) = [E − ϕ(n)]f(n) [1182]. F_{2} [1198]. γ [1085]. Γ_{[jk]}^{i} [1123]. IO(n) [1129]. IU(n) [1129]. j [344, 908, 823]. K_{n} = A_{n}: S_{n} [938]. L [320]. l < N [31]. λ [277]. Λ = 0 [399]. M [1319, 254]. N [470, 1415, 1282, 1376, 202, 537, 617, 658, 1263, 1223, 804, 231, 143, 254, 870]. N/D [468, 835, 758, 834, 1061, 422]. N × N [1319]. O(1, 1) [1130, 825]. O(2, 1) [633, 1241, 1130, 825, 869, 1399, 1400]. O(3, 1) [923]. O(5) [941]. O(n) ⊃ O(n − 1) ⊃ [710]. O(n, 1) [1129]. P [1061, 1223]. P (x, y, t = 0) = Φ(x, y) [1167]. q [391]. r [469]. S [1193, 1202, 1203, 990, 874, 320, 643, 1242, 1018, 101, 102, 103, 395, 628, 227, 204, 1291, 467, 1144, 1139, 538, 92, 313, 289, 925, 410, 817]. SL(2, C) [1148]. SL(2, C)/SU(2) [1197]. SL(2C) [711]. SL(3, C) [1246]. SL(3, R) [427]. SL(n, C) [981]. SO(4, 1) [1392, 1370]. SO(5) [1345]. SO(ν) [1254]. SO_{0}(p, q) [789]. SO_{0}(p, q − 1) [789]. SO_{0}(p − 1, q) [789]. SP(4) [1345]. Sp(n) [1015]. Sp(p, q) [1354]. SU(1, 1) [711]. SU(2) [726, 1044]. SU(2, 2) [790, 1070]. SU(3) [86, 4, 262, 639, 726, 98, 703, 1385, 659, 200, 557, 558, 408]. SU(6) [1286]. SU(N) [1410, 570, 301, 1373, 823, 556, 403]. SU(n) ⊃ SU(n − 1) ⊃ ⊃ SU(2) [1406]. SU(ν) [1254]. SU(ν) ⊗ SU(ν) [1254]. SU_{3} [239, 67, 210, 1199, 1221, 90, 93]. SU_{4} [666]. SU_{m+n} [177]. SU_{mn} [176]. SU_{m} ⊗ SU_{n} [176, 177]. SU_{n} [906, 564]. T [346, 765, 658]. θ [229]. U(2, 2) [761, 1398]. U(6) ⊗ U(6) [1108]. U(6, 6) [746]. U(N) [1410, 1338]. U(n, 1) [1129]. U_{4} [998]. U_{4} ⊃ U_{2} + U_{2} [998]. U_{n} [1163, 87]. U_{n−1} [87]. V [1397, 229]. V Θ [311, 147]. X [974]. x = 1 [1198]. Y [974]. y = 1 [1198].

-Body [1415, 1282]. -Bound [311]. -density [1223]. -Dimensional [231, 254]. -Equivalent [391]. -Fermion [1263]. -Function [788]. -Matrices [628]. -Matrix [643, 1242, 101, 102, 103, 395, 227, 204, 1291, 467, 1139, 538, 289, 410, 1193, 1202, 1203]. -Model [455]. -Operator [92, 313]. -Particle [469, 1139, 870, 1397, 202, 537, 617, 658]. -Representability [804, 1376, 1223]. -Space [1013, 1014, 1051]. -Spin-Wave [143]. -Spinors [481]. -Symbols [344]. -Vector [362]. -Wave [990, 874, 1061].

10 [1431, 1432].

2V [985].

3-Sphere [1000].

9 [1437].

= [587].

Abnormal [1180]. Absence [689]. Absorptive [583, 476]. Accidental [1403]. Accommodating [969]. According [67]. Acoustic [1356, 709, 517, 1186]. Action [578, 1265, 573, 1300, 1266, 165]. Action-at-a-Distance [573, 1300]. Addendum [1437]. Addition [496]. Adiabatic [365, 367, 43]. Adiabatically [972]. Adjacent [684]. Adjoint [868]. Admittance [1081]. Admitting [864]. Advanced [1079]. Affine [886, 1427]. Albedo [140]. Algebra [964, 743, 802, 1256, 205, 644, 718, 1371, 1372, 1221, 324, 746, 22, 172, 586, 1124, 882, 624, 328]. Algebraic [70, 131, 149, 1256, 532, 489, 1228, 90, 200, 357, 1264, 479, 181, 1230, 780]. Algebraically [742]. Algebras [504, 1129, 802, 1083, 394, 730, 708, 867, 868, 1174, 691, 791, 521, 1030].

Algorithm [1282, 1085]. Almost [1074]. Alpha [919]. Alpha-Function [919]. Alternate [853, 404, 1154]. Alternative [214, 1378]. among [908, 1164]. Amplitude [1381, 1361, 273, 583, 987, 447, 352, 1122, 662, 1309, 212, 135, 695, 1437]. Amplitudes [1240, 1296, 1179, 50, 855, 275, 1292, 1094, 1249, 1218]. Analysis [1279, 1410, 1343, 918, 599, 553, 1197, 1423, 2, 175, 836]. Analytic [211, 796, 1432, 39, 680, 579, 802, 473, 145, 769, 929, 1430, 1198, 1130, 1248, 662, 356, 1218, 522, 1045, 480, 897, 31]. Analytical [1244, 1254]. Analyticity [987, 855, 1395, 101, 102, 103, 244, 838, 352, 538, 1136, 1065]. Angle [1112, 82, 1072, 1299, 51, 428, 604, 1141]. Angles [1194]. Angular [973, 1001, 273, 776, 105, 178, 1, 5, 83, 84, 263, 112, 158, 745, 792, 217, 1189]. Angular-Momentum [1001, 1189]. Anharmonic [922, 621]. Anisotropic [813, 1279, 954, 881, 444, 323, 830, 251, 1113, 372]. Annihilation [1417]. Antenna [1192]. Antiferromagnet [1425, 40]. Antilinear [718]. Antisymmetric [1040]. Antisymmetrized [186]. Antiunitary [503]. Any [158, 1078]. Appell [1361, 1198]. Applicability [1177]. Application [1361, 758, 506, 1241, 788, 555, 544, 59, 282, 655, 808, 1309, 1412, 1021, 2, 175, 459, 1229]. Applications [292, 546, 767, 55, 242, 58, 718, 164, 1140, 895, 1071, 234]. Approach [75, 957, 677, 19, 1205, 581, 637, 419, 1243, 725, 970, 422, 574, 304, 1107, 119, 1097, 1188, 405, 406, 1021, 480, 719, 1186, 1149, 1105, 1223]. Approaches [173]. Approximant [610]. Approximants [269, 758, 866, 1238, 1367]. Approximate [1321, 327]. Approximation [265, 593, 269, 345, 927, 1260, 794, 1259, 875, 148, 1225, 764, 827, 1227, 1040]. Approximations [266, 318, 54, 1357, 314, 1416]. Arbitrarily [456, 763, 261]. Arbitrary [1297, 205, 344, 368, 570, 509, 1332, 51, 1390, 623, 1033, 258, 1366, 1096, 1244]. Area [591]. Arising [816, 1426]. Array [688]. Arrays [858, 1076]. Aspects [532, 274, 479]. Assemblies [65]. Associated [1323, 106, 693, 42]. Associative [743]. assumption [1118]. Asymmetrical [575]. Asymptotic [680, 342, 1104, 273, 141, 1043, 533, 510, 893, 1127, 1216, 709, 543, 277, 1298, 653, 912, 1429, 930, 1094, 63, 458, 51, 135, 361, 539, 73, 713, 1037, 1160]. Asymptotically [968]. Atmosphere [862, 1088]. Atom [828, 240, 1170, 597, 794, 1259]. Atomic [501, 1003, 218, 1409, 776, 153, 1313, 336, 335, 694, 38]. Atoms [46, 712]. Attenuation [350, 781]. Attractive [28, 136, 169, 1316, 982, 280]. Author [1431]. Autocorrelation [1191, 1255, 498]. Auxiliary [1131]. Average [1426]. Averages [393, 603, 685, 1210]. Axes [250]. Axially [1201]. Axiomatic [779, 809, 300, 405, 406]. Axisymmetric [561].

B [460]. Background [1056]. Backgrounds [1000]. Backward [185]. Baker [959, 1009]. Banach [437, 142]. Bands [972]. Bare [156]. Bargmann [683]. Baryon [1108, 97]. Based [1289]. Bases [292, 546, 998, 408, 941, 1267]. basi [1025]. Basic [718]. Basis [70, 320, 101, 102, 103, 1373, 1273, 1130, 1, 5, 83, 84, 263, 112, 158, 825, 869, 134, 1353, 1433, 539, 1148, 923, 1399, 1400]. BBGKY [193, 795]. BCS [384]. be [155]. Beam [490, 625]. Bég [119]. Behavior [1104, 1061, 1293, 273, 724, 82, 488, 552, 722, 303, 277, 948, 1252, 1303, 24, 104, 539, 728, 713, 1037, 1160]. Bergman [244]. Bernoulli [498]. Bessel [1348, 1071, 1312]. Best [1238]. Bethe [1403, 697, 1161, 535, 1055, 1209, 355, 476, 380, 805, 849]. Between [726, 173, 46, 1380, 578, 62, 921, 240, 542, 834, 1089, 508, 628, 9, 276, 517, 745, 336, 335, 772, 1065, 235, 712]. Beyond [345]. Bi [515]. Bi-Orthogonality [515]. Bianchi [1347]. Bias [1050, 1007]. Bifurcation [847]. Binding [1420, 1310]. Binomial [1401, 1153]. Bipolar [496]. Bloch [337, 972]. Bodies [748]. Body [293, 1415, 1414, 107, 169, 387, 855, 608, 1282, 651, 1107, 49, 788, 1382, 822, 164, 61, 54, 1357, 1412, 976, 1416, 669, 1006, 1337, 46, 712, 962, 655]. Body-Centered [655]. Bogoliubov [698, 528, 899]. Boltzmann [1120, 542, 530, 121, 724, 853, 450, 402, 785]. Bonded [1350, 454, 455]. Bondi [441, 762, 378]. Bootstrap [640, 1297, 97]. Bootstraps [548]. Born [827, 971, 927, 166, 287]. Born-Approximation [827]. Bose [1280, 881, 339, 715, 133, 1257, 64, 1359, 286]. Boson [1032, 1217, 48, 17, 1326, 1394, 902, 42, 1035, 1090, 1091]. Bosons [425, 127, 1274, 979]. Bound [76, 697, 549, 28, 136, 874, 222, 383, 13, 794, 1259, 353, 116, 741, 311, 1275, 1196]. Bound-State [697, 874, 222, 116]. Bound-States [383]. Boundary [1120, 821, 1104, 1119, 756, 612, 969, 1176, 1117, 785, 1281, 1052, 235]. Boundary-Value [1120, 821, 1104, 1119, 612, 1176, 785, 1281]. Bounded [749]. Bounds [1323, 318, 990, 1204, 866, 393, 940, 1004, 685, 173, 798, 1389, 1388, 1434, 723, 1190, 611, 1238, 1436]. Bra [407]. Bracket [910]. Branching [705, 607, 198]. Brans [1262]. Breaking [292, 546]. Breit [1038]. Bremmer [750]. Brillouin [282, 655]. Broglie [1278]. Broken [1257]. Brownian [192, 443, 942]. Bulk [1389, 1434]. Burgers [1010, 1200].

C [460]. Calculation [113, 507, 1282, 60, 554, 604, 982, 780, 1229]. Calculations [827, 661]. Calculus [88, 617, 1106, 329, 1370]. Campbell [1009]. Can [155]. Canonical [636, 720, 48, 715, 393, 1342, 770, 543, 155, 370, 1340, 1418, 1387, 1386, 1146, 312, 833, 1012, 803, 1145, 564, 1392]. Carrying [209]. Cartan [295]. Cartesian [242]. Cascade [820]. Case [320, 391, 1278, 1207]. Cases [5, 84]. Casimir [172]. Cauchy [983]. Causal [753, 535]. Causality [1240, 442, 841, 1103]. Cavity [171]. Cayley [1144]. Cell [1109, 1110]. Cells [59]. Center [664, 1027, 336, 919, 1038]. Centered [282, 655]. Central [975, 927, 1423]. Cerenkov [510, 893]. Certain [796, 548, 1104, 527, 685, 1210, 60, 396, 24, 689, 1281, 800]. CFP [652]. Chain [1068, 444, 998, 1151, 1305, 1306, 1159, 710, 1406, 498]. Chains [1341]. Chakrabarti [429]. Chandrasekhar [974]. Changes [1115]. Channel [842, 275, 1292, 1382]. Channels [221, 846]. Chapman [616]. Character [1410]. Characteristic [1341, 736, 1283, 820]. Characterization [592, 1396, 915]. Characterizing [512]. Charge [634, 57, 688, 396, 782, 1116, 165, 170, 410]. Charged [440, 549, 1301, 1093, 740, 1318, 189, 110, 989, 561, 1233, 747]. Chiang [1435]. Chiral [1208]. Choh [298]. Circular [157, 1101]. Class [504, 441, 1129, 343, 734, 766, 699, 929, 1430, 349, 811, 41, 17, 424, 544, 1139, 1111, 616, 702, 522, 71, 981, 1211]. Classes [883]. Classical [193, 341, 469, 1183, 1235, 926, 391, 270, 1168, 208, 505, 953, 121, 1152, 1282, 553, 755, 573, 574, 1291, 860, 1378, 1300, 871, 1111, 689, 691, 910, 1005, 1034, 1423, 443, 764, 426, 619, 757, 1234, 1344, 359, 737, 671, 1219, 898]. Classification [67, 493, 494, 1174, 1230]. Classifying [537]. Clebsch [678, 1028, 1029, 1286, 4, 262, 1320, 397, 90, 887, 403, 200, 824, 1124, 460, 674]. Clifford [565]. Closed [486, 1382, 399, 371, 1167, 435, 46]. Closed-Form [1167]. Closed-Shell [46]. Closure [654, 764]. Cluster [681, 149, 876, 366, 642, 8, 609, 491, 1359, 457, 150, 151, 812, 877, 37]. Cluster-Star [150, 151]. Clusters [463, 1263]. Co [1268]. Co-diagonal [1268]. Coated [1192]. Coefficient [1401, 1153, 1064, 350, 554, 575]. Coefficients [1029, 636, 199, 776, 811, 823, 1320, 397, 90, 703, 887, 1188, 200, 557, 558, 1028, 1286]. Coexistence [213]. Coherence [39, 754, 281, 774, 618]. Coherent [901, 512, 350, 898, 91, 1057, 1024, 1308]. Coherent-State [91, 1024, 1308]. Collapsing [416]. Collective [1364]. Collineations [1213]. Collision [85, 577, 1402]. Collisionless [759, 878, 1117]. Collisions [765]. Combinatorial [1163, 992, 1223, 1338]. Combined [766, 700]. Comment [1435, 1436]. Comments [1152, 936, 1255]. Common [550]. Commutation [720, 715, 1342, 370, 1340, 404, 793, 479]. Commutator [959]. Commutators [732, 1328]. Commuting [503]. Compact [1075, 615, 493, 494, 1041, 867, 868, 1174, 622, 1015]. Compactification [98]. Compass [180]. Compatibility [795]. Compensation [707, 1090, 1091]. Complementary [808, 1190]. Complete [679, 220, 24, 1174]. Completeness [473, 1314]. Complex [965, 319, 270, 272, 271, 273, 53, 511, 1058, 1284, 281, 1013, 1014, 1051, 745, 328, 836, 989, 988, 995, 257, 288]. Component [1135, 1390]. Components [701, 865]. Composite [920, 1119, 61]. Compton [506]. Concepts [201]. Concerning [1348, 754, 170]. Concise [448]. Condensate [1326]. Condensation [881, 715, 133, 1257, 1375]. Condition [587, 81, 1167, 1225, 1314, 1103]. Conditions [1325, 342, 28, 136, 1240, 584, 837, 1317, 371, 255, 967, 1052, 256, 660, 1270]. Conducting [593, 859, 688, 905]. Conductivity [226, 225, 448, 449, 350, 773, 434]. Cone [467, 1248, 958]. Cones [693]. Configuration [10, 945, 1149, 1095]. Configuration-Space [1149]. Configurations [1313, 1264, 623, 1189]. Confirmation [1177]. Conflicting [1309]. Conformal [1381, 1296, 668, 1237, 1211]. Conformal-Invariant [1381, 1296]. Conformally [562]. Conjecture [1256]. Conjugation [634, 257]. Connected [658, 972]. Connection [1319, 834, 1089, 508, 1364, 745, 1065, 173]. Consequent [253].

Conservation [392, 208, 505, 417, 864, 124, 453, 253, 1005, 430, 431, 170, 775, 1017]. Conserved [1205, 1333, 106, 1349]. Considerations [1235]. Consistency [584]. Consistent [10, 54, 1357, 1416]. Constant [766, 350, 830, 462, 781]. Constants [716, 818, 45, 1017]. Constructing [1059]. Construction [70, 484, 1001, 57, 183, 1005, 6, 174, 1030, 499, 666, 17]. Contact [921]. Contained [87, 1211]. Containing [218, 1204, 1277, 1436]. Contains [1284]. Content [850]. Continuation [680, 662, 255, 31, 1198]. Continuations [356]. Continuity [353, 1309]. Continuous [1427, 348, 1273, 17, 514, 846, 1362, 1400]. Continuous-Representation [17]. Continuum [389, 333]. Continuum-Ising [333]. Contracted [1085]. Contraction [708, 330]. Contribution [226]. Contributions [1322, 219, 1094]. Convergence [502, 750, 937, 971, 1024]. convergent [1238]. Converging [994]. Conversion [1186]. Convex [914, 709, 1244]. Coordinate [224, 779, 155]. Coordinates [218, 776, 936, 1311, 562, 1100, 1101]. Corben [117]. Core [926, 187, 582, 1146]. Corepresentations [1041]. Correction [614]. Corrections [953, 449, 1064]. Correlation [341, 814, 213, 735, 1326, 899, 309, 816, 33, 1263]. Correlations [530, 605, 606, 302, 966, 916, 1081, 375, 409, 361, 903]. Correspondence [851, 274]. Corresponding [522]. Cosets [1320]. Cosmological [471, 323, 399, 1000, 1277]. Cosmos [330]. Coulomb [389, 1132, 768, 346, 626, 1027, 496, 960, 336, 982, 673, 780]. Coulomb-Like [673]. Coulombic [79]. Counts [375]. Coupled [221, 223, 275, 65, 553, 1382, 316]. Coupling [638, 716, 602, 1102, 1405, 322, 178, 557, 1040, 1048]. Covariant [11, 1187, 392, 417, 509, 192, 1288, 430, 431, 1390, 1047]. Covering [569, 397]. Creeping [772]. Criteria [1239]. Criterion [715]. Critical [1339, 1293, 970, 751, 1303, 256]. Critical-point [1339]. criticality [1100, 1101]. Cross [118, 490, 1122, 51, 1194, 1437]. Cross-Section [118]. Crossed [1292]. Crossing [835, 1297, 1077, 1175, 368, 570, 1232, 212, 1065]. Crossing-Symmetric [835]. Crystal [922, 692, 1142]. Crystals [738, 739, 454, 455]. CTP [643]. Cubic [1269, 282, 655]. Current [720, 209, 1328, 791, 1031, 256, 925, 924]. Current-Carrying [209]. Currents [898, 924]. Curvature [439, 709, 1213, 1145]. Curved [1208, 772, 669]. Curves [349, 702]. Cut [1322]. Cut-Contributions [1322]. Cutoff [116]. Cutoffs [673]. Cuts [191]. Cycles [484]. Cyclic [1342]. Cyclotron [445]. Cylinder [157, 878]. Cylinders [416, 51]. Cylindrical [171, 432, 1192, 1101].

D [629, 460]. Dancoff [985]. Dangerous [707, 1090, 1091]. Data [1067]. Daughter [856]. Debye [401, 621, 670]. Decay [1397, 89]. Decomposition [152, 489, 711, 857, 737, 877]. Decompositions [789]. Decoupling [275]. Defined [1213]. Definite [988]. Definition [636, 446, 410]. Deformable [682]. Deformation [708]. Deformations [1129]. degeneracies [1408]. Degeneracy [1403, 445, 999]. Degenerate [470, 1108, 1254, 514, 283, 789, 1015, 1354, 497]. Degree [281, 774]. Delta [1274]. Delta-Function [1274]. Democratic [537]. Dense [193, 580, 298, 385]. Densities [721, 385, 988, 1263]. Density [854, 186, 1226, 463, 36, 37, 187, 724, 533, 601, 1224, 598, 1376, 91, 804, 517, 357, 675, 1351, 1263, 982, 1223]. Dependence [1056, 771]. Dependent [341, 593, 859, 1206, 529, 551, 831, 194, 206, 1210, 140, 1111, 1318, 1377, 627, 1018, 1113]. Derivation [1103, 447, 448, 609, 986, 707, 536, 5, 84, 402, 1352, 40, 181, 974, 1154, 1200]. Derivations [617]. Derivative [1213]. Derivatives [874, 845, 853]. Described [1260]. Description [748, 509, 1291, 1162, 125, 1251, 942]. Descriptions [896]. Determinant [267]. Determinantal [846]. Determinants [1239]. Determination [842, 642, 383, 1249, 375, 1122, 1437]. Determine [591]. Detonations [994]. Development [392, 173]. Development-Evaluation [173]. Developments [55]. Deviations [1159]. Diagonal [133, 91, 1024, 1308, 556, 1268]. Diagram [68, 254]. Diagrams [113, 56, 349, 707, 79, 702, 1136, 1090, 1091]. Dicke [1262]. Didactic [775]. Dielectric [1295, 350, 1299, 830, 1192, 781]. Dielectric-Coated [1192]. Dielectrics [1247]. Difference [1182]. Different [517]. Differential [500, 223, 141, 23, 1023, 115, 1138, 819, 400, 850, 1167, 1122, 329, 803, 1133, 974, 1437]. Differentiation [1009, 676]. Difficulties [131, 580]. Diffraction [1060, 1374, 171, 1356, 709, 653, 137, 1299, 1429, 1383, 669, 261, 171]. Diffuse [975]. Diffusion [1365, 1180, 1100, 1101]. Dilute [576]. Dimension [1414, 52, 280, 816, 890]. Dimensional [123, 1350, 68, 1415, 1119, 1112, 1011, 215, 1170, 1099, 596, 391, 880, 645, 394, 955, 1185, 651, 41, 1171, 788, 1019, 582, 425, 858, 231, 1076, 1417, 1146, 495, 690, 232, 723, 134, 254, 1142, 1330, 563, 882, 861, 181, 333, 1230, 1020, 1274, 1100, 1101]. Dimensionality [207, 1066]. Dimensions [1138, 1307, 189, 747, 27]. Dimer [1242, 487, 1185, 844, 1368]. Dimers [939, 843]. Dipoles [418]. Dirac [1155, 1428, 1392, 295, 752, 1250, 533, 1140, 111, 910, 44, 146, 407, 801, 882]. Direct [489, 301, 370, 1157]. Direct-Product [370]. Directed [390]. Direction [1112]. Disconnected [120]. Discontinuities [829, 513]. Discontinuity [1179]. Discontinuous [43]. Discrete [389, 569, 853, 5, 84, 279, 789, 497, 286, 1399]. Discrete-Continuum [389]. discussion [1203]. Disordered [1142, 1426]. Dispersion [583, 738, 1201]. Dispersion-Energy [1201]. Dissimilar [684]. Dissipation [813]. Distance [62, 573, 1300]. Distorted [827]. Distorted-Wave [827]. Distortions [686]. Distribution [1220, 469, 363, 714, 953, 1169, 81, 880, 268, 755, 246, 575, 1007, 1284, 604, 1031, 1402, 660]. Distributions [899, 396, 326, 1164, 135, 695, 1047]. Disturbances [1169]. Divergences [898, 897, 673]. Divergent [116]. Domain [547, 347]. Domains [1119]. Doppler [439]. Double [76, 700, 647, 755, 204, 23, 256]. Double-Time [755]. Drop [561]. Drukarev [752]. Drum [540]. Duality [814]. Dumbbell [1417]. Dumbbells [858, 1076]. during [330]. Dust [471, 839, 399, 20]. Dust-Filled [471, 399, 20]. Dyadic [918]. Dynamical [131, 907, 442, 507, 509, 21, 864, 850, 1073]. Dynamics [1269, 704, 1260, 49, 1378, 956, 396, 1308, 1116, 233, 727, 852].

Edge [1429]. Effect [506, 507, 143]. Effective [350, 1382, 830, 1389, 1388, 1434]. Effective-Range [1382]. Effects [501, 1003, 920, 806, 1080, 1143, 875, 1420, 32, 827]. Ehrenfest [589]. Ehrenfests [1180]. Eigenfunction [693]. Eigenfunctions [654, 1231, 1207]. Eigenstates [128, 1351]. Eigenvalue [613, 784, 1126, 545, 567, 999]. Eigenvalues [766, 173, 309, 723, 1362, 1436]. Eight [645]. Eight-Dimensional [645]. Einstein [766, 1363, 1036, 1333, 715, 1087, 1253, 325, 656, 160, 159, 72, 894, 1347, 286, 1227, 1002, 672, 382, 20, 260]. Elastic [211, 297, 390, 927, 145, 137, 1058]. Elastic-Wave [1058]. Electric [773, 969, 104]. Electric-Field [969]. Electrical [1388, 434]. Electricity [306]. Electro [806, 1080, 1143]. Electro-Optical [806, 1080, 1143]. Electrodynamics [241, 156, 1300, 612, 1346, 129, 313]. Electromagnetic [632, 731, 593, 859, 1279, 1206, 1128, 529, 1099, 1082, 418, 209, 709, 246, 1079, 1299, 1318, 738, 739, 878, 1117, 160, 184, 1288, 146, 976, 332, 623, 1222, 728, 261, 905]. Electromagnetism [272, 656]. Electron [1295, 1206, 414, 1235, 599, 930, 576, 456, 763, 903, 1088]. Electron-Phonon [414]. Electrons [542, 153, 433]. Electrostatic [945]. Elementary [1171, 287, 288]. Elements [1380, 796, 1134, 94, 768, 776, 746, 659, 6, 174, 361, 786, 981, 1133]. Ellipsoids [560]. Elsasser [435]. Embedding [343, 641, 607]. Embeddings [909]. Emission [568, 241, 575]. Empty [343, 742, 637]. Encountered [913]. Energies [383, 1190]. Energy [588, 632, 1204, 854, 766, 740, 584, 82, 488, 552, 183, 1072, 1023, 1224, 173, 24, 741, 892, 745, 662, 777, 961, 988, 559, 560, 1351, 458, 667, 1310, 1194, 775, 1201, 1436]. Energy-Integral [1023]. Energy-Level [1224]. Energy-Momentum [775]. Ensemble [35, 393, 1050, 188, 1007, 1146]. Ensembles [53, 543, 122]. Enskog [616]. Entropy [1008, 454, 461]. Enumeration [464]. Envelope [1136]. Envelopes [798]. Enveloping [521]. Equal [1328]. Equal-Time [1328]. Equation [1084, 1403, 750, 697, 954, 295, 1161, 752, 267, 414, 542, 1182, 873, 121, 243, 77, 194, 751, 450, 23, 115, 1171, 229, 535, 1055, 1209, 1321, 706, 654, 536, 819, 648, 1140, 545, 744, 610, 400, 616, 310, 1016, 840, 402, 1177, 1167, 723, 476, 114, 785, 327, 44, 1255, 1010, 380, 849, 358, 913, 147, 801, 949, 978, 1113, 1392, 1233, 1244, 1017, 1200]. Equations [500, 293, 1323, 193, 468, 835, 197, 1322, 813, 11, 220, 548, 1343, 526, 527, 1250, 907, 223, 758, 834, 1061, 1179, 683, 734, 766, 1363, 141, 106, 1283, 875, 422, 1087, 1282, 1253, 1371, 1372, 305, 1023, 491, 424, 1384, 1278, 1138, 656, 1139, 284, 658, 117, 847, 160, 159, 72, 1150, 567, 1049, 146, 1039, 619, 1309, 803, 1347, 1227, 1002, 1383, 483, 671, 673, 877, 20, 1270, 1411, 1078, 974, 1088, 1036, 1133, 331]. Equilibrium [419, 1260, 940, 192, 574, 122, 425, 576, 561, 621]. Equivalence [482]. Equivalent [391, 590, 153, 1189]. Ergodic [722, 1195, 1053]. Errata [83, 84, 175]. Erratum [546, 262, 1405, 436, 547, 1434, 263, 763, 174, 1433, 437]. Error [318, 940, 1004, 1238]. Estimates [1043, 73]. Euclidean [1013, 1014, 1051, 331]. Euler [307, 803]. Evaluating [46, 712]. Evaluation [635, 797, 3, 436, 1271, 1210, 173, 401, 694, 1344, 621, 670, 139, 1109]. Evolution [243, 899]. Exact [1084, 1297, 766, 97, 451, 1317, 424, 1111, 1318, 858, 42, 1076, 284, 79, 128, 967, 27, 1231, 801, 695, 1277, 1113]. Exactly [247, 1035]. Example [1273, 353]. Exceptional [1187, 824, 674]. Exchange [534, 554, 380]. Excitation [679, 74]. Excitations [551]. Excited [232]. exciton [1229]. Excluded [620]. Exhibit [1002]. Existence [468, 835, 963, 756, 937, 486, 629, 26, 371, 847, 519, 1281, 480, 1017]. Expanding [931, 249, 562]. Expansion [681, 149, 62, 633, 821, 267, 1361, 296, 932, 463, 132, 1292, 225, 609, 248, 1359, 648, 795, 616, 765, 1024, 1246, 252, 496, 51, 1263, 89]. Expansions [876, 642, 141, 533, 543, 653, 1304, 693, 991, 771, 40, 877, 1038, 1109]. Expectation [318, 154, 381]. Experiment [921]. Experimental [1249, 1177]. Explicit [239, 823, 1016, 1370]. Exploding [417]. Explosives [994]. Exponential [35, 664, 249, 1009, 676]. Expressions [401, 621, 670]. Extended [748, 1023]. Extension [634, 579, 1022, 398, 933]. Extensions [644]. Extent [596]. Exterior [416]. External [928, 508, 615, 606]. Extraction [452].

Face [282]. Face-Centered-Cubic [282]. Factor [876]. Factor-Cluster [876]. Factorization [541, 1022, 1058]. Factors [1422, 123, 720, 642, 408]. Faddeev [293, 1321, 1270]. Faddeev-Type [293]. Failure [366]. Features [1364]. Feinberg [29]. Fermi [436, 1335, 528, 3, 64, 1224, 353]. Fermi-Gas [528]. Fermion [753, 156, 876, 186, 788, 491, 614, 42, 1263, 861, 877, 1339, 1118]. Fermions [52, 280, 127]. Ferromagnet [40, 237]. Ferromagnetism [247]. Ferromagnets [605, 606, 1332, 916]. Feynman [415, 113, 1043, 1424, 488, 552, 56, 349, 41, 702, 1311, 254, 144, 139]. Field [66, 950, 957, 578, 1295, 96, 632, 731, 677, 342, 1206, 1011, 442, 1256, 1301, 734, 270, 766, 889, 1093, 928, 50, 209, 383, 572, 447, 214, 769, 422, 1342, 227, 1127, 1216, 1253, 148, 512, 246, 17, 216, 1318, 782, 283, 492, 930, 969, 42, 1272, 284, 1116, 375, 456, 763, 117, 160, 72, 126, 328, 836, 1262, 1034, 525, 146, 63, 426, 519, 520, 619, 852, 989, 961, 988, 285, 862, 191, 856, 331, 480, 897, 672, 1369, 972, 20, 260, 410, 1196, 1172, 1096, 979, 1233, 1215]. Field-Plus [578]. Field-Theoretic [950, 447]. Field-Theoretical [410]. Field-Theory [957, 191]. Fields [7, 963, 845, 271, 1082, 418, 872, 64, 1067, 606, 1342, 649, 686, 1127, 1216, 1298, 1079, 124, 761, 810, 773, 495, 690, 253, 184, 1288, 1234, 43, 1135, 360, 1390, 672, 71, 334, 256, 260]. Figure [59]. Figures [561]. Filled [471, 399, 20]. Films [881]. Finding [1185, 232]. Fine [501, 1003, 766]. Fine-Structure [766]. Finite [813, 993, 596, 740, 883, 1243, 394, 1228, 1199, 1373, 543, 230, 582, 1146, 263, 112, 376, 659, 32, 1214, 961, 162, 386, 827, 1192, 972, 1399, 1400]. Finite-Dimensional [394]. Finite-Energy [961]. Finite-Range [827]. Finiteness [749]. First [1132, 864, 310, 326, 803, 1227, 1229]. First-Order [1132, 310, 803, 1227]. First-Passage [326, 1229]. Five [82, 1019, 1020]. Five-Dimensional [1019, 1020]. Five-Point [82]. Fixed [82, 423, 1072, 545, 745, 458, 667]. Fixed-Angle [1072]. Flat [562, 968]. Flexible [714]. Flow [611]. Fluctuation [21]. Fluctuations [854, 122, 601, 461]. Fluid [596, 601, 582, 1146, 1181, 1002, 1109, 1110]. Fluids [1350, 1316, 164, 1277, 671]. Fock [1280, 631, 1260]. Fokker [578, 77, 648, 1177, 165, 577]. Fokker-Type [578]. Force [682, 1423, 128]. Forces [596, 164, 1236, 1218, 982]. Form [731, 720, 250, 401, 1167, 1347, 771, 621, 820, 670, 435]. Formal [807, 23, 991, 1360]. Formalism [18, 34, 290, 291, 834, 1409, 154, 1394, 355, 405, 406, 477, 407, 285, 288, 410, 1155, 1428, 1215, 925, 924, 673]. Formalisms [876]. Formalization [201]. Forming [131]. Forms [346, 518, 1145]. Formula [634, 29, 959, 1382, 1009, 150]. Formulas [743, 448, 1342, 351, 792, 1021]. Formulation [950, 1325, 1187, 392, 506, 384, 837, 815, 340, 85, 770, 1300, 111, 495, 690, 357, 525, 1336, 92]. Formulations [214]. Forward [476]. Foundations [904, 300]. Four [293, 753, 1230, 1149]. Four-Body [293]. Four-Dimensional [1230]. Four-Fermion [753]. Four-Particle [1149]. Fourier [62]. Fourth [39, 409]. Fourth-Order [39, 409]. Fractional [1188]. Fractionally [969]. Frames [295]. Fredholm [305, 483]. Free [1295, 271, 704, 106, 183, 124, 253, 357, 1088, 1096, 979]. Free-Electron [1295, 1088]. Free-field [979]. Free-Particle [106]. Frequency [1191, 1157, 1377, 944, 1158]. Fresnel [706]. Friedmann [1079, 1288]. Froissart [255]. Fronted [909]. Function [341, 1191, 593, 859, 76, 469, 529, 842, 814, 953, 928, 81, 880, 82, 853, 268, 626, 735, 513, 244, 246, 788, 535, 491, 248, 614, 746, 355, 783, 78, 498, 919, 862, 627, 861, 181, 1390, 1274, 1109]. Functional [1220, 1219, 1032, 845, 132, 340, 574, 230, 1024, 617, 329, 1181, 820, 139, 1411]. Functionals [769, 88, 33, 151, 897]. Functions [1084, 389, 86, 813, 39, 1068, 1220, 1348, 802, 1361, 1128, 296, 363, 1099, 932, 1108, 485, 873, 369, 213, 145, 1212, 755, 245, 324, 1320, 1304, 1058, 899, 1248, 1197, 798, 1071, 309, 1385, 1013, 1014, 1051, 659, 356, 1044, 816, 1255, 522, 1344, 604, 1312, 826, 539, 565, 660, 1113, 974, 1338, 1254, 962]. Fundamental [1213, 868]. Further [881, 1401, 801].

G [824]. Galilean [1106]. Galilei [592, 1292, 1012, 196, 236, 539, 624]. Gamma [294]. Gamma-Ray [294]. Gas [193, 1295, 414, 528, 339, 880, 530, 121, 243, 724, 853, 788, 451, 554, 902, 353, 805, 317]. Gases [580, 298, 36, 37, 187, 189, 385, 459]. Gauge [384, 872, 782, 314, 1369]. Gaussian [712]. Gel’fand [546, 292, 1256, 354]. Geminal [186]. General [62, 1325, 219, 19, 295, 1205, 416, 1082, 736, 392, 417, 704, 837, 839, 1265, 646, 915, 1086, 276, 450, 986, 1213, 997, 916, 706, 397, 284, 996, 1188, 326, 659, 1346, 557, 793, 1347, 667, 563, 1125, 168, 235, 625, 1155, 1203, 1073, 164, 1215]. Generalization [298, 1178, 555, 840, 846, 1033, 978, 1338]. Generalizations [1016, 359, 1017, 1200, 1367]. Generalized [767, 1319, 1358, 389, 1280, 441, 762, 881, 1361, 363, 369, 121, 133, 1257, 970, 1331, 453, 1329, 402, 253, 477, 1049, 1164, 1402, 673, 1411, 1062]. Generating [1084, 1113, 230, 1344, 151]. Generators [802, 657, 786, 884, 958]. Geodesic [864, 334]. Geometric [935, 228, 167, 826]. Geometrical [787]. Geometries [1257, 310]. Geometrization [272, 126, 328, 836, 1262]. Geometry [993, 954, 19, 873, 850, 238, 1244]. Ginzburg [847]. Given [1206, 802, 742, 13, 6, 174]. Global [963, 537, 430, 431]. Glory [1158]. Gordan [678, 1028, 1029, 1286, 4, 262, 1320, 397, 90, 887, 403, 200, 824, 1124, 460, 674]. Gordon [766, 907, 744, 1251, 146]. Grand [268, 1146]. Graph [513, 991, 462, 459]. Graphs [540, 415, 1315, 464, 24, 465, 992, 710]. Gravitation [986, 426]. Gravitational [7, 963, 983, 917, 931, 108, 325, 1236, 930, 1074, 495, 690, 233, 184, 27, 432, 180, 360, 719, 71, 334, 1215]. Greatly [517]. Green [593, 859, 1068, 1128, 529, 1099, 932, 485, 928, 873, 943, 1178, 755, 626, 245, 535, 491, 1304, 783, 862, 627, 861]. Gribov [255]. Griffiths [916]. Ground [1415, 1204, 740, 1217, 1364, 1165, 1190, 1310, 1339]. Ground-State [1204, 740, 1217, 1190, 1310]. Group [70, 832, 1422, 886, 1427, 828, 86, 592, 964, 441, 762, 1112, 1343, 1205, 320, 388, 205, 344, 641, 569, 815, 597, 507, 700, 984, 1355, 1379, 1292, 474, 1273, 1019, 397, 202, 1, 5, 83, 84, 657, 492, 231, 1140, 1041, 1013, 1014, 1051, 1313, 701, 865, 354, 825, 869, 607, 659, 378, 134, 1044, 710, 833, 1012, 824, 727, 44, 557, 558, 711, 1135, 630, 882, 80, 196, 236, 539, 624, 565, 1020, 786, 362, 1078, 1237, 1025, 1211, 938, 1121, 923, 1015, 1354, 958, 1267, 1211]. Group-Theoretic [1205]. Group-Theoretical [828, 202]. Groups [70, 832, 239, 993, 413, 199, 1156, 638, 1168, 641, 368, 570, 113, 369, 883, 1243, 647, 489, 1228, 705, 1059, 120, 1075, 514, 887, 810, 615, 493, 494, 1041, 556, 1174, 911, 1246, 376, 622, 312, 497, 521, 1030, 163, 162, 386, 499, 1344, 170, 737, 198, 674, 1073, 1062, 1015, 1354]. Growth [1310, 1275]. Guiding [664]. Gyration [714]. Gyroelectric [14]. Gyrotropic [1383, 235].

H [853]. Haag [760, 1048]. Half [273, 688, 251, 515, 1421, 862, 1383, 1133]. Half-off-shell [1133]. Half-Plane [273, 1383]. Half-Planes [688]. Half-Space [251, 515, 1421, 862]. Hamilton [1250, 1052]. Hamiltonian [201, 128, 1038, 590, 1419, 1331, 1165, 1284, 799, 1214, 361, 1042, 410]. Hamiltonians [391, 753, 74, 800]. Hard [926, 339, 880, 49, 582, 554, 1146, 805, 187]. Hard-Core [926, 582, 1146, 187]. Hard-Sphere [339, 554]. Harmonic [86, 240, 450, 1387, 1111, 1318, 231, 746, 1197, 659, 1044, 692, 1142, 1377, 944, 938]. harmonic-oscillator [938]. Harmonics [817, 496, 771]. Hartree [1260]. Hausdorff [959, 1009]. Having [145, 58, 1041]. Heisenberg [1425]. Helicity [1290, 1094, 1121]. Helicon [26]. Helmholtz [750, 496]. Hepp [698]. Herglotz [668]. Hermite [88, 89, 1189]. Hermitian [538, 1065, 800]. Heterogeneous [1389, 1388, 1434]. Hidden [1046]. Hidden-Variable [1046]. Hierarchies [617, 721]. High [588, 345, 584, 64, 82, 488, 552, 1064, 1072, 24, 892, 1157, 1158, 777, 1194]. High-Energy [588, 584, 82, 488, 552, 1072, 24, 892, 777, 1194]. High-Frequency [1157, 1158]. High-Order [64]. High-Temperature [345]. Higher [11, 215, 1293, 959, 57, 395, 449, 120, 774, 377, 356]. Higher-Dimensional [215]. Higher-Neighbor [1293]. Higher-Order [959, 395, 449, 774, 356]. Higher-Spin [11]. Hilbert [16, 616, 525, 1276]. Hill [327]. Histories [1311]. Hole [1364]. Holomorphy [798]. Holonomy [810]. Homeomorphically [464]. Homogeneous [70, 1407, 918, 471, 323, 556, 869, 160, 159, 72, 585, 539, 624]. Homology [113]. Hook [1163]. Hopf [905, 1358, 1383]. Horizons [1156]. Houston [282, 655]. Huygens [1123]. Hybrid [1027]. Hydrodynamics [245]. Hydrogen [828, 240, 1170, 597, 454, 455, 38]. Hydromagnetic [474, 195]. Hyperbolic [1138]. Hyperboloids [693]. Hypergeometric [1361, 356, 1338, 1198]. Hyperplane [1290]. Hyperspheres [484]. Hypersurface [742]. Hypersurfaces [736, 1283]. Hysteresis [663].

Ice [454]. Icecream [467]. Icecream-Cone [467]. Ideal [902, 189]. Identical [1204, 1436]. Identifying [465]. Identities [1401, 1153]. Identity [453]. II [1405, 1113, 34, 1029, 1428, 1003, 1391, 266, 784, 241, 1202, 136, 1080, 1096, 171, 1338, 186, 271, 505, 1175, 387, 570, 1102, 984, 725, 37, 552, 606, 102, 1430, 1259, 721, 1216, 1372, 276, 1252, 513, 1091, 555, 1317, 947, 514, 1306, 280, 494, 655, 373, 1014, 690, 1017, 30, 865, 1400, 455, 1158, 1354, 172, 836, 406, 313, 747, 128, 329, 1181, 431, 288, 1416, 523, 459, 409, 672, 1110, 236, 877, 1042, 1006, 1101, 1020, 259, 1267, 1037, 925, 1070, 38]. III [290, 1126, 1203, 1143, 272, 962, 1355, 187, 103, 693, 1051, 1229, 1200, 317]. Imbedding [55, 813, 840, 1207]. Impenetrable [425]. Implications [377]. Implies [66, 775]. Improved [169, 387, 962, 10]. Impurities [516]. Impurity [728]. Incidence [235]. Including [875]. Incoherent [160, 72]. Incompressible [1181]. Increase [1310, 1275]. Indefinite [338, 242, 348, 129]. Index [1431, 1432, 1041, 781]. Induced [568, 638]. Inelastic [182]. Inequalities [197, 1226, 916, 946]. Inequality [190, 144, 234]. Infinite [68, 684, 36, 1419, 451, 860, 140, 688, 773, 1135, 433, 890, 1390]. Infinite-Component [1135, 1390]. Infinite-Dimensional [68]. Infinite-Spin [890]. Infinitesimal [657]. Infrared [338, 898]. Inhomogeneous [813, 975, 750, 107, 893, 1326, 397, 536, 1162, 1, 5, 83, 84, 657, 556, 867, 868, 1174, 701, 865, 878, 44, 1030, 669, 1105]. Inhomogenizations [965, 964]. Initial [1325, 778, 372, 373, 1167, 1052]. Initial-Value [778, 372, 373]. Inner [1168, 726]. Instabilities [625]. Instability [1322, 1147]. Instance [999]. Instantaneous [573, 1300]. insulation [903]. Insulators [95]. Integral [211, 951, 1323, 197, 1358, 218, 526, 527, 365, 797, 145, 566, 794, 1259, 1342, 305, 1023, 244, 864, 229, 1384, 730, 610, 518, 1039, 913, 147, 982, 671, 949, 771]. Integral-Operator [566]. Integrals [412, 389, 1220, 484, 635, 1032, 1043, 1424, 3, 436, 1271, 488, 552, 41, 115, 60, 554, 307, 116, 1027, 336, 335, 694, 919, 435, 139, 1193, 1202]. Integrated [575]. Integration [1321, 1181, 1411]. integro [974]. integro-differential [974]. Intense [456, 763]. Intensity [575, 51]. Interacting [951, 542, 508, 52, 280, 259, 317, 1215]. Interaction [753, 889, 10, 1265, 13, 1425, 1316, 1303, 164, 1305, 1306, 1284, 143, 14, 1034, 1137, 459, 1274]. Interactions [834, 1293, 1269, 1396, 1395, 872, 929, 1430, 1291, 788, 1159, 79, 945, 27, 46, 712, 944, 1095]. Interelectronic [218, 776]. Interface [517, 772]. Internal [766, 641, 368, 570, 508, 700, 179, 513, 810, 615, 824, 1406, 852, 315, 1393, 80]. Internal-Labeling [1393]. Interpolation [822, 792, 1021]. Interpretation [550, 900, 1123, 977]. Intersect [604]. Intrinsic [12, 1140, 895, 74]. Invariance [364, 596, 208, 505, 643, 179, 936, 582, 1146, 1005, 1066, 775, 1237, 1096, 979, 1118]. Invariant [55, 1381, 813, 19, 1296, 384, 1292, 818, 943, 1371, 1372, 730, 693, 250, 840, 1249, 1207, 212, 1078]. Invariants [1183, 365, 1093, 424, 1111, 799, 867, 868, 22, 172, 1030, 499, 1042, 968]. Inverse [681, 506, 1114, 1071, 745, 1360, 458, 667, 1312]. Inverse-Scattering [458]. Inversion [716, 305, 1173, 150, 151]. Involutional [1289]. Involving [776, 946, 358]. Ionization [38]. Ions [46, 712]. Irreducibility [1398]. Irreducible [70, 504, 366, 906, 600, 301, 464, 1019, 938, 746, 493, 494, 1041, 701, 865, 354, 87, 200, 427, 163, 386, 711, 666, 857, 1020, 790, 1070, 1267, 196]. Irregularities [914]. Ising [333, 123, 591, 1068, 68, 1293, 1151, 487, 725, 1261, 605, 606, 1332, 966, 9, 304, 276, 609, 916, 1081, 1272, 1329, 1137, 409, 563, 1125, 40, 181, 890]. Ising-Model [966]. Isometries [839]. Isoparity [630]. Isoperimetric [1184, 59]. Isoscalar [1422, 408]. Isospectral [540]. Isospin [650, 1232, 1208, 289, 1096, 979]. Isotopic [257]. Isotopically [1142]. Isotropic [975, 1036, 627]. Iterative [1054]. IV [291, 1379, 771]. IX [173, 1347].

J [1437, 639]. Jost [522]. Junction [837].

Kahan [555]. Källén [114]. Kampen [1103]. KDP [455]. Kepler [807, 936]. Kernel [1342]. Kernels [658]. Kerr [579, 733, 973, 109]. Ket [407]. Killing [1156]. Kind [1196]. Kinematic [1425, 1094]. Kinematics [1069, 202, 203, 870]. Kinetic [756, 937, 414, 581, 580, 298, 875, 536, 1162, 741, 1352, 1255, 665, 1166, 1105]. Kinetics [1076, 1417]. Kinks [421, 1089]. Kirillov [1256, 1241]. Kirkwood [1323]. Klein [907, 766, 744, 1251, 146]. Knudsen [611]. Kondo [1271]. Korteweg [1016, 1017, 1200]. Kronecker [857]. l [460]. Labeling [906, 1393]. Ladder [936, 952, 1398, 79]. Lagrange [787, 803]. Lagrangian [201, 96, 613, 784, 1126]. Laminas [1141]. Landau [1203, 100, 349, 847, 702, 749]. Laplace [680]. Lapse [1325]. Large [756, 1194]. Larger [702]. Laser [456, 763]. Lasers [317]. Late [89]. Lateral [772]. Laterally [773]. Lattice [1191, 939, 1287, 1269, 642, 843, 226, 225, 448, 449, 9, 276, 1303, 451, 306, 655, 1329, 454, 455, 992, 991, 805, 459, 409, 462]. Lattices [1269, 282, 30, 1426, 944, 1229]. Law [986, 1005, 1189]. Laws [438, 392, 208, 505, 417, 864, 124, 453, 253, 430, 431, 1017]. Leading [349, 117, 702]. Lebowitz [398]. Lee [920, 340, 1397, 1298, 311, 891, 985, 147, 1048]. Left [273]. Legendre [1384]. Lehmann [285]. length [1100, 1101]. Lenz [626]. Lepton [50]. Level [1224, 1165]. Lichnerowicz [585]. Lie [239, 964, 802, 1256, 644, 1243, 394, 1371, 1372, 1213, 324, 708, 615, 493, 494, 1041, 691, 1246, 607, 312, 22, 172, 1124, 1030, 499, 630, 624, 198]. Lie-Algebra [964]. Lie-Algebra-Invariant [1371, 1372]. Lie-Group [964]. Lie-like [1243]. Lifetime [85]. Light [975, 1248, 575, 958]. Light-Cone [1248]. Like [673, 438, 1243]. Limit [602, 1102, 1405, 64, 871, 36]. Limitable [1073]. Limiting [346]. Line [742]. Linear [197, 594, 542, 1341, 1226, 444, 1023, 115, 555, 1305, 1306, 1159, 263, 112, 158, 465, 764, 803, 577]. Linearized [648, 45, 785, 104]. Linearly [251]. Lines [399]. Linked [642, 491]. Linked-Cluster [642]. Liouville [1341, 1036, 723]. Lippmann [673]. Liquid [1335, 278, 561]. Liquid-Vapor [278]. Local [265, 266, 591, 1096, 979, 643, 1083, 915, 1321, 475, 1056, 1057, 791, 499, 1002, 192]. Localized [520, 1412]. Logarithmic [138, 224, 724]. Logarithmically [531]. Long [1269, 420, 133, 1316, 1303, 164, 385, 1218, 104, 1137, 459, 944]. Long-Range [1269, 420, 133, 1316, 1303, 164, 1218, 1137, 459, 944]. Long-Time [385, 104]. Loop [572, 254]. Loops [56]. Lorentz [83, 70, 152, 993, 1327, 1025, 179, 397, 1, 5, 84, 657, 1140, 701, 865, 869, 911, 1399, 1400, 254, 727, 44, 1135, 482, 882, 624, 1047, 362]. Lossy [1128]. Low [1425, 143, 40, 463]. Low-Density [463]. Low-Temperature [143, 40]. Lower [1204, 393, 741, 87, 611, 1436]. Lowering [710, 786]. Lowest [226, 614]. Lowest-Order [226, 614]. Lucas [1319].

Macroscopic [1240, 461]. Magnetic [1295, 1301, 444, 1093, 486, 1152, 606, 1272, 189, 308, 1388, 1116, 663, 146, 165, 1233, 903]. Magnetic-Monopole [1116]. Magnetization [1272, 143]. Magnetoelastic [95]. Magnetogasdynamics [829]. Magnetostatic [661]. Majorana [1135]. Majorization [415]. Manifestly [509]. Manifolds [362]. Many [1032, 876, 1179, 842, 1151, 107, 1217, 48, 855, 299, 1115, 49, 155, 822, 491, 614, 164, 42, 374, 54, 1357, 1035, 1412, 1416, 861, 877, 1339]. Many-Body [107, 855, 49, 822, 164, 54, 1357, 1412, 1416]. Many-Boson [1032, 1217, 48, 1035]. Many-Channel [842]. Many-Fermion [876, 491, 614, 42, 861, 877, 1339]. Many-Neighbored [1151]. Many-Particle [1179, 299, 1115, 155, 374]. Mappings [955]. Marchenko [834]. Markovian [419]. Mass [634, 743, 156, 320, 388, 700, 101, 102, 103, 513, 1055, 351, 5, 84, 263, 112, 158, 701, 865, 110, 1150, 315, 1234, 1390, 1237, 923]. Masses [733]. Massive [126, 380, 1215, 1078, 1366]. Massless [1283, 761, 879, 253]. Master [1130]. Matched [653]. Materials [1389, 1388, 1434]. Math [1437]. Mathematical [324, 116, 274, 46, 712, 111]. Matrices [35, 186, 1297, 1226, 1050, 36, 37, 187, 53, 628, 629, 1085, 598, 1376, 1289, 91, 804, 1007, 309, 1399, 1400, 357, 675, 1351, 1264, 1223, 570]. Matrix [1380, 796, 681, 294, 1134, 1319, 152, 842, 94, 643, 768, 368, 570, 1242, 420, 346, 981, 1364, 85, 533, 101, 102, 103, 395, 776, 227, 204, 1291, 305, 823, 1058, 452, 467, 1139, 1232, 844, 746, 765, 659, 1214, 538, 841, 6, 174, 217, 1263, 982, 289, 361, 1402, 258, 786, 410, 1193, 1202, 1203, 1018, 1144, 1133, 925]. Matrix-Valued [1058]. Matter [417, 595, 704, 947, 160, 72, 360, 259, 317]. Matthews [1411]. Maximal [579, 1163]. Maximum [1090]. Maxwell [766, 1343, 1363, 551, 1333, 1087, 278, 159, 672, 260]. Mayer [366]. Mean [463, 1008, 741, 1195]. Means [680, 1054, 617, 151]. Measure [955, 1251]. Measure-Preserving [955]. Measure-Theoretical [1251]. Measurement [439, 32]. Mechanical [131, 169, 387, 91, 154, 398, 130, 518, 1276, 942, 962]. Mechanics [950, 149, 589, 1220, 1063, 814, 366, 345, 208, 505, 446, 980, 48, 843, 509, 65, 699, 1265, 787, 590, 940, 1004, 188, 779, 809, 717, 721, 58, 755, 770, 573, 543, 574, 696, 1418, 1008, 871, 729, 111, 402, 1195, 691, 910, 1005, 405, 406, 477, 443, 764, 757, 359, 1131, 459, 234, 977, 1336, 1155, 1428, 1219, 1073]. Media [767, 859, 1365, 1245, 1128, 529, 1099, 684, 893, 517, 235]. Medium [593, 1279, 350, 140, 830, 773, 627, 1383, 781]. Meromorphy [269]. Mesodynamics [314]. Meson [1060, 440, 549, 845, 97, 1215]. Meson-Baryon [97]. Method [341, 294, 1391, 767, 1358, 821, 169, 387, 1242, 298, 321, 566, 590, 489, 1210, 1022, 696, 21, 864, 822, 1059, 555, 844, 575, 371, 282, 655, 991, 252, 885, 846, 764, 667, 985, 1302, 316, 261, 962]. Methods [984, 1355, 1379, 9, 276, 652, 537, 307, 93, 2, 175, 46, 712, 1219]. Metric [579, 733, 973, 338, 242, 348, 622, 110, 109, 129]. Metrical [306]. Metrics [416, 839, 1285, 1033, 411]. Metzner [378, 441, 762]. Microcanonical [188]. Microcausality [63]. Microscopic [581, 1105]. Microwave [14]. Mills [810]. Minimal [634, 545]. Minimum [901, 183, 1091]. Minkowski [98, 895, 288]. Mirror [832]. Miscellanea [524]. Mixed [600, 1221, 612, 1313]. Mixed-Shell [1221]. Mixing [191]. Mixture [463, 556]. Möbius [150]. Mode [1186]. Model [589, 123, 591, 1287, 68, 920, 440, 549, 994, 640, 1297, 1293, 528, 1077, 1175, 384, 419, 50, 507, 383, 340, 214, 725, 1261, 1375, 1397, 966, 1180, 609, 1298, 247, 1081, 216, 1307, 1272, 1123, 455, 311, 1034, 1352, 61, 1035, 285, 891, 985, 330, 89, 1137, 147, 409, 563, 1125, 191, 856, 40, 331, 337, 181, 333, 890, 999, 259, 317, 237, 410, 1048, 1073, 938, 1109, 1110]. Models [937, 584, 550, 918, 487, 1127, 1216, 97, 1303, 323]. Moderately [385]. Modes [470, 26, 621]. Modification [589, 79, 1052]. Modified [927, 1081].

Modulated [1093]. Modulation [1391, 1302]. Moduli [6, 174]. Modulus [1389, 1434]. Molecular [738, 739]. Molecules [805, 1201]. Moment [165]. Momenta [178, 745]. Moments [748, 325]. Momentum [1191, 973, 1001, 273, 1395, 118, 182, 779, 794, 1259, 244, 105, 1, 5, 83, 84, 263, 112, 158, 518, 793, 498, 792, 849, 217, 775, 1189]. Momentum-Space [244]. Momentum-Transfer [182]. Monomer [302]. Monopole [1116]. Monopoles [1152]. Moshinsky [546, 292]. Most [397, 789]. Motion [1301, 889, 445, 1093, 106, 192, 1282, 491, 45, 1423, 443, 619, 433, 1017, 1006]. Motions [637, 382, 565]. Mott [903]. Moving [593, 859, 1279, 1128, 529, 1099, 418, 688, 627]. Multichannel [1077, 1175]. Multicomponent [47]. Multidimensional [240]. Multigroup [1421]. Multimode [317]. Multiparticle [8, 1251]. Multiple [1097, 623, 1222, 1193]. Multiplicative [1365, 326]. Multiplicity [1168, 726, 1075, 615, 824, 1406]. Multipliers [787]. Multipole [94, 748, 547, 347, 1283, 325]. Multipoles [682, 418]. Multiregion [310].

n [503, 189, 629]. Nakanishi [190]. Narrow [1098]. Natural [1052]. Nature [228, 888]. Near [759, 1260]. Near-Steady [759]. Nearest [1261, 1305, 1306, 1159, 1007]. Nearest-Neighbor [1007]. Nearly [799, 1042]. Necessary [81, 1419, 816, 660]. Negative [1029, 1185]. Neighbor [1293, 1261, 1305, 1306, 1159, 1007]. Neighbored [1151]. Nelson [1154]. Neumann [111]. Neutrinos [935, 167, 125]. Neutron [294, 684, 751, 555, 251, 808, 1362, 1100, 1101, 1244]. Neutron-Transport [555]. Neutrons [1169]. Newman [818, 1349]. News [432]. Newton [438, 986]. Next [1261, 1305, 1306, 1159]. Next-Nearest-Neighbor [1261, 1305, 1306, 1159]. No [1040]. Noether [668]. Non [588, 419, 100, 394, 574, 305, 664, 525, 800]. Non-Equilibrium [574]. Non-Fredholm [305]. Non-Hermitian [800]. Non-Hilbert-Space [525]. Non-Landau [100]. Non-Markovian [419]. Non-Semisimple [394]. Non-uniform [664]. Non-Usual [588]. Nonanalytic [699]. Noncanonical [1253, 1248]. Noncommuting [248]. Noncompact [369, 1075, 514, 911, 497, 521, 1399, 1400, 1354]. Nondegenerate [544]. Nondispersive [627]. Nonequilibrium [980, 590, 724, 1004, 122, 692]. Nonexistence [520, 961]. Nonhomogeneous [767, 23, 400]. Nonideal [829]. Noninteracting [735]. Noninvariant [872]. Nonlinear [1120, 1391, 821, 1104, 1119, 1365, 854, 1258, 297, 390, 734, 889, 424, 1176, 1016, 1177, 785, 1346, 1034, 161, 519, 520, 619, 852, 989, 961, 988, 1281, 1302, 316]. Nonlocal [222, 929, 1430]. Nonorthogonal [1353, 1433]. Nonperiodic [1224]. Nonradiative [184]. Nonrandom [1007]. Nonrelativistic [593, 628, 534, 956, 203, 456, 763, 812, 780, 1103, 202]. Nonrenormalizable [265, 266, 422]. Nonsaturation [1236]. Nonuniform [1186, 373]. Nonuniqueness [1098]. Nonunitary [1134, 376]. Nonzero [320, 5, 84, 1272, 158, 701]. Nordsieck [337]. Normal [1180, 466, 371]. Note [412, 1063, 1401, 733, 935, 100, 571, 3, 436, 1153, 1059, 119, 250, 885, 233, 109, 378, 69, 434]. Nucleon [827, 168]. Null [632, 731, 637, 1067, 811, 495, 690, 996, 672, 260]. Number [631, 463, 13, 901, 868, 1275, 749, 1091]. Number-Phase [901]. Numerical [305, 1384, 913, 483].

Obeying [121]. Object [171, 914]. Oblique [235]. Observables [1404, 643, 779, 809, 1249, 1366]. Observations [521, 237]. Obstacles [261]. Obtained [766, 72]. Obtaining [864, 575, 991]. Occupation [858, 1076]. Odd [218, 1284]. ODLRO [881, 902]. Off [133, 662, 1133]. Off-Diagonal [133]. Off-Energy-Shell [662]. Old [1087]. On-Mass-Shell [101, 102, 103]. One [1350, 1415, 1414, 1119, 1099, 596, 391, 873, 880, 1038, 227, 651, 652, 788, 140, 654, 582, 425, 858, 492, 52, 280, 1076, 1417, 310, 1146, 232, 723, 1368, 254, 816, 747, 1142, 27, 1330, 563, 861, 890, 1274, 1100, 1101, 1172]. One- [1119, 1099, 1038]. One-Dimensional [1350, 1415, 596, 391, 880, 651, 788, 582, 425, 858, 1076, 1417, 1146, 232, 723, 1142, 1330, 563, 861, 1274]. One-Loop [254]. One-Parameter [492]. One-Particle [227, 652, 1172]. One-Speed [873, 140, 654, 310, 1100, 1101]. Onsager [9, 276]. Onset [902]. Open [980]. Operational [2, 175]. Operations [863]. Operator [1280, 94, 1083, 57, 566, 693, 92, 313, 379, 217, 577, 168, 258, 949]. Operators [131, 239, 503, 634, 318, 12, 1280, 296, 1001, 1054, 718, 936, 952, 555, 248, 1387, 1386, 1009, 279, 946, 556, 87, 74, 658, 710, 172, 518, 329, 1276, 386, 286, 676, 786, 981]. Optical [806, 1080, 1143, 900, 281, 774, 375]. Orbit [1021, 1040, 1095]. Orbitals [788, 336, 335]. Orbits [888, 308, 43]. Order [39, 1132, 223, 959, 113, 883, 133, 64, 395, 226, 449, 115, 614, 693, 400, 774, 310, 356, 1049, 803, 1227, 409, 712, 1201]. Ordered [1011]. Ordering [371, 689, 465, 478, 1046, 903, 924]. Orders [1183]. Ordinary [803]. Orear [948, 1252]. Oriented [1141]. Origin [874, 620]. Ornstein [970]. Orthochronous [657]. Orthogonal [250, 710, 1030, 786, 1267]. Orthogonality [515]. Orthogonally [1156]. Orthonormal [1001]. Oscillating [330]. Oscillations [26, 778]. Oscillator [1387, 1318, 1177, 1377, 316, 938]. Oscillators [240, 65, 553, 25, 1111, 231]. Other [909, 1166, 524]. Outer [726]. Overcomplete [1353, 1433]. Overlap [681, 335, 1090].

Packet [447]. Packets [608, 490, 73]. Padé [269, 758, 866, 610, 1238, 1367]. Pair [341, 953, 880, 302, 1326]. Pair-Correlation [1326]. Pairing [128]. Pairing-Force [128]. Pairs [652]. Pais [29]. paper [1436, 1435]. Para [64]. Para-Bose [64]. Para-Fermi [64]. Parallel [146]. Paramagnetic [516]. Parameter [492, 1009, 676]. Parameters [874, 506, 383]. Parasiuk [698]. Parentage [1188]. Parity [503]. Part [583, 865]. Partial [500, 828, 633, 15, 1233, 797, 1315, 599, 1292, 1197, 610]. Partial-Wave [633, 797, 1292, 1197, 610]. Partially [1154]. Particle [469, 1301, 1179, 391, 889, 67, 210, 1093, 928, 299, 106, 1364, 1115, 1397, 227, 652, 735, 155, 888, 1326, 1139, 202, 203, 537, 1318, 1413, 374, 617, 109, 960, 311, 783, 816, 727, 146, 870, 43, 861, 236, 1189, 1149, 942, 1172, 658]. Particle-Hole [1364]. Particlelike [734, 1034, 161, 519, 852, 989, 988]. Particles [951, 1381, 1204, 388, 631, 740, 508, 1283, 1316, 860, 1378, 1420, 351, 741, 263, 112, 158, 61, 380, 1412, 165, 976, 287, 288, 257, 1078, 1366, 1096, 1233, 1436]. Partition [268, 1344, 181, 1109]. partitioning [1408]. Parts [230, 164, 476, 459]. Passage [326, 1229]. Path [1424, 554, 982]. Path-Integral [982]. patterns [1163]. Pauli [1038, 114]. Peculiarities [645]. Penetrable [261]. Penetration [294, 969]. Penrose [818, 398, 1349]. Peratization [132, 531, 687, 119, 524]. Percolation [463]. Periodic [215, 1301, 799, 520, 1309, 1042]. Permeability [350, 781]. Permutation [357]. Permutation-Algebraic [357]. Permutational [675]. Perpendicular [1081]. Perturbation [341, 1391, 502, 1183, 932, 855, 82, 206, 225, 1356, 751, 1304, 912, 544, 173, 1176, 14, 252, 1412, 1302, 261, 1339, 1408, 925]. Perturbations [1258, 699, 1224, 1268]. Perturbed [907]. Petit [543]. Petrov [1145]. Pfaffian [9, 276]. Phase [1380, 990, 874, 635, 1132, 363, 754, 550, 383, 794, 1259, 228, 901, 277, 511, 246, 60, 202, 203, 1307, 902, 618, 777, 443, 805, 1426, 1330, 1164, 1125, 333, 780]. Phase-Integral [794, 1259]. Phase-Space [635, 363, 246, 60, 1164]. Phases [213]. Phenomena [970]. Phenomenological [663]. Phonon [414, 922]. Phonons [542]. Photon [898, 375, 38]. photosynthetic [1229]. Phys [1437]. Physical [1341, 1179, 275, 1139, 1413, 374, 749, 1193, 1202, 1203]. Physical-Region [1179]. Physics [469, 324, 457, 913, 676]. Picture [1265]. Pion [220]. Pion-Pion [220]. Planar [487, 24]. Planck [77, 648, 1177, 577]. Plane [909, 273, 1169, 1356, 830, 310, 282, 655, 146, 792, 665, 1383, 235, 1369, 1141]. Plane-Fronted [909]. Plane-Symmetric [1369]. Plane-Wave [1169]. Planes [688]. Plasma [75, 1245, 854, 953, 664, 899, 536, 1162, 778, 1147, 969, 576, 759, 878, 1117, 1309, 235, 1222]. Plasmas [1132, 1184, 345, 14, 45, 2, 175]. Platelets [428]. Plethysm [995]. Plus [578]. Poincaré [1237, 320, 388, 1025, 884, 644, 1211, 1398, 1121, 923, 1062, 958, 80, 1078]. Poincaré-group [958]. Point [926, 1293, 339, 684, 572, 82, 23, 244, 688, 782, 1248, 400, 1116, 1390, 1048, 1339, 1198]. Point-Coupling [1048]. Point-Loop [572]. Points [62]. Poisson [1120]. Polar [1311]. Polarizable [781]. Polarization [1315]. Polarized [862]. Pole [1094]. Poles [221, 420, 395, 56, 355, 191, 1402]. Polycrystalline [1247]. Polymer [714, 913]. Polynomial [868, 408, 142, 437, 941, 1267]. Polynomials [1319, 1341, 423, 1384, 250, 258]. Populations [1309]. Portions [749]. Position [12, 155]. Positive [388, 988, 749]. Positive- [749]. Positive-Definite [988]. Positronium [69]. Possess [28, 136]. Possessing [839]. Possibility [179]. Possible [1069, 1322, 1396, 377, 977]. Postulate [442, 1294]. Potential [138, 264, 99, 76, 28, 136, 990, 874, 222, 926, 842, 814, 815, 602, 1102, 1405, 531, 85, 321, 566, 1072, 948, 1252, 1209, 164, 838, 1071, 52, 280, 78, 104, 1310, 1275, 459, 168]. Potential-Correlation [814]. Potentials [587, 957, 224, 987, 927, 169, 387, 1217, 420, 831, 687, 187, 277, 1321, 545, 1173, 892, 496, 777, 1288, 522, 458, 1021, 524, 848, 1194, 459, 673, 1337, 1018]. Power [62, 1319]. Powers [218, 186, 1212]. Prediction [640]. Presence [1279, 1083]. Present [1293]. Preserving [131, 955]. Pressure [704]. Pressure-Free [704]. Primitive [877]. Principle [1327, 1265, 194, 851, 707, 1090, 1091, 137, 555, 1266, 729, 567, 165, 166, 287, 1131, 95, 1052]. Principles [578, 68, 603, 808, 1039, 661, 332]. Probability [714, 899, 385, 461]. Problem [293, 613, 784, 1126, 99, 697, 219, 743, 1414, 807, 983, 754, 1217, 1242, 716, 1185, 1356, 651, 9, 304, 276, 1107, 534, 936, 822, 1075, 804, 140, 306, 612, 778, 844, 1159, 59, 372, 373, 1173, 745, 618, 1368, 1309, 458, 667, 976, 1393, 1053, 1186, 1149, 962]. Problems [1120, 767, 821, 1104, 1119, 756, 169, 387, 116, 515, 1176, 991, 252, 785, 808, 1360, 1021, 1281, 1166, 1383, 1006, 1155, 1428, 1100, 1101]. Proca [500]. Proca-Type [500]. Procedure [275, 116]. Process [88]. Processes [463, 182, 1092, 1413, 374, 326, 820]. Produced [777]. Product [412, 503, 901, 1321, 370, 90, 200, 1044, 1426, 1286, 938]. Product-Integration [1321]. Production [66]. Products [678, 1011, 959, 489, 301, 1085, 823, 1071, 1312]. Projected [105]. Projection [217, 1231, 1040]. Projective [504, 162]. Prolate [936]. Prolate-Spheroidal [936]. Proof [66, 190, 583, 1294, 853, 1118, 183, 144]. Propagation [767, 1245, 1169, 530, 1058, 830, 665, 1166, 360]. Propagator [1114, 922]. Propagators [342, 907, 230]. Proper [26, 657, 355]. Properties [211, 796, 921, 807, 1240, 1367, 596, 344, 639, 987, 959, 831, 1018, 473, 929, 1430, 8, 244, 535, 1340, 952, 912, 582, 189, 1146, 1388, 352, 1247, 385, 692, 1218, 1353, 1433, 1142, 522, 458, 848, 812, 381, 1110, 1193]. Property [37, 1213, 78]. Propositions [1046]. Pseudo [1030]. Pseudo-Orthogonal [1030]. Pseudo-Unitary [1030]. Pseudolimits [1328]. Pseudonull [1067]. Pulses [1356].

Quadratic [843, 1340, 1266]. Quadrature [307]. Quadrupole [256]. Quantities [1205, 1333, 106, 32, 1349]. Quantization [219, 472, 1331, 936, 1317, 967]. Quantized [271]. Quantum [18, 34, 290, 291, 66, 950, 131, 12, 241, 156, 1132, 442, 271, 345, 953, 446, 169, 387, 48, 509, 699, 121, 243, 904, 36, 37, 187, 769, 188, 779, 809, 300, 58, 1342, 227, 228, 1064, 1127, 1216, 696, 735, 148, 534, 91, 1418, 1008, 154, 871, 956, 1386, 1111, 1318, 398, 283, 492, 111, 189, 896, 47, 402, 1195, 456, 763, 117, 129, 405, 406, 477, 525, 130, 443, 518, 63, 961, 1276, 1377, 1053, 982, 331, 897, 434, 1046, 381, 73, 676, 1196, 977, 942, 1336, 1155, 1428, 1073, 962, 313]. Quantum-Coulomb [1132]. Quantum-Ergodic [1053]. Quantum-Mechanical [131, 91, 398, 518, 1276, 942, 962]. Quantum-Statistical [1418, 1008, 982]. Quantumelectrodynamics [230].

Quasi [553, 652]. Quasi-Classical [553]. Quasi-Spin [652]. Quasibinomial [1028, 1029]. Quasiparticle [1409, 1239, 1091]. Quasistatic [1115]. Quaternian [309]. Quaternion [53]. Quaternionic [622].

r [264, 1422]. Racah [205]. Radial [796, 389, 81, 771, 660]. Radiating [997]. Radiation [731, 593, 1279, 241, 875, 510, 893, 931, 512, 688, 1074, 160, 862, 719, 259, 317]. Radiationless [807]. Radiative [523, 1088]. Radius [502, 714]. Rainbow [1158]. Raise [87]. Raising [710, 786]. Random [35, 1063, 463, 1050, 1026, 350, 830, 1007, 30, 1229, 620]. Randomly [1141]. Range [1269, 420, 1396, 1395, 133, 1316, 1303, 1382, 164, 1218, 827, 1137, 459, 944]. Rank [96]. Rates [611]. Ray [294, 503, 592, 1187, 1228, 376, 51, 428, 604, 1141]. Rays [334]. Reactions [827]. Reactor [722]. Real [932, 53, 649, 1304, 309, 207]. Real-Time [932, 1304]. Realization [802, 1208, 958]. Realizations [691, 312, 833, 1012]. Rearrangements [877]. Reciprocity [1327, 166, 287, 1189]. Recoupling [322, 558]. Rectangular [939, 9, 1100]. Recurrences [532]. Recursion [199, 768]. Recursive [694]. Reduced [36, 37, 187, 721, 1171, 598, 1376, 804, 675, 1351, 135, 695]. Reduced-Width [135, 695]. Reducibility [1174]. Reducible [657, 492, 1121]. Reducing [1085]. Reduction [29, 762, 984, 1355, 1379, 176, 177, 301, 1055, 1130, 657, 701, 1121, 865, 911, 200, 254]. reductions [1254]. Refinements [144]. Reflection [832, 1157, 1117]. Refractive [781]. Regge [633, 221, 532, 1094, 212, 191, 856]. Regge-Pole [1094]. Reggeon [255]. Region [1179, 338, 209, 275, 749, 1193, 1202, 1203]. Regions [269, 1413, 374]. Regular [834, 823]. Regularization [698, 338, 687, 897]. Regularized [572]. Rejuvenates [330]. Related [504, 1348, 548, 1184, 864, 351, 201]. Relating [179]. Relation [240, 633, 726, 628, 9, 276, 1340, 793, 772, 624]. Relations [1380, 578, 131, 199, 720, 583, 768, 715, 1342, 370, 654, 515, 404, 479]. Relationship [542, 615]. Relationships [908, 1164]. Relativistic [18, 34, 290, 291, 501, 1003, 828, 697, 753, 295, 807, 506, 210, 417, 839, 599, 509, 106, 85, 474, 322, 192, 717, 721, 875, 1092, 628, 770, 573, 323, 1378, 1420, 155, 888, 778, 1147, 1116, 158, 701, 865, 1117, 1049, 61, 1277, 539, 1048, 1078, 1366]. Relativity [1325, 219, 19, 1205, 416, 1082, 736, 392, 748, 704, 837, 242, 646, 915, 1086, 986, 1213, 997, 706, 1266, 895, 284, 996, 1346]. Relaxation [576]. Relevant [113]. Remark [145, 1307, 492]. Remarkable [1016]. Remarks [717, 721, 228, 1371, 1372, 888, 1010, 499, 775]. Renormalizability [342]. Renormalizable [265, 266]. Renormalization [1335, 68, 698, 920, 572, 1419, 1214, 1045, 410]. Renormalized [716]. Representability [1376, 804, 1223]. Representation [951, 1427, 504, 633, 594, 388, 1108, 631, 48, 715, 348, 1345, 1370, 1273, 823, 1289, 17, 1130, 575, 279, 1385, 1308, 93, 791, 212, 217, 1390, 565, 1006, 1237, 1254]. Representations [70, 1029, 886, 239, 504, 503, 592, 965, 441, 762, 1112, 1129, 638, 998, 906, 1241, 600, 984, 1355, 1379, 883, 1243, 394, 145, 647, 489, 176, 177, 301, 1228, 1019, 17, 91, 370, 1059, 1340, 90, 761, 730, 514, 1, 5, 83, 84, 657, 1248, 1398, 746, 493, 494, 556, 1174, 701, 865, 354, 825, 869, 911, 1246, 376, 622, 789, 200, 134, 497, 727, 44, 427, 163, 162, 933, 711, 666, 857, 1135, 196, 236, 539, 624, 1020, 786, 1267, 790, 1070, 362, 1028, 1148, 981, 1211, 938, 1121, 923, 1062, 1015, 1354, 1342]. Representing [1220]. Repulsion [1316]. Repulsive [1217, 788, 777, 1274, 387, 52]. Requirement [854]. Requirements [1061, 795]. Residual [454]. Resistive [625]. Resolvent [840]. Resolvents [473]. Resonance [1098, 765]. Resonances [1060]. Resonant [316]. respect [176, 177]. Respective [1129]. Response [542, 1115]. Rest [320, 388, 1150]. Restricted [1257, 1398, 789, 711]. Restriction [1168, 816]. result [1154]. Resulting [338]. Results [855, 1332, 800, 1026]. Retarded [1079]. Reversal [1284, 1118]. Rideau [555]. Riemann [1213, 1230]. Riemannian [864, 430, 431]. Right [1299]. Right-Angle [1299]. Rigid [382]. Rigorous [1332, 278, 1026]. Ring [1322, 345, 142, 437, 256]. Ring-Current [256]. Riseman [1323]. Robertson [1277]. Robinson [571]. Rod [880]. Rods [49, 428]. Role [442, 566]. Root [564]. Rosenbaum [1401, 1153]. Rotating [733, 110, 561, 20]. Rotation [70, 993, 1112, 369, 1019, 514, 134, 833, 497, 258, 1020]. Rotational [22, 172, 1002]. Rotationally [216]. Rotationally-Symmetric [216]. Rotations [250]. Roussopoulos [555]. Ruelle [760]. Rule [551, 1118]. Rules [1098, 413, 1083, 544, 607, 857, 198, 1103]. Runge [626]. Rys [455].

S [842, 420, 16, 452, 841, 379]. Sachs [378]. Salam [1411]. Salpeter [1403, 697, 1161, 535, 1055, 1209, 355, 476, 380, 849]. Sample [32]. Saturated [95]. Saturation [1098]. Scalar [440, 549, 914, 983, 1067, 997, 17, 879, 1044, 22, 172, 126, 328, 836, 1262, 161, 989, 988, 331]. Scalar-Tensor [983, 997]. Scalars [900, 867]. Scattered [433]. Scattering [1380, 66, 138, 264, 588, 1060, 1381, 813, 975, 99, 76, 954, 220, 1240, 1361, 914, 1296, 1179, 273, 583, 367, 987, 927, 584, 50, 855, 815, 599, 275, 447, 214, 1424, 1292, 321, 566, 608, 97, 204, 1107, 534, 1072, 948, 1252, 490, 229, 1382, 157, 1097, 879, 1071, 739, 251, 372, 352, 517, 892, 238, 878, 745, 1157, 476, 662, 1360, 61, 1218, 212, 458, 667, 1021, 51, 428, 604, 147, 1194, 760, 812, 719, 623, 1222, 673, 73, 949, 728, 1141, 1113, 934, 1158]. Scheme [1334, 652, 1224, 232, 457]. Schemes [57]. Schrödinger [367, 194, 303, 545, 134, 358, 1233, 1408, 934, 173, 978]. Schwarzschild [1033, 1324, 879]. Schwinger [732, 1328, 673]. Screened [346, 496]. Screening [934]. Second [96, 223, 1064, 115, 554, 693, 400, 1049, 712, 1201]. Second-Order [223, 115, 693, 400, 1049, 712, 1201]. Second-Rank [96]. Secondary [1382]. Section [118, 490, 1122, 51, 1194, 1437]. Sections [1000]. Sector [311, 891, 985]. Secular [267]. Selection [413, 857]. Self [10, 54, 1357, 1416]. Self-Consistent [10, 54, 1357, 1416]. Semi [684, 451, 140, 773]. Semi-Infinite [684, 451, 140, 773]. Semiclassical [896].

Semiconductors [516]. semidirect [938]. semidirect-product [938]. Semigroups [722]. semimaximal [1163]. Semisimple [239, 394, 1124, 499]. Seniority [652]. Separability [936]. Separable [1173]. Separation [218, 164, 459, 1233]. Sequences [269]. Series [678, 62, 502, 750, 680, 267, 4, 262, 866, 971, 1108, 569, 575, 911, 991, 403, 824, 1124, 460, 674, 1201, 1238, 1339, 1198, 1399, 1400, 1109]. Set [86, 223, 659]. Several [56, 1303, 1058]. Shape [183]. Shaped [261]. Shapes [540]. Shell [975, 1409, 101, 102, 103, 1221, 662, 46, 938, 1133]. Shells [1234]. Shift [1325, 990, 874, 277]. Shifts [383, 777, 780]. Shocks [195]. Short [1396, 1395, 1316, 164, 401, 621]. Short-Range [1396, 1395, 1316, 164]. Showing [247]. sided [905]. Sign [853]. Signals [39]. Simple [682, 1191, 529, 1099, 1269, 49, 1387, 164, 615, 493, 494, 1041, 401, 607, 621, 630, 198, 410, 978, 1203, 282]. Simple-Cubic [282]. Single [1280, 842, 1326, 861]. Single-Particle [1326, 861]. Singular [138, 264, 587, 957, 526, 527, 224, 987, 531, 687, 422, 1356, 277, 305, 252, 777, 913, 524, 848, 1194, 483, 1198]. Singularities [1296, 100, 915, 145, 227, 452, 1248, 1094, 1339, 1203]. Singularity [1324, 467, 879, 538, 1193, 1202, 1203]. Sinusoidally [1245]. Sitter [1172, 569, 426]. Six [1413, 882]. Six-Dimensional [882]. Six-Particle [1413]. Sixth [113]. Sixth-Order [113]. Size [32]. Skew [152]. Skew-Symmetric [152]. Slab [813, 954, 751, 372, 373, 433, 905]. Slater [455, 1239]. Slit [1374]. Slow [1420]. Small [1284, 1309, 51, 428, 604, 1141]. Small-Amplitude [1309]. Small-Angle [51, 428, 1141]. Smooth [171, 709]. Soft [898]. Soft-Photon [898]. Solid [994, 496]. Solids [681, 903]. Soluble [640, 528, 1375, 247, 353]. Solution [1084, 193, 1088, 954, 440, 549, 218, 1414, 807, 267, 758, 1242, 487, 507, 422, 534, 305, 23, 229, 1209, 1321, 1384, 819, 844, 399, 42, 1000, 1167, 114, 327, 146, 358, 913, 147, 1227, 337, 316, 20, 1113, 905]. Solutions [1120, 468, 835, 197, 527, 937, 1184, 1297, 734, 766, 1363, 471, 141, 571, 1054, 77, 1283, 1087, 751, 1253, 97, 1171, 451, 1139, 371, 616, 310, 284, 996, 253, 160, 159, 72, 1150, 1034, 1360, 161, 519, 520, 852, 989, 988, 1309, 1281, 801, 1002, 483, 1277, 1036, 1244]. Solvable [222, 1035]. Solving [9, 276, 515]. Some [55, 921, 1348, 762, 416, 754, 839, 113, 394, 3, 436, 718, 1127, 1216, 97, 323, 952, 281, 946, 284, 45, 385, 1039, 521, 1218, 694, 499, 196, 775, 968, 1160, 411, 239, 1026]. Source [578, 975, 684]. Sources [108, 575, 233]. Space [18, 34, 290, 291, 588, 439, 413, 1280, 635, 1156, 363, 343, 742, 637, 909, 224, 270, 272, 271, 641, 472, 700, 645, 1086, 647, 348, 794, 1259, 98, 228, 686, 303, 1221, 244, 41, 246, 1213, 1059, 60, 202, 203, 887, 231, 895, 746, 1000, 251, 515, 1208, 1013, 1014, 1051, 87, 1421, 254, 525, 443, 1066, 793, 1276, 163, 849, 287, 288, 862, 170, 1164, 257, 668, 565, 1006, 968, 1149]. Space-Momentum [793]. Spacelike [698, 1234]. Spaces [1407, 19, 242, 16, 864, 1074, 87, 430, 431, 1230, 382]. Spacing [1007]. Spatially [750, 1301, 1093, 107, 918, 323]. Special [832, 742, 748, 41, 324, 283, 895, 1013, 1014, 1051, 71, 539]. Specialized [659]. Spectra [679, 1386, 995, 944]. Spectral [76, 1182, 1423, 1166, 1390].

Spectroscopic [811]. Spectroscopy [1004]. Spectrum [1191, 379, 1390]. Speed [873, 140, 654, 310, 1100, 1101]. Sphere [339, 554, 1000, 1157, 1158]. Spheres [997]. Spherical [731, 975, 873, 369, 817, 450, 1278, 396, 1307, 894, 771, 999, 539, 237]. Spherically [7, 1227]. Spheroidal [936, 411]. Spin [11, 12, 963, 1089, 817, 488, 552, 1332, 966, 650, 652, 5, 84, 1159, 143, 263, 112, 158, 357, 1095, 1021, 165, 976, 1231, 1065, 826, 459, 409, 1040, 563, 1125, 289, 482, 890, 258, 1078, 1366, 1096, 1026]. Spin- [817]. Spin-Free [357]. Spin-Matrix [258]. Spin-Orbit [1021, 1040, 1095]. Spin-Weighted [826]. Spinless [1381, 979, 1073]. Spinning [1235, 109]. Spinning-Particle [109]. Spinor [506, 1086, 649, 686, 1140, 72, 134, 1370]. Spinors [594, 744, 481]. Splitting [700, 315]. Spontaneous [568, 143]. Square [451, 1209, 857, 1028]. Square-Well [1209]. Stability [75, 595, 299, 664, 947, 561, 104, 1270]. Stage [89]. standard [1121]. Star [464, 150, 151]. Stark [507]. Stark-Effect [507]. State [697, 874, 1415, 1204, 222, 906, 740, 1217, 1364, 77, 91, 116, 1024, 1308, 311, 692, 1190, 1310, 1339, 1163, 1338]. States [549, 28, 136, 67, 210, 107, 446, 383, 13, 322, 794, 1259, 1373, 901, 512, 105, 898, 1418, 730, 1008, 537, 475, 1056, 1057, 353, 1290, 232, 354, 1239, 783, 1251, 961, 757, 1275, 760, 1189, 1196, 1163, 1338]. Static [440, 549, 640, 1297, 1077, 1175, 97, 1139, 1033]. Static-Model [1297]. Stationary [1082, 1253, 399, 692, 961, 1033, 20]. Statistical [950, 149, 589, 1220, 1063, 914, 814, 345, 446, 980, 843, 65, 243, 53, 590, 940, 1004, 188, 717, 721, 755, 543, 574, 696, 246, 1418, 1008, 729, 1247, 402, 991, 457, 764, 757, 1426, 27, 359, 1131, 459, 982, 234, 381, 1219]. Statistics [1089, 121, 904, 858, 47, 454, 455, 1065]. Steady [77, 759]. Steady-State [77]. Stieltjes [1238, 866, 713, 1037, 1160]. Stimulated [241]. Stochastic [1092, 1336]. Stokes [506]. Störmer [256]. Stratified [1245]. Strength [838, 1310, 1275]. Stress [732]. Stress-Tensor [732]. Stretched [1422]. Strip [1374, 238, 1031]. Strong [638, 602, 1102, 1405, 1420]. Strong-Coupling [1102, 1405]. Strongly [508, 277]. Structure [18, 34, 290, 291, 149, 501, 1003, 319, 186, 766, 368, 570, 700, 884, 1086, 108, 230, 615, 922, 910, 538, 824, 1406, 481, 1193, 1202, 1203, 1163, 1338]. Structures [963]. Studies [173]. Study [116, 1426, 1192]. Sturm [1341, 723]. Sturmian [962]. Subenergy [466]. Subgroup [176, 177, 1398, 1041, 911, 711]. Subgroups [1075]. Subject [1432, 1167]. Submerged [157]. Substitution [832, 1422]. Subtractions [99]. Successive [853]. Sudarshan [1024]. Sufficiency [1225]. Sufficient [28, 136, 1270]. Sum [1098, 544, 282, 655, 1103]. Summation [15, 79, 1311]. Sums [1315, 1426]. Super [1276]. Superconducting [185, 1031]. Superconductivity [1394, 747]. Superconductors [728]. Superfluid [245]. Supermultiplet [1334]. Superposition [1057]. Superselection [1083, 1118]. Supplementary [911]. Surface [914, 183, 157, 559, 560, 1033]. Surface-Energy [559, 560]. Surfaces [682, 297, 390, 486, 709, 749]. Susceptibilities [1081]. Susceptibility [1335]. Switch [195]. Switch-on [195]. Switching [185, 367]. Symanzik [285]. symbol [1028]. Symbolic [88]. Symbols [344, 639, 908]. Symmetric [835, 152, 7, 216, 1074, 354, 212, 1227, 1369, 1006, 1201, 1172]. Symmetrical [178]. Symmetries [697, 508, 1364, 179, 864, 377]. Symmetrization [1294]. Symmetrized [678, 863]. Symmetry [292, 546, 832, 1134, 1170, 640, 344, 639, 641, 368, 570, 597, 700, 106, 57, 1257, 98, 1213, 25, 1165, 120, 1278, 283, 492, 279, 1208, 312, 894, 675, 315, 166, 1002, 170, 257, 482, 80, 1155, 1428, 1428]. Symplectic [705, 1015]. sympletic [1354]. System [500, 1415, 542, 740, 48, 1115, 1282, 735, 49, 1224, 1055, 491, 1359, 42, 741, 1284, 22, 172, 1263, 168, 861, 1274, 1339, 1110]. Systems [951, 1134, 613, 784, 1126, 1063, 1204, 1032, 215, 107, 980, 642, 1260, 299, 809, 543, 8, 860, 864, 1326, 164, 799, 689, 47, 232, 617, 960, 816, 54, 1357, 747, 1412, 803, 1416, 459, 462, 142, 437, 236, 316, 1042, 1337, 1436, 1105, 1026, 1090, 1091].

Tabulation [90, 200]. Tail [917]. Talmi [25, 351]. Tamm [985]. Taub [1000]. TCP [1066]. Technique [1023, 72, 1368, 919, 1408]. Techniques [1140, 895]. Telegraph [616]. Temperateness [583]. Temperature [345, 551, 1425, 143, 40]. Temperature-Dependent [551]. Temperatures [1064, 511]. Tensor [678, 1295, 96, 632, 634, 732, 983, 1213, 997, 830, 895, 93, 1106, 894, 386, 863]. Tensorial [125]. Tensors [152, 594, 242, 450, 559, 560, 737]. Terms [732, 990, 874, 1112, 153, 1, 5, 83, 84, 1197, 1328, 1263]. Tetrad [637, 837]. th [1319]. Their [808, 1309, 624, 416, 1254, 496, 710]. Theorem [1098, 698, 296, 366, 118, 182, 943, 1178, 705, 864, 598, 1376, 491, 648, 398, 992, 403, 184, 933, 524, 334, 668, 460, 1048]. Theorems [541, 754, 281, 1195, 496, 847, 478, 359, 1125, 924]. Theoretic [950, 1205, 447]. Theoretical [828, 202, 1251, 1192, 410]. Theories [265, 266, 753, 1011, 889, 50, 383, 572, 214, 300, 101, 102, 103, 148, 216, 117, 161, 519, 520, 619, 852, 989, 961, 988, 1046]. Theory [18, 34, 290, 291, 66, 957, 1427, 682, 96, 921, 12, 677, 1250, 756, 937, 581, 442, 542, 876, 983, 1183, 297, 390, 935, 758, 1409, 270, 271, 1182, 1404, 754, 367, 580, 1169, 1241, 643, 1333, 855, 602, 1102, 1405, 132, 1243, 1152, 298, 1424, 85, 321, 566, 769, 547, 347, 82, 192, 422, 206, 510, 893, 1228, 722, 751, 227, 1127, 1216, 709, 204, 246, 653, 997, 155, 1289, 17, 1326, 555, 912, 325, 544, 467, 164, 278, 1139, 1318, 1429, 398, 283, 492, 231, 173, 111, 516, 1176, 1313, 326, 93, 14, 456]. Theory [763, 618, 1421, 167, 1034, 663, 525, 92, 313, 1352, 443, 78, 791, 63, 426, 1181, 1412, 104, 285, 287, 288, 665, 1166, 995, 826, 459, 577, 191, 856, 462, 331, 332, 480, 289, 897, 434, 672, 669, 625, 671, 877, 73, 1141, 260, 1196, 1336, 1172, 1219, 1096, 979, 1105, 1408, 1073, 934, 1345, 1370, 1100, 1101, 925, 1158]. Thermal [1169, 1260, 226, 225, 448, 449, 1210, 425, 1388, 621]. Thermalization [1362]. Thermodynamic [243, 685, 1110]. Thermodynamically [590]. Thermodynamics [1184, 27, 1274]. Thin [881, 1031]. Third [39, 365]. Third- [39]. Those [972]. Three [218, 1414, 1112, 67, 210, 169, 387, 1185, 651, 1107, 244, 1171, 178, 203, 495, 690, 960, 311, 134, 783, 1006, 1337, 46, 712, 962, 1100, 1101]. Three-Body [1414, 169, 387, 651, 1107, 1006, 1337, 46, 712, 962]. Three-Dimensional [1112, 1185, 1171, 495, 690, 134, 1100, 1101]. Three-Particle [67, 210, 203, 960, 311, 783]. Threshold [221, 1061]. Threshold-Behavior [1061]. Thresholds [466]. Time [341, 593, 859, 1206, 1011, 529, 932, 547, 347, 194, 206, 722, 755, 1210, 423, 1304, 140, 1111, 1318, 1284, 1328, 385, 130, 1276, 104, 358, 1377, 627, 588, 439, 1156, 637, 270, 272, 271, 641, 700, 1118, 686, 303, 41, 1066, 287, 288, 170, 668, 1113]. Time-Dependent [341, 593, 859, 1206, 529, 194, 206, 1210, 140, 1111, 1318, 1377, 627, 1113]. Time-Ordered [1011]. Time-Reversal [1284, 1118]. Timelike [399, 812]. times [343, 742, 909, 1086, 1213, 1229, 968]. Topologies [588]. Topology [646, 411]. Torsion [472]. Total [631, 968]. Trace [154, 863]. Traces [946]. Trajectories [864]. Trajectory [856]. Transfer [294, 1395, 118, 182, 844, 827, 523, 1088]. Transfer-Matrix [294]. Transform [956, 307, 1144]. Transformation [500, 152, 752, 926, 58, 25, 1081, 351, 871, 1016, 263, 112, 158, 659, 429]. Transformations [1327, 528, 339, 106, 818, 943, 1199, 1373, 1386, 1392, 1399, 1400]. Transforms [1358, 680, 713, 1037, 1160]. Transition [23, 278, 1307, 902, 759, 357, 805, 1330, 1125, 333]. Transitions [1132, 550, 511]. Transitive [1156]. Translation [94, 775]. Translational [596, 582, 1146]. Translations [286]. Transmission [905, 975, 1157, 1117, 1142]. Transparent [171, 914, 1157, 1158]. Transport [1084, 954, 684, 873, 751, 1171, 555, 654, 251, 515, 310, 326, 32, 1421, 808, 1113, 1244, 1100, 1101]. Transverse [1295, 1147]. Transversity [1094]. trapping [1229]. Trautman [571]. Treating [21, 164, 252]. Treatment [828, 1408, 278, 202, 453]. Tree [1180]. Trial [1054]. Triangle [513]. Triangular [1287, 1329, 409]. Trilinear [479]. Turbulence [1181, 89]. Turning [400]. Twistor [586]. Two [62, 123, 440, 549, 1119, 1011, 1361, 1401, 1170, 343, 1099, 684, 50, 1153, 955, 608, 41, 788, 1382, 1138, 164, 1248, 1159, 1071, 1027, 960, 336, 61, 747, 1309, 27, 976, 661, 919, 1312, 459, 168, 181, 333, 1390, 235, 38, 905, 1233, 1038, 1100, 1101, 543, 236]. Two- [960, 1100, 1101]. Two-Body [608, 788, 1382, 61, 976]. Two-Center [1027, 336, 919, 1038]. Two-Dimensional [123, 1119, 1011, 1170, 1099, 955, 41, 181, 333]. Two-Meson [440, 549]. Two-Nucleon [168]. Two-Particle [236]. Two-Photon [38]. Two-Point [1248, 1390]. two-sided [905]. Two-Variable [1361]. Type [500, 293, 578, 1323, 528, 16, 1285, 283, 996, 1347, 71]. Type-I [71]. Type-Null [996].

U [634, 636, 600, 257]. Uhlenbeck [298]. Ultraviolet [897]. Unbound [353]. Unbounded [318]. Uncertainty [901]. Uncoupling [1383]. Underdense [1222]. Undirected [465]. Unequal [351]. Unequal-Mass [351]. Uniaxially [1279]. Unified [677, 388, 453, 111, 93]. Uniform [1295, 928, 1429, 2, 175, 73, 664]. Uniformly [1279, 688]. Unifying [729]. Unimodular [281]. Uniqueness [756, 937, 77, 183, 311, 760]. Unitarity [211, 220, 907, 797, 1077, 1175, 145, 101, 102, 103, 1139, 377, 1122, 255, 1437, 1202]. Unitary [292, 546, 832, 886, 503, 199, 388, 1148, 25, 1211, 761, 493, 494, 867, 868, 825, 869, 911, 1062, 727, 44, 6, 174, 427, 1030, 711, 166, 482, 1402, 196, 236, 790, 1070, 362]. units [1229]. Universal [569, 397]. Universe [319, 931, 308, 1347, 562, 20]. Universes [1079, 585, 1288]. Upper [76, 393, 173, 798, 723, 611]. Use [39, 962]. Using [1427, 1384, 1146]. Usual [588].

V [239, 17]. Vacancy [1417]. Vacuum [416, 364, 1315, 1253, 1285, 283, 996, 1239]. Valid [621]. Validity [502, 528, 1225]. Value [1120, 821, 1104, 1119, 756, 612, 778, 372, 373, 1176, 785, 1281]. Valued [1058]. Values [318, 874, 154, 385, 381]. Vanishing [156, 484, 1213]. Vapor [278]. Variable [1361, 709, 155, 1046]. Variables [48, 509, 21, 1058, 466, 1131]. Variation [485]. Variational [1120, 68, 990, 603, 194, 598, 137, 555, 1097, 232, 808, 567, 1039, 661, 1131, 332, 95]. Varies [773]. Various [550, 173]. Vector [1060, 845, 626, 895, 87, 362, 1215]. Vector-Meson [1060]. Vectors [1353, 1433, 564, 1163, 1338]. Velocity [1187, 831, 474, 853, 350, 1018]. Velocity-Dependent [831, 1018]. Versa [1005]. Vertex [614, 355, 79, 78]. Vertices [254]. very [422]. Vice [1005]. vicinity [1198]. View [787]. Viewpoint [324]. Violation [1046]. Virial [1064, 554, 359]. Vlasov [854, 875, 617, 45, 104]. Volterra [1309]. Volume [1115, 36, 1431, 1432]. Vries [1016, 1017, 1200].

Waals [278, 551, 398]. Walk [620]. Walker [1277]. Walks [30, 1229]. Wall [359]. Waller [401, 621, 670]. Ward [453]. Wave [1391, 633, 990, 874, 1245, 1061, 917, 366, 797, 1169, 599, 447, 1292, 1283, 608, 490, 1058, 424, 1114, 1197, 610, 143, 238, 1049, 146, 443, 161, 827, 665, 1166, 1302, 73]. Wave-Packet [447]. Wavefield [1278]. Wavefunction [848]. Wavefunctions [1407, 222, 303, 701, 865, 712]. Waveguides [1186]. Waves [859, 1187, 15, 1128, 914, 909, 529, 1099, 474, 1356, 350, 137, 1299, 1147, 157, 830, 879, 738, 739, 759, 282, 655, 1117, 772, 623, 781, 1222, 261, 905]. Weak [927, 1316, 851, 991, 944]. Weak-Graph [991]. Weakly [542, 785]. Weber [612]. Wedge [1299]. Weight [123, 642, 1024]. Weighted [603, 826]. Weights [591]. Weil [244]. Well [1209]. Weyl [871, 956, 703, 274]. Which [591, 1002, 250, 330]. Whose [773, 1284]. Wick [423]. Width [1284, 135, 695]. Wiener [905, 1358, 88, 89, 1383]. Wiggles [551]. Wightman [798]. Wigner [239, 636, 199, 1103, 683, 1334, 696, 908]. Wind [1180]. Wind-Tree [1180]. Without [129, 1118, 534, 399, 673]. WKB [767, 885, 1225]. World [364, 742, 918, 207, 426]. Wulff [183].

X [51, 428, 604, 1141]. X-Ray [51, 428, 604, 1141].

Yang [810]. Yukawa [892].

Z [587]. Zak [1059]. Zernike [970]. Zero [502, 1237, 388, 1055, 1307, 5, 84, 263, 112, 923, 1150, 972, 865]. Zero-Mass [1055, 1237, 923, 865]. Zero-Rest-Mass [1150]. Zeros [1348, 268]. Zimmermann [285]. Zone [282, 655].

References

Lomont:1964:RILa

[1] J. S. Lomont and H. E. Moses. The representations of the inhomogeneous Lorentz group in terms of an angular momentum basis. Journal of Mathematical Physics, 5(2):294–298, February 1964. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i2/p294_s1. See errata

[83].

Silven:1964:AOM

[2] Saul Silven. Application of operational methods to the analysis of uniform plasmas. Journal of Mathematical Physics, 5(4):557–560, April 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v5/i4/p557_s1. See errata [175].

Glasser:1964:NES

[3] M. L. Glasser. Note on the evaluation of some Fermi integrals. Journal of Mathematical Physics, 5(8):1150–1152, August 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i8/p1150_s1.

Coleman:1964:CGS

[4] Sidney Coleman. The Clebsch–Gordan series for SU(3). Journal of Mathematical Physics, 5(9):1343–1344, September 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v5/i9/p1343_s1. See erratum [262].

Lomont:1964:RILb

[5] J. S. Lomont and H. E. Moses. Representations of the inhomogeneous Lorentz group in terms of an angular momentum basis: Derivation for the cases of nonzero mass and zero mass, discrete spin. Journal of Mathematical Physics, 5(10):1438–1457, October 1964. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v5/i10/p1438_s1. See [84].

Roos:1964:CUM

[6] Matts Roos. On the construction of a unitary matrix with elements of given moduli. Journal of Mathematical Physics, 5(11):1609–1611, November 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v5/i11/p1609_s1. See erratum [174].

Bergmann:1965:SSG

[7] P. G. Bergmann, M. Cahen, and A. B. Komar. Spherically symmetric gravitational fields. Journal of Mathematical Physics, 6(1):1–5, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i1/p1_s1.

Hunziker:1965:CPM

[8] W. Hunziker. Cluster properties of multiparticle systems. Journal of Mathematical Physics, 6(1):6–10, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i1/p6_s1.

Hurst:1965:RBO

[9] C. A. Hurst. Relation between the Onsager and Pfaffian methods for solving the Ising problem. I. The rectangular lattice. Journal of Mathematical Physics, 6(1):11–18, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i1/p11_s1.

Fraga:1965:ISC

[10] Serafin Fraga. Improved self-consistent configuration interaction. Journal of Mathematical Physics, 6(1):18–23, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i1/p18_s1.

Berg:1965:CHS

[11] R. A. Berg. Covariant higher-spin equations. Journal of Mathematical Physics, 6(1):24–33, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p24_s1.

Berg:1965:PIS

[12] R. A. Berg. Position and intrinsic spin operators in quantum theory. Journal of Mathematical Physics, 6(1):34–39, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i1/p34_s1.

Ghirardi:1965:NBS

[13] Gian Carlo Ghirardi and Alberto Rimini. On the number of bound states of a given interaction. Journal of Mathematical Physics, 6(1): 40–44, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i1/p40_s1.

Nagelberg:1965:PTM

[14] Elliott R. Nagelberg. Perturbation theory of microwave interaction with gyroelectric plasmas. Journal of Mathematical Physics, 6(1):44–50, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i1/p44_s1.

Brysk:1965:SPW

[15] Henry Brysk. Summation of partial waves. Journal of Mathematical Physics, 6(1):51–54, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p51_s1.

Grossmann:1965:HST

[16] Alexander Grossmann. Hilbert spaces of type S. Journal of Mathematical Physics, 6(1):54–67, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p54_s1.

Klauder:1965:CRT

[17] John R. Klauder and James McKenna. Continuous-representation theory. V. Construction of a class of scalar boson field continuous representations. Journal of Mathematical Physics, 6(1):68–87, January 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i1/p68_s1.

Ahmavaara:1965:SSFa

[18] Yrjö Ahmavaara. The structure of space and the formalism of relativistic quantum theory. I. Journal of Mathematical Physics, 6(1):87– 93, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i1/p87_s1.

Brans:1965:IAG

[19] Carl H. Brans. Invariant approach to the geometry of spaces in general relativity. Journal of Mathematical Physics, 6(1):94–102, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i1/p94_s1.

Wright:1965:SEF

[20] James P. Wright. Solution of Einstein’s field equations for a rotating, stationary, and dust-filled universe. Journal of Mathematical Physics, 6(1):103–105, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p103_s1.

Jones:1965:MTF

[21] Gerald L. Jones. Method for treating fluctuation of dynamical variables. Journal of Mathematical Physics, 6(1):106–111, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i1/p106_s1.

Pease:1965:SIRa

[22] M. C. Pease. Scalar invariants of a rotational system in a Lie algebra. Journal of Mathematical Physics, 6(1):111–117, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i1/p111_s1.

Jordan:1965:FSN

[23] P. F. Jordan and P. E. Shelley. Formal solution of a nonhomogeneous differential equation with a double transition point. Journal of Mathematical Physics, 6(1):118–135, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i1/p118_s1.

Martinis:1965:CHE

[24] M. Martinis. Complete high-energy behavior for certain planar graphs. Journal of Mathematical Physics, 6(1):136–141, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i1/p136_s1.

Kaufman:1965:USO

[25] Bruria Kaufman and Cornelius Noack. Unitary symmetry of oscillators and the Talmi transformation. Journal of Mathematical Physics, 6(1):142–152, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p142_s1.

Legendy:1965:EPM

[26] C. R. Legéndy. Existence of proper modes of helicon oscillations. Journal of Mathematical Physics, 6(1):153–157, January 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i1/p153_s1.

Salzberg:1965:EST

[27] Allan M. Salzberg. Exact statistical thermodynamics of gravitational interactions in one and two dimensions. Journal of Mathematical Physics, 6(1):158–160, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i1/p158_s1.

Calogero:1965:SCAa

[28] Francesco Calogero. Sufficient conditions for an attractive potential to possess bound states. Journal of Mathematical Physics, 6(1):161– 164, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i1/p161_s1.

Bollini:1965:RFF

[29] C. G. Bollini, J. J. Giambiagi, and A. González Domínguez. On the reduction formula of Feinberg and pais. Journal of Mathematical Physics, 6(1):165–166, January 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v6/i1/p165_s1.

Montroll:1965:RWL

[30] Elliott W. Montroll and George H. Weiss. Random walks on lattices. II. Journal of Mathematical Physics, 6(2):167–181, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i2/p167_s1.

Zimmerman:1965:ACF

[31] Robert L. Zimmerman. Analytic continuation of a (l, s) for ℜl < N. Journal of Mathematical Physics, 6(2):182–192, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i2/p182_s1.

Nossal:1965:EFS

[32] Ralph Nossal. Effects of finite sample size on the measurement of transport quantities. Journal of Mathematical Physics, 6(2):193–200, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i2/p193_s1.

Ruelle:1965:CF

[33] D. Ruelle. Correlation functionals. Journal of Mathematical Physics, 6(2):201–220, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i2/p201_s1.

Ahmavaara:1965:SSFb

[34] Yrjö Ahmavaara. The structure of space and the formalism of relativistic quantum theory. II. Journal of Mathematical Physics, 6(2):220– 227, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i2/p220_s1.

Bronk:1965:EER

[35] Burt V. Bronk. Exponential ensemble for random matrices. Journal of Mathematical Physics, 6(2):228–237, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i2/p228_s1.

Ginibre:1965:RDMa

[36] Jean Ginibre. Reduced density matrices of quantum gases. I. Limit of infinite volume. Journal of Mathematical Physics, 6(2):238–251, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i2/p238_s1.

Ginibre:1965:RDMb

[37] Jean Ginibre. Reduced density matrices of quantum gases. II. Cluster property. Journal of Mathematical Physics, 6(2):252–262, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i2/p252_s1.

Zernik:1965:TPI

[38] W. Zernik and R. W. Klopfenstein. Two-photon ionization of atomic hydrogen. II. Journal of Mathematical Physics, 6(2):262–270, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i2/p262_s1.

Beran:1965:UAS

[39] M. Beran and P. Corson. Use of analytic signals in third- and fourthorder coherence functions. Journal of Mathematical Physics, 6(2):271– 282, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i2/p271_s1.

Sykes:1965:DLT

[40] M. F. Sykes, J. W. Essam, and D. S. Gaunt. Derivation of lowtemperature expansions for the Ising model of a ferromagnet and an antiferromagnet. Journal of Mathematical Physics, 6(2):283–298, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i2/p283_s1.

Kallen:1965:SCF

[41] G. Källén and J. Toll. Special class of Feynman integrals in twodimensional space–time. Journal of Mathematical Physics, 6(2):299– 303, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i2/p299_s1.

Mattis:1965:ESM

[42] Daniel C. Mattis and Elliott H. Lieb. Exact solution of a many-fermion system and its associated boson field. Journal of Mathematical Physics, 6(2):304–312, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i2/p304_s1.

Schindler:1965:APO

[43] Karl Schindler. Adiabatic particle orbits in discontinuous fields. Journal of Mathematical Physics, 6(2):313–321, February 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i2/p313_s1.

Redei:1965:DEU

[44] L. B. Redei. The Dirac equation and the unitary representations of the inhomogeneous Lorentz group. Journal of Mathematical Physics, 6(2):322– 326, February 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i2/p322_s1.

Oberman:1965:SCL

[45] C. Oberman and M. Kruskal. Some constants of the linearized motion of Vlasov plasmas. Journal of Mathematical Physics, 6(3):327–335, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i3/p327_s1.

Zimering:1965:MME

[46] Samson Zimering. Mathematical methods for evaluating three-body interactions between closed-shell atoms or ions. Journal of Mathematical Physics, 6(3):336–347, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i3/p336_s1.

Mohling:1965:QSM

[47] Franz Mohling and W. T. Grandy. Quantum statistics of multicomponent systems. Journal of Mathematical Physics, 6(3):348–379, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i3/p348_s1.

Ezawa:1965:QMM

[48] Hiroshi Ezawa. Quantum mechanics of a many-boson system and the representation of canonical variables. Journal of Mathematical Physics, 6(3): 380–404, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i3/p380_s1.

Jepsen:1965:DSM

[49] D. W. Jepsen. Dynamics of a simple many-body system of hard rods. Journal of Mathematical Physics, 6(3):405–413, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i3/p405_s1.

Erickson:1965:LSA

[50] G. W. Erickson and H. M. Fried. Lepton scattering amplitudes in two model field theories. Journal of Mathematical Physics, 6(3):414– 423, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i3/p414_s1.

Schmidt:1965:AEI

[51] Paul W. Schmidt. Asymptotic expansion of the intensity of the smallangle X-ray scattering from cylinders of arbitrary cross section. Journal of Mathematical Physics, 6(3):424–431, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i3/p424_s1.

McGuire:1965:IFO

[52] J. B. McGuire. Interacting fermions in one dimension. I. Repulsive potential. Journal of Mathematical Physics, 6(3):432–439, March 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i3/p432_s1.

Ginibre:1965:SEC

[53] Jean Ginibre. Statistical ensembles of complex, quaternion, and real matrices. Journal of Mathematical Physics, 6(3):440–449, March 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i3/p440_s1.

Revzen:1965:SCA

[54] Michael Revzen. Self-consistent approximations in many-body systems. Journal of Mathematical Physics, 6(3):450–453, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i3/p450_s1.

Bailey:1965:SRD

[55] Paul B. Bailey and G. Milton Wing. Some recent developments in invariant imbedding with applications. Journal of Mathematical Physics, 6(3): 453–462, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i3/p453_s1.

Halpern:1965:PFD

[56] Francis R. Halpern and William Wilson. Poles in Feynman diagrams with several loops. Journal of Mathematical Physics, 6(3):463–469, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i3/p463_s1.

Gerstein:1965:CCO

[57] I. S. Gerstein. Construction of the charge operator for higher symmetry schemes. Journal of Mathematical Physics, 6(3):469–473, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i3/p469_s1.

Heffner:1965:THA

[58] H. Heffner and W. H. Louisell. Transformation having applications in quantum mechanics. Journal of Mathematical Physics, 6(3):474– 478, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i3/p474_s1.

McKean:1965:IPA

[59] H. P. McKean, M. Schreiber, and G. H. Weiss. Isoperimetric problem with application to the figure of cells. Journal of Mathematical Physics, 6(3):479–484, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i3/p479_s1.

Krzywicki:1965:CCP

[60] A. Krzywicki. Calculation of certain phase-space integrals. Journal of Mathematical Physics, 6(3):485–487, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i3/p485_s1.

Redei:1965:CPR

[61] L. B. Redei. Composite particles in a relativistic model of two-body scattering. Journal of Mathematical Physics, 6(3):487–491, March 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i3/p487_s1.

Ashour:1965:FSE

[62] Attia A. Ashour. Fourier series expansion for the general power of the distance between two points. Journal of Mathematical Physics, 6(3): 492–494, March 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i3/p492_s1.

Rohrlich:1965:MAQ

[63] F. Rohrlich and J. C. Stoddart. Microcausality in asymptotic quantum field theory. Journal of Mathematical Physics, 6(4):495–499, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i4/p495_s1.

Greenberg:1965:HOL

[64] O. W. Greenberg and A. M. L. Messiah. High-order limit of para-Bose and para-Fermi fields. Journal of Mathematical Physics, 6(4):500–504, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i4/p500_s1.

Ford:1965:SMA

[65] G. W. Ford, M. Kac, and P. Mazur. Statistical mechanics of assemblies of coupled oscillators. Journal of Mathematical Physics, 6(4):504– 515, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i4/p504_s1.

Aks:1965:PSI

[66] Stanley Ø. Aks. Proof that scattering implies production in quantum field theory. Journal of Mathematical Physics, 6(4):516–532, April 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i4/p516_s1.

Dragt:1965:CTP

[67] A. J. Dragt. Classification of three-particle states according to SU_{3}. Journal of Mathematical Physics, 6(4):533–553, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i4/p533_s1.

Bloch:1965:DRV

[68] Claude Bloch and James S. Langer. Diagram renormalization, variational principles, and the infinite-dimensional Ising model. Journal of Mathematical Physics, 6(4):554–572, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i4/p554_s1.

Tiktopoulos:1965:NP

[69] George Tiktopoulos. Note on positronium. Journal of Mathematical Physics, 6(4):573–577, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i4/p573_s1.

Alcaras:1965:ACB

[70] J. A. Castilho Alcarás and P. Leal Ferreira. Algebraic construction of the basis for the irreducible representations of rotation groups and for the homogeneous Lorentz group. Journal of Mathematical Physics, 6(4): 578–583, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i4/p578_s1.

Trumper:1965:SCT

[71] Manfred Trümper. On a special class of type-I gravitational fields. Journal of Mathematical Physics, 6(4):584–589, April 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i4/p584_s1.

Ozsvath:1965:NHS

[72] István Ozsváth. New homogeneous solutions of Einstein’s field equations with incoherent matter obtained by a spinor technique. Journal of Mathematical Physics, 6(4):590–610, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i4/p590_s1.

Wilcox:1965:UAE

[73] Calvin H. Wilcox. Uniform asymptotic estimates for wave packets in the quantum theory of scattering. Journal of Mathematical Physics, 6(4): 611–620, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i4/p611_s1.

Nesbet:1965:EOI

[74] R. K. Nesbet. Excitation operators and intrinsic Hamiltonians. Journal of Mathematical Physics, 6(4):621–629, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i4/p621_s1.

Abraham:1965:ASP

[75] B. W. Abraham. Approach to stability of a plasma. Journal of Mathematical Physics, 6(4):630–636, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i4/p630_s1.

Bessis:1965:UBD

[76] D. Bessis. Upper bound for the double spectral function in potential scattering. Journal of Mathematical Physics, 6(4):637–643, April 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i4/p637_s1.

Gray:1965:USS

[77] A. H. Gray. Uniqueness of steady-state solutions to the Fokker– Planck equation. Journal of Mathematical Physics, 6(4):644–647, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i4/p644_s1.

Rockmore:1965:PVF

[78] R. Rockmore. Property of the vertex function in potential theory. Journal of Mathematical Physics, 6(4):648–651, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i4/p648_s1.

Pao:1965:VMC

[79] Yoh-Han Pao and H. L. Frisch. Vertex modification for Coulombic interactions by exact summation of ladder diagrams. Journal of Mathematical Physics, 6(4):652–655, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i4/p652_s1.

Tomozawa:1965:ISP

[80] Yukio Tomozawa. Internal symmetry and the Poincaré group. Journal of Mathematical Physics, 6(4):656–657, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i4/p656_s1.

Feenberg:1965:NCR

[81] Eugene Feenberg. Necessary condition on the radial distribution function. Journal of Mathematical Physics, 6(4):658–659, April 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i4/p658_s1.

Greenman:1965:HEB

[82] J. V. Greenman. High-energy behavior at fixed angle for the five-point function in perturbation theory. Journal of Mathematical Physics, 6(4):

660–670, April 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i4/p660_s1.

Lomont:1965:ERIa

[83] J. S. Lomont and H. E. Moses. Errata: “The Representations of the Inhomogeneous Lorentz Group in Terms of an Angular Momentum Basis”. Journal of Mathematical Physics, 6(4):670, April 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i4/p670_s1. See [1].

Lomont:1965:ERIb

[84] J. S. Lomont and H. E. Moses. Errata: “Representations of the Inhomogeneous Lorentz Group in Terms of an Angular Momentum Basis: Derivation for the Cases of Nonzero Mass and Zero Mass, Discrete Spin”. Journal of Mathematical Physics, 6(4):670, April 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i4/p670_s2. See [5].

Gien:1965:RFL

[85] Tran Trong Gien. Relativistic formulation of the lifetime matrix in the potential theory of collision. Journal of Mathematical Physics, 6(5): 671–676, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i5/p671_s1.

Beg:1965:SHF

[86] Mirza A. B. Bég and Henri Ruegg. A set of harmonic functions for the group SU(3). Journal of Mathematical Physics, 6(5):677–682, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i5/p677_s1.

Nagel:1965:OLR

[87] John G. Nagel and Marcos Moshinsky. Operators that lower or raise the irreducible vector spaces of U_{n−1} contained in an irreducible vector space of U_{n}. Journal of Mathematical Physics, 6(5):682–694, May 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i5/p682_s1.

Imamura:1965:SCW

[88] Tsutomu Imamura, William C. Meecham, and Armand Siegel. Symbolic calculus of the Wiener process and Wiener–Hermite functionals. Journal of Mathematical Physics, 6(5):695–706, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i5/p695_s1.

Siegel:1965:WHE

[89] Armand Siegel, Tsutomu Imamura, and William C. Meecham. Wiener– Hermite expansion in model turbulence in the late decay stage. Journal of Mathematical Physics, 6(5):707–721, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i5/p707_s1.

Kuriyan:1965:ATC

[90] J. G. Kuriyan, D. Lurié, and A. J. Macfarlane. Algebraic tabulation of Clebsch–Gordan coefficients of SU_{3} for the product (λ, μ) ⊗ (1, 1) of representations of SU_{3}. Journal of Mathematical Physics, 6(5):722–733, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i5/p722_s1.

Klauder:1965:DCS

[91] John R. Klauder, James McKenna, and Douglas G. Currie. On “diagonal” coherent-state representations for quantum-mechanical density matrices. Journal of Mathematical Physics, 6(5):734–739, May 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i5/p734_s1.

Pugh:1965:OTF

[92] Robert E. Pugh. S-operator theory. I. Formulation. Journal of Mathematical Physics, 6(5):740–746, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i5/p740_s1.

Mukunda:1965:TMU

[93] N. Mukunda and L. K. Pandit. Tensor methods and a unified representation theory of SU_{3}. Journal of Mathematical Physics, 6(5):746–765, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i5/p746_s1.

Danos:1965:MME

[94] M. Danos and L. C. Maximon. Multipole matrix elements of the translation operator. Journal of Mathematical Physics, 6(5):766–778, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i5/p766_s1.

Tiersten:1965:VPS

[95] H. F. Tiersten. Variational principle for saturated magnetoelastic insulators. Journal of Mathematical Physics, 6(5):779–787, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i5/p779_s1.

Barnes:1965:LTS

[96] K. J. Barnes. Lagrangian theory for the second-rank tensor field. Journal of Mathematical Physics, 6(5):788–794, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i5/p788_s1.

Huang:1965:EBS

[97] Kerson Huang and Francis E. Low. Exact bootstrap solutions in some static models of meson-baryon scattering. Journal of Mathematical Physics, 6(5):795–816, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i5/p795_s1.

Hermann:1965:CMS

[98] Robert Hermann. Compactification of Minkowski space and SU(3) symmetry. Journal of Mathematical Physics, 6(5):817–819, May 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i5/p817_s1.

Bessis:1965:PSP

[99] D. Bessis. Problem of subtractions in potential scattering. Journal of Mathematical Physics, 6(5):820–824, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i5/p820_s1.

Federbush:1965:NNL

[100] Paul Federbush. Note on non-Landau singularities. Journal of Mathematical Physics, 6(5):825–826, May 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i5/p825_s1.

Gunson:1965:UMSa

[101] J. Gunson. Unitarity and on-mass-shell analyticity as a basis for Smatrix theories. I. Journal of Mathematical Physics, 6(6):827–844, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p827_s1.

Gunson:1965:UMSb

[102] J. Gunson. Unitarity and on-mass-shell analyticity as a basis for Smatrix theories. II. Journal of Mathematical Physics, 6(6):845–851, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p845_s1.

Gunson:1965:UMSc

[103] J. Gunson. Unitarity and on-mass-shell analyticity as a basis for Smatrix theories. III. Journal of Mathematical Physics, 6(6):852–858, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p852_s1.

Saenz:1965:LTB

[104] Albert W. Sáenz. Long-time behavior of the electric potential and stability in the linearized Vlasov theory. Journal of Mathematical Physics, 6(6):859–875, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i6/p859_s1.

Karayianis:1965:PAM

[105] N. Karayianis and C. A. Morrison. Projected angular momentum states. Journal of Mathematical Physics, 6(6):876–879, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i6/p876_s1.

Fradkin:1965:CQA

[106] D. M. Fradkin. Conserved quantities associated with symmetry transformations of relativistic free-particle equations of motion. Journal of Mathematical Physics, 6(6):879–890, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i6/p879_s1.

Eger:1965:SIS

[107] M. Eger and E. P. Gross. Spatially inhomogeneous states of many-body systems. Journal of Mathematical Physics, 6(6):891–901, June 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i6/p891_s1.

Janis:1965:SGS

[108] A. I. Janis and E. T. Newman. Structure of gravitational sources. Journal of Mathematical Physics, 6(6):902–914, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i6/p902_s1.

Newman:1965:NKS

[109] E. T. Newman and A. I. Janis. Note on the Kerr spinning-particle metric. Journal of Mathematical Physics, 6(6):915–917, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i6/p915_s1.

Newman:1965:MRC

[110] E. T. Newman, E. Couch, K. Chinnapared, A. Exton, A. Prakash, and R. Torrence. Metric of a rotating, charged mass. Journal of Mathematical Physics, 6(6):918–919, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i6/p918_s1.

Marlow:1965:UDN

[111] A. R. Marlow. Unified Dirac–von Neumann formulation of quantum mechanics. I. Mathematical theory. Journal of Mathematical Physics, 6(6): 919–927, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i6/p919_s1.

Moses:1965:TLMa

[112] H. E. Moses. Transformation from a linear momentum to an angular momentum basis for particles of zero mass and finite spin. Journal of Mathematical Physics, 6(6):928–940, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i6/p928_s1. See erratum [263].

Federbush:1965:CSH

[113] Paul Federbush. Calculation of some homology groups relevant to sixthorder Feynman diagrams. Journal of Mathematical Physics, 6(6):941– 954, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i6/p941_s1.

Pagnamenta:1965:SKP

[114] Antonio Pagnamenta. Solution of the Källén–Pauli equation. Journal of Mathematical Physics, 6(6):955–965, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i6/p955_s1.

Kane:1965:ISO

[115] J. Kane and E. R. Surynarayan. Integrals of the second-order linear differential equation. Journal of Mathematical Physics, 6(6):966–974, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p966_s1.

Mandelbrojt:1965:MSC

[116] J. Mandelbrojt and S. Mandelbrojt. Mathematical study of the cutoff procedure for divergent integrals in bound-state problems. Journal of Mathematical Physics, 6(6):974–980, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i6/p974_s1.

Norton:1965:QFT

[117] Richard E. Norton. On the quantum field theories leading to the Corben equations. Journal of Mathematical Physics, 6(6):981–992, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p981_s1.

Gerjuoy:1965:MTC

[118] E. Gerjuoy. Momentum transfer cross-section theorem. Journal of Mathematical Physics, 6(6):993–996, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i6/p993_s1.

Leader:1965:NBA

[119] Elliot Leader. Note on Bég’s approach to peratization. Journal of Mathematical Physics, 6(6):996–998, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i6/p996_s1.

Klink:1965:DGH

[120] W. H. Klink. Disconnected groups as higher symmetry groups. Journal of Mathematical Physics, 6(6):999–1004, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i6/p999_s1.

Fujita:1965:GBE

[121] S. Fujita. Generalized Boltzmann equation for a quantum gas obeying classical statistics. Journal of Mathematical Physics, 6(6):1004–1008, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i6/p1004_s1.

Jones:1965:FEN

[122] Gerald L. Jones and Michael Davies. Fluctuations in equilibrium and nonequilibrium ensembles. Journal of Mathematical Physics, 6(6):1009– 1014, June 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i6/p1009_s1.

Baxter:1965:WFT

[123] Glen Baxter. Weight factors for the two-dimensional Ising model. Journal of Mathematical Physics, 6(7):1015–1021, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i7/p1015_s1.

Kibble:1965:CLF

[124] T. W. B. Kibble. Conservation laws for free fields. Journal of Mathematical Physics, 6(7):1022–1026, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i7/p1022_s1.

Penney:1965:TDN

[125] R. Penney. Tensorial description of neutrinos. Journal of Mathematical Physics, 6(7):1026–1028, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i7/p1026_s1.

Penney:1965:GMS

[126] R. Penney. Geometrization of a massive scalar field. Journal of Mathematical Physics, 6(7):1029–1031, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i7/p1029_s1.

Penney:1965:BF

[127] R. Penney. Bosons and fermions. Journal of Mathematical Physics, 6(7): 1031–1034, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i7/p1031_s1.

Richardson:1965:EEP

[128] R. W. Richardson. Exact eigenstates of the pairing-force hamiltonian. II. Journal of Mathematical Physics, 6(7):1034–1051, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1034_s1.

Peres:1965:QEI

[129] Asher Peres. Quantum electrodynamics without indefinite metric. Journal of Mathematical Physics, 6(7):1052–1056, July 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1052_s1.

Rankin:1965:QMT

[130] Bayard Rankin. Quantum mechanical time. Journal of Mathematical Physics, 6(7):1057–1071, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i7/p1057_s1.

Arens:1965:ADP

[131] Richard Arens and Donald Babbitt. Algebraic difficulties of preserving dynamical relations when forming quantum-mechanical operators. Journal of Mathematical Physics, 6(7):1071–1075, July 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1071_s1.

Fried:1965:FEP

[132] H. M. Fried. Functional expansion in peratization theory. Journal of Mathematical Physics, 6(7):1076–1082, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i7/p1076_s1.

Girardeau:1965:DLR

[133] M. D. Girardeau. Off-diagonal long-range order and generalized Bose condensation. Journal of Mathematical Physics, 6(7):1083–1098, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i7/p1083_s1.

Pandres:1965:SBS

[134] Dave Pandres. Schrödinger basis for spinor representations of the threedimensional rotation group. Journal of Mathematical Physics, 6(7):1098– 1102, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i7/p1098_s1.

Ullah:1965:ARW

[135] Nazakat Ullah. Asymptotic reduced-width amplitude distributions. Journal of Mathematical Physics, 6(7):1102–1105, July 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1102_s1.

Calogero:1965:SCAb

[136] F. Calogero. Sufficient conditions for an attractive potential to possess bound states. II. Journal of Mathematical Physics, 6(7):1105–1107, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i7/p1105_s1.

Korringa:1965:VPD

[137] J. Korringa. Variational principle for diffraction of elastic waves. Journal of Mathematical Physics, 6(7):1107–1114, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i7/p1107_s1.

Aly:1965:SSL

[138] H. H. Aly, Riazuddin, and A. H. Zimerman. Scattering by singular logarithmic potential. Journal of Mathematical Physics, 6(7):1115–1117, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i7/p1115_s1.

Zimmerman:1965:EFF

[139] Robert L. Zimmerman. Evaluation of Feynman’s functional integrals. Journal of Mathematical Physics, 6(7):1117–1124, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1117_s1.

Kuscer:1965:TDO

[140] I. Kuscer and P. F. Zweifel. Time-dependent one-speed albedo problem for a semi-infinite medium. Journal of Mathematical Physics, 6(7):1125– 1130, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i7/p1125_s1.

Fikioris:1965:AES

[141] J. G. Fikioris. Asymptotic expansions of solutions of differential equations. Journal of Mathematical Physics, 6(7):1131–1148, July 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1131_
s1.

Taylor:1965:PSB

[142] J. G. Taylor. On polynomial systems in a Banach ring. Journal of Mathematical Physics, 6(7):1148–1151, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v6/i7/p1148_s1. See erratum [437].

Morita:1965:ESW

[143] Tohru Morita and Tomoyasu Tanaka. Effect of n-spin-wave interaction on the low-temperature spontaneous magnetization. Journal of Mathematical Physics, 6(7):1152–1154, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i7/p1152_s1.

Symanzik:1965:PRI

[144] Kurt Symanzik. Proof and refinements of an inequality of Feynman. Journal of Mathematical Physics, 6(7):1155–1156, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1155_s1.

Gilbert:1965:SAF

[145] R. P. Gilbert, H. C. Howard, and S. Aks. Singularities of analytic functions having integral representations, with a remark about the elastic unitarity integral. Journal of Mathematical Physics, 6(7):1157–1162, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i7/p1157_s1.

Redmond:1965:SKG

[146] Peter J. Redmond. Solution of the Klein–Gordon and Dirac equations for a particle with a plane electromagnetic wave and a parallel magnetic field. Journal of Mathematical Physics, 6(7):1163–1169, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i7/p1163_s1.

Sommerfield:1965:SIE

[147] Charles M. Sommerfield. Solution of the integral equation for V θ scattering in the Lee model. Journal of Mathematical Physics, 6(7):1170– 1172, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i7/p1170_s1.

Jaffe:1965:AQF

[148] Arthur Jaffe. On the approximation of quantum field theories. Journal of Mathematical Physics, 6(7):1172–1178, July 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i7/p1172_s1.

Arf:1965:ASC

[149] Cahït Arf, Kaya Ïmre, and Ercüment Özïzmïr. On the algebraic structure of the cluster expansion in statistical mechanics. Journal of Mathematical Physics, 6(8):1179–1188, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v6/i8/p1179_s1.

Sherman:1965:CSI

[150] S. Sherman. Cluster-star inversion and the Möbius formula. Journal of Mathematical Physics, 6(8):1189–1192, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i8/p1189_s1.

Stell:1965:CSI

[151] G. Stell. Cluster-star inversion by means of generating functionals. Journal of Mathematical Physics, 6(8):1193–1200, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1193_s1.

Bazanski:1965:DLT

[152] Stanislaw L. Bazanski. Decomposition of the Lorentz transformation matrix into skew-symmetric tensors. Journal of Mathematical Physics, 6(8):1201–1203, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i8/p1201_s1.

Karayianis:1965:ATE

[153] N. Karayianis. Atomic terms for equivalent electrons. Journal of Mathematical Physics, 6(8):1204–1209, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i8/p1204_s1.

Langerholc:1965:TFQ

[154] John Langerholc. Trace formalism for quantum mechanical expectation values. Journal of Mathematical Physics, 6(8):1210–1218, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i8/p1210_s1.

Kerner:1965:CPV

[155] Edward H. Kerner. Can the position variable be a canonical coordinate in a relativistic many-particle theory? Journal of Mathematical Physics, 6(8):1218–1227, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v6/i8/p1218_s1.

Bose:1965:QEV

[156] S. K. Bose and S. N. Biswas. Quantum electrodynamics with vanishing bare fermion mass. Journal of Mathematical Physics, 6(8):1227– 1230, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1227_s1.

Levine:1965:SSW

[157] Harold Levine. Scattering of surface waves by a submerged circular cylinder. Journal of Mathematical Physics, 6(8):1231–1243, August 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1231_
s1.

Moses:1965:TLMb

[158] H. E. Moses. Transformation from a linear momentum to an angular momentum basis for relativistic particles of nonzero mass and any spin. Journal of Mathematical Physics, 6(8):1244–1254, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1244_s1.

Ozsvath:1965:HSE

[159] István Ozsváth. Homogeneous solutions of the Einstein–Maxwell equations. Journal of Mathematical Physics, 6(8):1255–1265, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/
p1255_s1.

Ozsvath:1965:AHS

[160] István Ozsváth. All homogeneous solutions of Einstein’s field equations with incoherent matter and electromagnetic radiation. Journal of Mathematical Physics, 6(8):1265–1268, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i8/p1265_s1.

Rosen:1965:PSN

[161] Gerald Rosen. Particlelike solutions to nonlinear scalar wave theories. Journal of Mathematical Physics, 6(8):1269–1272, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1269_s1.

Rudra:1965:PRF

[162] P. Rudra. On projective representations of finite groups. Journal of Mathematical Physics, 6(8):1273–1277, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i8/p1273_s1.

Rudra:1965:IRS

[163] P. Rudra. On irreducible representations of space groups. Journal of Mathematical Physics, 6(8):1278–1282, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i8/p1278_s1.

Lebowitz:1965:SIP

[164] J. L. Lebowitz, G. Stell, and S. Baer. Separation of the interaction potential into two parts in treating many-body systems. I. General theory and applications to simple fluids with short-range and long-range forces. Journal of Mathematical Physics, 6(8):1282–1298, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1282_s1.

Schild:1965:FAP

[165] A. Schild and J. A. Schlosser. Fokker action principle for particles with charge, spin, and magnetic moment. Journal of Mathematical Physics, 6(8):1299–1306, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i8/p1299_s1.

Shin:1965:BRP

[166] Ernest E. H. Shin. Born reciprocity principle and unitary symmetry. Journal of Mathematical Physics, 6(8):1307–1309, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1307_s1.

Penney:1965:GTN

[167] R. Penney. Geometric theory of neutrinos. Journal of Mathematical Physics, 6(8):1309–1314, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v6/i8/p1309_s1.

Tharrats:1965:GOP

[168] J. Tharrats, O. Cerceau, and O. Rojo. General operator potential for the two-nucleon system. Journal of Mathematical Physics, 6(8):1315– 1320, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1315_s1.

Eyges:1965:IMQ

[169] Leonard Eyges. Improved method for quantum mechanical three-body problems. I. attractive potentials. Journal of Mathematical Physics, 6(8):1320–1328, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i8/p1320_s1.

Sudarshan:1965:CST

[170] E. C. G. Sudarshan. Concerning space–time, symmetry groups, and charge conservation. Journal of Mathematical Physics, 6(8):1329–1331, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1329_s1.

Chen:1965:DST

[171] Yung Ming Chen. Diffraction by a smooth transparent object. II. Diffraction by a cylindrical cavity. Journal of Mathematical Physics, 6(8):1332– 1333, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1332_s1.

Pease:1965:SIRb

[172] M. C. Pease. Scalar invariants of a rotational system in a Lie algebra. II. Casimir operators. Journal of Mathematical Physics, 6(8):1334– 1340, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1334_s1.

Lowdin:1965:SPTa

[173] Per-Olov Löwdin. Studies in perturbation theory. IX. Connection between various approaches in the recent development-evaluation of upper bounds to energy eigenvalues in Schrödinger’s perturbation theory. Journal of Mathematical Physics, 6(8):1341–1353, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i8/p1341_s1.

Roos:1965:ECU

[174] Matts Roos. Erratum: “On the Construction of a Unitary Matrix with Elements of Given Moduli”. Journal of Mathematical Physics, 6(8): 1354, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1354_s1. See [6].

Silven:1965:EAO

[175] Saul Silven. Errata: “Application of Operational Methods to The Analysis of Uniform Plasmas”. Journal of Mathematical Physics, 6(8): 1354, August 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i8/p1354_s2. See [2].

Hagen:1965:RRRa

[176] C. R. Hagen and A. J. Macfarlane. Reduction of representations of SU_{mn} with respect to the subgroup SU_{m} ⊗ SU_{n}. Journal of Mathematical Physics, 6(9):1355–1365, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i9/p1355_s1.

Hagen:1965:RRRb

[177] C. R. Hagen and A. J. Macfarlane. Reduction of representations of SU_{m+n} with respect to the subgroup SU_{m} ⊗ SU_{n}. Journal of Mathematical Physics, 6(9):1366–1371, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i9/p1366_s1.

Levy-Leblond:1965:SCT

[178] Jean-Marc Lévy-Leblond and Monique Lévy-Nahas. Symmetrical coupling of three angular momenta. Journal of Mathematical Physics, 6 (9):1372–1380, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i9/p1372_s1.

Itzykson:1965:PRI

[179] Claude Itzykson. On the possibility of relating internal symmetries and Lorentz invariance. Journal of Mathematical Physics, 6(9):1381–1386, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i9/p1381_s1.

Szekeres:1965:GC

[180] P. Szekeres. The gravitational compass. Journal of Mathematical Physics, 6(9):1387–1391, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i9/p1387_s1.

Thompson:1965:ADP

[181] Colin J. Thompson. Algebraic derivation of the partition function of a two-dimensional Ising model. Journal of Mathematical Physics, 6 (9):1392–1395, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i9/p1392_s1.

Gerjuoy:1965:MTT

[182] E. Gerjuoy. Momentum-transfer theorem for inelastic processes. Journal of Mathematical Physics, 6(9):1396–1402, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i9/p1396_s1.

Johnson:1965:PUW

[183] Charles A. Johnson and G. D. Chakerian. On the proof and uniqueness of Wulff’s construction of the shape of minimum surface free energy. Journal of Mathematical Physics, 6(9):1403–1404, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i9/p1403_s1.

Papapetrou:1965:TNE

[184] A. Papapetrou. Theorem on nonradiative electromagnetic and gravitational fields. Journal of Mathematical Physics, 6(9):1405–1410, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i9/p1405_s1.

Cohen:1965:BSS

[185] Hirsh Cohen and Farouk Odeh. Backward superconducting switching. Journal of Mathematical Physics, 6(9):1411–1424, September 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i9/p1411_
s1.

Coleman:1965:SFD

[186] A. J. Coleman. Structure of fermion density matrices. II. Antisymmetrized geminal powers. Journal of Mathematical Physics, 6(9):1425– 1431, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i9/p1425_s1.

Ginibre:1965:RDMc

[187] Jean Ginibre. Reduced density matrices of quantum gases. III. Hard-core potentials. Journal of Mathematical Physics, 6(9):1432–1446, September 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i9/p1432_s1.

Griffiths:1965:MEQ

[188] Robert B. Griffiths. Microcanonical ensemble in quantum statistical mechanics. Journal of Mathematical Physics, 6(10):1447–1461, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1447_s1.

May:1965:MPC

[189] Robert M. May. Magnetic properties of charged ideal quantum gases in n dimensions. Journal of Mathematical Physics, 6(10):1462–1468, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1462_s1.

Boyling:1965:PNI

[190] J. B. Boyling. A proof of Nakanishi’s inequality. Journal of Mathematical Physics, 6(10):1469–1471, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v6/i10/p1469_s1.

Swift:1965:MRP

[191] Arthur R. Swift. Mixing of Regge poles and cuts in a field-theory model. Journal of Mathematical Physics, 6(10):1472–1481, October 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1472_
s1.

Hakim:1965:CTR

[192] Rémi Hakim. A covariant theory of relativistic Brownian motion I. Local equilibrium. Journal of Mathematical Physics, 6(10):1482–1495, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1482_s1.

Andrews:1965:SBE

[193] Frank C. Andrews. On the solution of the BBGKY equations for a dense classical gas. Journal of Mathematical Physics, 6(10):1496–1505, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1496_s1.

Gurtin:1965:VPT

[194] Morton E. Gurtin. Variational principle for the time-dependent schrödinger equation. Journal of Mathematical Physics, 6(10):1506– 1507, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i10/p1506_s1.

Sarason:1965:HSS

[195] Leonard Sarason. Hydromagnetic switch-on shocks. Journal of Mathematical Physics, 6(10):1508–1518, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i10/p1508_s1.

Voisin:1965:SUR

[196] J. Voisin. On some unitary representations of the Galilei group I. Irreducible representations. Journal of Mathematical Physics, 6(10):1519– 1529, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i10/p1519_s1.

Balachandran:1965:ISL

[197] A. P. Balachandran. Inequalities for the solutions of linear integral equations. Journal of Mathematical Physics, 6(10):1530–1533, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1530_s1.

Whippman:1965:BRS

[198] M. L. Whippman. Branching rules for simple Lie groups. Journal of Mathematical Physics, 6(10):1534–1539, October 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1534_s1.

Brody:1965:RRW

[199] T. A. Brody, M. Moshinsky, and I. Renero. Recursion relations for the Wigner coefficients of unitary groups. Journal of Mathematical Physics, 6(10):1540–1546, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i10/p1540_s1.

Pandit:1965:ATC

[200] L. K. Pandit and N. Mukunda. Algebraic tabulation of the Clebsch– Gordan coefficients for reduction of the product (λ, μ) ⊕ (3, 0) of irreducible representations of SU(3). Journal of Mathematical Physics, 6(10):1547–1557, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i10/p1547_s1.

Pease:1965:FLH

[201] M. C. Pease. Formalization of the lagrangian, the hamiltonian, and related concepts. Journal of Mathematical Physics, 6(10):1558–1563, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1558_s1.

Levy-Leblond:1965:PKG

[202] Jean-Marc Lévy-Leblond and François Lurçat. N-particle kinematics and group-theoretical treatment of phase space I. Nonrelativistic. Journal of Mathematical Physics, 6(10):1564–1570, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1564_s1.

Levy-Leblond:1965:TPN

[203] Jean-Marc Lévy-Leblond and Monique Lévy-Nahas. Three-particle nonrelativistic kinematics and phase space. Journal of Mathematical Physics, 6(10):1571–1575, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i10/p1571_s1.

Iagolnitzer:1965:MTD

[204] Daniel Iagolnitzer. S-matrix theory and double scattering. Journal of Mathematical Physics, 6(10):1576–1583, October 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1576_s1.

Derome:1965:RAA

[205] Jean-Robert Derome and W. T. Sharp. Racah algebra for an arbitrary group. Journal of Mathematical Physics, 6(10):1584–1590, October 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i10/p1584_s1.

Hammer:1965:TDP

[206] C. L. Hammer and T. A. Weber. Time-dependent perturbation theory. Journal of Mathematical Physics, 6(10):1591–1606, October 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i10/p1591_
s1.

Penney:1965:DRW

[207] R. Penney. On the dimensionality of the real world. Journal of Mathematical Physics, 6(11):1607–1611, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i11/p1607_s1.

Denman:1965:ICL

[208] Harry H. Denman. Invariance and conservation laws in classical mechanics. Journal of Mathematical Physics, 6(11):1611–1616, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1611_s1.

Fikioris:1965:EFI

[209] John G. Fikioris. Electromagnetic field inside a current-carrying region. Journal of Mathematical Physics, 6(11):1617–1620, November 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1617_
s1.

Dragt:1965:RTP

[210] A. J. Dragt. Relativistic three-particle SU_{3} states. Journal of Mathematical Physics, 6(11):1621–1625, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i11/p1621_s1.

Aks:1965:APE

[211] S. Ø. Aks, R. P. Gilbert, and H. C. Howard. Analytic properties of the elastic unitarity integral. Journal of Mathematical Physics, 6(11):1626– 1634, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1626_s1.

Rothe:1965:CSR

[212] Heinz J. Rothe. Crossing symmetric Regge representation for the invariant scattering amplitude. Journal of Mathematical Physics, 6(11):1635– 1642, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1635_s1.

Fisher:1965:CFC

[213] Michael E. Fisher. Correlation functions and the coexistence of phases. Journal of Mathematical Physics, 6(11):1643–1653, November 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1643_
s1.

Gelman:1965:AFS

[214] Harry Gelman and Kurt Haller. Alternative formulations of scattering in model field theories. Journal of Mathematical Physics, 6(11):1653– 1663, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1653_s1.

Chambers:1965:HDP

[215] Charles M. Chambers. Higher-dimensional periodic systems. Journal of Mathematical Physics, 6(11):1664–1666, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1664_s1.

Klauder:1965:RSM

[216] John R. Klauder. Rotationally-symmetric model field theories. Journal of Mathematical Physics, 6(11):1666–1679, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1666_s1.

Shapiro:1965:MRA

[217] Joseph Shapiro. Matrix representation of the angular momentum projection operator. Journal of Mathematical Physics, 6(11):1680–1691, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1680_s1.

Burke:1965:SAI

[218] E. A. Burke. Solution of an atomic integral containing three odd powers of interelectronic separation coordinates. Journal of Mathematical Physics, 6(11):1691–1695, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i11/p1691_s1.

Boardman:1965:CQP

[219] John Boardman. Contributions to the quantization problem in general relativity. Journal of Mathematical Physics, 6(11):1696–1701, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1696_s1.

Broido:1965:CUE

[220] Michael M. Broido. On the complete unitarity equations for pion-pion scattering. Journal of Mathematical Physics, 6(11):1702–1708, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1702_s1.

Carnahan:1965:TRP

[221] W. H. Carnahan. Threshold Regge poles for coupled channels. Journal of Mathematical Physics, 6(11):1709–1719, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1709_s1.

Catara:1965:BSW

[222] F. Catara and M. Di Toro. On the bound-state wavefunctions of a nonlocal solvable potential. Journal of Mathematical Physics, 6(11):1720– 1722, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1720_s1.

Chen:1965:SCS

[223] Joseph C. Y. Chen. On a set of coupled second-order differential equations. Journal of Mathematical Physics, 6(11):1723–1729, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1723_s1.

Cornille:1965:SLP

[224] Henri Cornille and Enrico Predazzi. Singular logarithmic potentials in coordinate space. Journal of Mathematical Physics, 6(11):1730–1740, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1730_s1.

Hardy:1965:PEL

[225] Robert J. Hardy, Robert J. Swenson, and William C. Schieve. Perturbation expansion for lattice thermal conductivity. Journal of Mathematical Physics, 6(11):1741–1748, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i11/p1741_s1.

Hardy:1965:LOC

[226] Robert J. Hardy. Lowest-order contribution to the lattice thermal conductivity. Journal of Mathematical Physics, 6(11):1749–1761, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1749_s1.

Hepp:1965:OPS

[227] Klaus Hepp. One-particle singularities of the S-matrix in quantum field theory. Journal of Mathematical Physics, 6(11):1762–1767, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1762_s1.

Hermann:1965:RGN

[228] Robert Hermann. Remarks on the geometric nature of quantum phase space. Journal of Mathematical Physics, 6(11):1768–1771, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1768_s1.

Kazes:1965:SIE

[229] E. Kazes. Solution of an integral equation in V –θ scattering. Journal of Mathematical Physics, 6(11):1772–1773, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1772_s1.

LeGaillard:1965:SFP

[230] Y. Le Gaillard and A. Visconti. On the structure of the finite parts of the generating functional of propagators in quantumelectrodynamics. Journal of Mathematical Physics, 6(11):1774–1785, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1774_s1.

Louck:1965:GTH

[231] James D. Louck. Group theory of harmonic oscillators in n-dimensional space. Journal of Mathematical Physics, 6(11):1786–1804, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i11/p1786_s1.

More:1965:NVS

[232] Richard M. More. New variational scheme for finding the excited states of one-dimensional systems. Journal of Mathematical Physics, 6(11):1804– 1805, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1804_s1.

Newman:1965:NDG

[233] Ezra T. Newman and T. W. J. Unti. Note on the dynamics of gravitational sources. Journal of Mathematical Physics, 6(11):1806–1811, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1806_s1.

Thompson:1965:IAS

[234] Colin J. Thompson. Inequality with applications in statistical mechanics. Journal of Mathematical Physics, 6(11):1812–1813, November 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1812_
s1.

Unz:1965:OIP

[235] H. Unz. Oblique incidence on plane boundary between two general gyrotropic plasma media. Journal of Mathematical Physics, 6(11):1813– 1821, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1813_s1.

Voisin:1965:URG

[236] J. Voisin. On the unitary representations of the Galilei group. II. Twoparticle systems. Journal of Mathematical Physics, 6(11):1822–1832, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1822_s1.

Yan:1965:OSM

[237] C. C. Yan and G. H. Wannier. Observations on the spherical model of a ferromagnet. Journal of Mathematical Physics, 6(11):1833–1838, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i11/p1833_s1.

Myers:1965:WSG

[238] John M. Myers. Wave scattering and the geometry of a strip. Journal of Mathematical Physics, 6(11):1839–1846, November 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i11/p1839_s1.

Baird:1965:RSL

[239] G. E. Baird and L. C. Biedenharn. On the representations of the semisimple Lie groups. V. Some explicit Wigner operators for SU_{3}. Journal of Mathematical Physics, 6(12):1847–1854, December 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i12/p1847_s1.

Bergmann:1965:RBH

[240] D. Bergmann and Y. Frishman. A relation between the hydrogen atom and multidimensional harmonic oscillators. Journal of Mathematical Physics, 6(12):1855–1856, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i12/p1855_s1.

Bevensee:1965:QES

[241] R. M. Bevensee. Quantum electrodynamics of the stimulated emission of radiation. II. Journal of Mathematical Physics, 6(12):1857–1870, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i12/p1857_s1.

Foldy:1965:CTS

[242] Leslie L. Foldy. Cartesian tensors in spaces with an indefinite metric with applications to relativity. Journal of Mathematical Physics, 6(12):1871– 1876, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1871_s1.

Fujita:1965:TEE

[243] S. Fujita. Thermodynamic evolution equation for a quantum statistical gas. Journal of Mathematical Physics, 6(12):1877–1885, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i12/p1877_s1.

Kallen:1965:MSA

[244] G. Källén and J. S. Toll. Momentum-space analyticity properties of the Bergman–Weil integral for the three point function. Journal of Mathematical Physics, 6(12):1885–1901, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v6/i12/p1885_s1.

Kane:1965:GFS

[245] Joseph W. Kane and Leo P. Kadanoff. Green’s functions and superfluid hydrodynamics. Journal of Mathematical Physics, 6(12):1902–1912, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1902_s1.

Kano:1965:NPS

[246] Yutaka Kano. A new phase-space distribution function in the statistical theory of the electromagnetic field. Journal of Mathematical Physics, 6 (12):1913–1915, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i12/p1913_s1.

Katsura:1965:ESM

[247] Shigetoshi Katsura and Sakari Inawashiro. An exactly soluble model showing ferromagnetism. Journal of Mathematical Physics, 6(12):1916– 1922, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1916_s1.

Kumar:1965:EFN

[248] Kailash Kumar. Expansion of a function of noncommuting operators. Journal of Mathematical Physics, 6(12):1923–1927, December 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i12/p1923_
s1.

Kumar:1965:EE

[249] Kailash Kumar. On expanding the exponential. Journal of Mathematical Physics, 6(12):1928–1934, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i12/p1928_s1.

Maldonado:1965:NOP

[250] C. D. Maldonado. Note on orthogonal polynomials which are “invariant in form” to rotations of axes. Journal of Mathematical Physics, 6 (12):1935–1938, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i12/p1935_s1.

McCormick:1965:HSN

[251] N. J. McCormick and I. Kuscer. Half-space neutron transport with linearly anisotropic scattering. Journal of Mathematical Physics, 6 (12):1939–1945, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i12/p1939_s1.

Nayfeh:1965:EMT

[252] Ali Hasan Nayfeh. An expansion method for treating singular perturbation problems. Journal of Mathematical Physics, 6(12):1946–1951, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1946_s1.

OConnell:1965:GSM

[253] R. F. O’Connell and D. R. Tompkins. Generalized solutions for massless free fields and consequent generalized conservation laws. Journal of Mathematical Physics, 6(12):1952–1954, December 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i12/p1952_s1.

Petersson:1965:ROL

[254] Bengt Petersson. Reduction of a one-loop Feynman diagram with n vertices in m-dimensional Lorentz space. Journal of Mathematical Physics, 6(12):1955–1959, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i12/p1955_s1.

Polkinghorne:1965:FGC

[255] J. C. Polkinghorne. The Froissart–Gribov continuation and Reggeon unitarity conditions. I. Journal of Mathematical Physics, 6(12):1960– 1966, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1960_s1.

Urban:1965:CSC

[256] Eugene W. Urban. Critical störmer conditions in quadrupole and double ring-current fields. Journal of Mathematical Physics, 6(12):1966–1975, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v6/i12/p1966_s1.

Teitler:1965:ISC

[257] S. Teitler. Isotopic space, complex conjugation, and U_{3} symmetry for particles. Journal of Mathematical Physics, 6(12):1976–1979, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v6/i12/p1976_s1.

Weber:1965:SMP

[258] T. A. Weber and S. A. Williams. Spin-matrix polynomials and the rotation operator for arbitrary spin. Journal of Mathematical Physics, 6 (12):1980–1983, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v6/i12/p1980_s1.

Willis:1965:MIR

[259] Charles R. Willis. A model of interacting radiation and matter. II. Journal of Mathematical Physics, 6(12):1984–1994, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v6/i12/p1984_s1.

Wyman:1965:NFE

[260] M. Wyman and R. Trollope. Null fields in Einstein–Maxwell field theory. Journal of Mathematical Physics, 6(12):1995–2007, December 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i12/p1995_
s1.

Yeh:1965:PMD

[261] C. Yeh. Perturbation method in the diffraction of electromagnetic waves by arbitrarily shaped penetrable obstacles. Journal of Mathematical Physics, 6(12):2008–2013, December 1965. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v6/i12/p2008_s1.

Coleman:1965:ECG

[262] Sidney Coleman. Erratum: The Clebsch–Gordan series for SU(3). Journal of Mathematical Physics, 6(12):2014, December 1965. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v6/i12/p2014_s1. See [4].

Moses:1965:ETL

[263] H. E. Moses. Erratum: “Transformation from a Linear Momentum to an Angular Momentum Basis for Particles of Zero Mass and Finite Spin”. Journal of Mathematical Physics, 6(12):2014, December 1965. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v6/i12/p2014_

s2. See [112].

Aly:1966:SSP

[264] H. H. Aly and H. J. W. Müller. Scattering by the singular potential r^{−4}. Journal of Mathematical Physics, 7(1):1–9, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p1_s1.

Bardakci:1966:LARa

[265] K. Bardakci and B. Schroer. Local approximation in renormalizable and nonrenormalizable theories. I. Journal of Mathematical Physics, 7(1): 10–15, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p10_s1.

Bardakci:1966:LARb

[266] K. Bardakci and B. Schroer. Local approximations in renormalizable and nonrenormalizable theories. II. Journal of Mathematical Physics, 7(1): 16–26, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p16_s1.

Chan:1966:SED

[267] Yau W. Chan. On series expansion of a determinant and solution of the secular equation. Journal of Mathematical Physics, 7(1):27–34, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i1/p27_s1.

Hemmer:1966:DZG

[268] P. C. Hemmer, E. Hiis Hauge, and J. O. Aasen. Distribution of zeros of the grand partition function. Journal of Mathematical Physics, 7(1): 35–39, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p35_s1.

Chisholm:1966:ASP

[269] J. S. R. Chisholm. Approximation by sequences of padé approximants in regions of meromorphy. Journal of Mathematical Physics, 7(1):39– 44, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p39_s1.

Das:1966:CSTa

[270] A. Das. Complex space–time and classical field theory. I. Journal of Mathematical Physics, 7(1):45–51, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i1/p45_s1.

Das:1966:QCS

[271] A. Das. The quantized complex space–time and quantum theory of free fields. II. Journal of Mathematical Physics, 7(1):52–60, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i1/p52_s1.

Das:1966:CSTb

[272] A. Das. Complex space–time and geometrization of electromagnetism. III. Journal of Mathematical Physics, 7(1):61–63, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p61_s1.

Dilley:1966:ABS

[273] J. Dilley and A. O. Barut. Asymptotic behavior of the scattering amplitude in the left half-plane of complex angular momentum. Journal of Mathematical Physics, 7(1):64–65, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i1/p64_s1.

Pool:1966:MAW

[274] James C. T. Pool. Mathematical aspects of the Weyl correspondence. Journal of Mathematical Physics, 7(1):66–76, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p66_s1.

Fleming:1966:PRD

[275] Gordon N. Fleming. Physical region decoupling procedure for coupled channel scattering amplitudes. Journal of Mathematical Physics, 7(1): 76–80, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p76_s1.

Hurst:1966:RBO

[276] C. A. Hurst. Relation between the Onsager and Pfaffian methods for solving the Ising problem. II. The general lattice. Journal of Mathematical Physics, 7(1):81–87, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i1/p81_s1.

Jaksic:1966:ABP

[277] B. Jaksić and N. Limić. Asymptotic behavior of phase shift in λ for strongly singular potentials. Journal of Mathematical Physics, 7(1):88– 97, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p88_s1.

Lebowitz:1966:RTV

[278] J. L. Lebowitz and O. Penrose. Rigorous treatment of the van der Waals– Maxwell theory of the liquid-vapor transition. Journal of Mathematical Physics, 7(1):98–113, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i1/p98_s1.

Mariwalla:1966:RDS

[279] K. H. Mariwalla. Representation of discrete symmetry operators. Journal of Mathematical Physics, 7(1):114–122, January 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i1/p114_s1.

McGuire:1966:IFO

[280] J. B. McGuire. Interacting fermions in one dimension. II. Attractive potential. Journal of Mathematical Physics, 7(1):123–132, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i1/p123_s1.

Mehta:1966:STU

[281] C. L. Mehta, E. Wolf, and A. P. Balachandran. Some theorems on the unimodular complex degree of optical coherence. Journal of Mathematical Physics, 7(1):133–138, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i1/p133_s1.

Miasek:1966:AHM

[282] Maria Miasek. Application of Houston’s method to the sum of plane waves over the Brillouin zone. I. Simple-cubic and face-centered-cubic lattices. Journal of Mathematical Physics, 7(1):139–147, January 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p139_s1.

Lopuszanski:1966:QFT

[283] Jan Lopuszanski and Helmut Reeh. On a quantum field theory with degenerate vacuum and a special type of symmetry. Journal of Mathematical Physics, 7(1):148–154, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i1/p148_s1.

Misra:1966:SES

[284] M. Misra. Some exact solutions of the field equations of general relativity. Journal of Mathematical Physics, 7(1):155–158, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p155_s1.

Scarfone:1966:LSZ

[285] Leonard M. Scarfone. Lehmann–Symanzik–Zimmermann formalism in a model field theory. Journal of Mathematical Physics, 7(1):159–162, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i1/p159_s1.

Shelupsky:1966:TDB

[286] David Shelupsky. Translations of the discrete Bose–Einstein operators. Journal of Mathematical Physics, 7(1):163–166, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i1/p163_s1.

Shin:1966:STTa

[287] Ernest E. H. Shin. Space–time theory of elementary particles. I. Born reciprocity principle. Journal of Mathematical Physics, 7(1):167–173, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i1/p167_s1.

Shin:1966:STTb

[288] Ernest E. H. Shin. Space–time theory of elementary particles. II. Complex Minkowski space formalism. Journal of Mathematical Physics, 7(1):174–180, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i1/p174_s1.

Taylor:1966:SIM

[289] John R. Taylor. Spin and isospin in S-matrix theory. Journal of Mathematical Physics, 7(1):181–196, January 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i1/p181_s1.

Ahmavaara:1966:SSFa

[290] Yrjö Ahmavaara. Structure of space and the formalism of relativistic quantum theory. III. Journal of Mathematical Physics, 7(2):197–201, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i2/p197_s1.

Ahmavaara:1966:SSFb

[291] Yrjö Ahmavaara. Structure of space and the formalism of relativistic quantum theory. IV. Journal of Mathematical Physics, 7(2):201–204, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i2/p201_s1.

Alcaras:1966:AGM

[292] J. A. Castilho Alcarás, V. C. Aguilera Navarro, and P. Leal Ferreira. Applications of the Gel’fand–Moshinsky bases in unitary symmetry and its breaking. Journal of Mathematical Physics, 7(2):205–214, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v7/i2/p205_s1. See erratum [546].

Alessandrini:1966:FTE

[293] Victor A. Alessandrini. Faddeev-type equations for the four-body problem. Journal of Mathematical Physics, 7(2):215–220, February 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i2/p215_s1.

Aronson:1966:TMM

[294] Raphael Aronson and David Leon Yarmush. Transfer-matrix method for gamma-ray and neutron penetration. Journal of Mathematical Physics, 7(2):221–237, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i2/p221_s1.

Brill:1966:CFG

[295] Dieter R. Brill and Jeffrey M. Cohen. Cartan frames and the general relativistic Dirac equation. Journal of Mathematical Physics, 7(2):238– 243, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i2/p238_s1.

Cohen:1966:ETF

[296] Leon Cohen. Expansion theorem for functions of operators. Journal of Mathematical Physics, 7(2):244–245, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i2/p244_s1.

Cohen:1966:NTEa

[297] H. Cohen and C. N. DeSilva. Nonlinear theory of elastic surfaces. Journal of Mathematical Physics, 7(2):246–253, February 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i2/p246_s1.

Garcia-Colin:1966:GCU

[298] Leopoldo S. García-Colín and Asdrúbal Flores. The generalization of Choh–Uhlenbeck’s method in the kinetic theory of dense gases. Journal of Mathematical Physics, 7(2):254–259, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i2/p254_s1.

Fisher:1966:SMP

[299] Michael E. Fisher and David Ruelle. The stability of many-particle systems. Journal of Mathematical Physics, 7(2):260–270, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i2/p260_s1.

Guenin:1966:AFQ

[300] Marcel Guenin. Axiomatic foundations of quantum theories. Journal of Mathematical Physics, 7(2):271–282, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i2/p271_s1.

Harari:1966:RDP

[301] H. Harari. On the reduction of direct products of the irreducible representations of SU(n). Journal of Mathematical Physics, 7(2):283–285, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i2/p283_s1.

Hartwig:1966:MPC

[302] Robert E. Hartwig. Monomer pair correlations. Journal of Mathematical Physics, 7(2):286–299, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v7/i2/p286_s1.

Hunziker:1966:STB

[303] W. Hunziker. On the space–time behavior of schrödinger wavefunctions. Journal of Mathematical Physics, 7(2):300–304, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i2/p300_s1.

Hurst:1966:NAI

[304] C. A. Hurst. New approach to the Ising problem. Journal of Mathematical Physics, 7(2):305–310, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i2/p305_s1.

Jones:1966:NSN

[305] C. Edward Jones and George Tiktopoulos. Numerical solution of non-Fredholm (singular) integral equations by matrix inversion. Journal of Mathematical Physics, 7(2):311–315, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i2/p311_s1.

Lanczos:1966:MLP

[306] C. Lanczos. Metrical lattice and the problem of electricity. Journal of Mathematical Physics, 7(2):316–324, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i2/p316_s1.

Livingston:1966:MQE

[307] Peter M. Livingston. Methods of quadrature for Euler transform integrals. Journal of Mathematical Physics, 7(2):325–333, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i2/p325_s1.

Melvin:1966:OMU

[308] M. A. Melvin and J. S. Wallingford. Orbits in a magnetic universe. Journal of Mathematical Physics, 7(2):333–340, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i2/p333_s1.

Mehta:1966:CFE

[309] M. L. Mehta and P. K. Srivastava. Correlation functions for eigenvalues of real quaternian matrices. Journal of Mathematical Physics, 7(2):341– 344, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i2/p341_s1.

Mendelson:1966:FOS

[310] M. R. Mendelson. First-order solutions of the one-speed transport equation in plane multiregion geometries. Journal of Mathematical Physics, 7(2):345–355, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i2/p345_s1.

Pagnamenta:1966:BSU

[311] Antonio Pagnamenta. V Θ-bound state and uniqueness in the threeparticle sector of the Lee model. Journal of Mathematical Physics, 7(2):356–365, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i2/p356_s1.

Pauri:1966:CRL

[312] M. Pauri and G. M. Prosperi. Canonical realizations of Lie symmetry groups. Journal of Mathematical Physics, 7(2):366–375, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i2/p366_s1.

Pugh:1966:OTI

[313] Robert E. Pugh. S-operator theory. II. Quantum electrodynamics. Journal of Mathematical Physics, 7(2):376–384, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i2/p376_s1.

Rivers:1966:GAM

[314] R. J. Rivers. Gauge approximations in mesodynamics. Journal of Mathematical Physics, 7(2):385–394, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i2/p385_s1.

Roskies:1966:ISM

[315] Ralph Roskies. Internal symmetry and mass splitting. Journal of Mathematical Physics, 7(2):395–398, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i2/p395_s1.

Waters:1966:MSR

[316] John Waters and Joseph Ford. A method of solution for resonant nonlinear coupled oscillator systems. Journal of Mathematical Physics, 7(2):399–403, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i2/p399_s1.

Willis:1966:MIR

[317] Charles R. Willis. Model of interacting radiation and matter. III. Multimode gas lasers. Journal of Mathematical Physics, 7(2):404–412, February 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i2/p404_s1.

Bazley:1966:EBA

[318] Norman W. Bazley and David W. Fox. Error bounds for approximations to expectation values of unbounded operators. Journal of Mathematical Physics, 7(3):413–416, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i3/p413_s1.

Brown:1966:CSU

[319] Edmund H. Brown. On the complex structure of the universe. Journal of Mathematical Physics, 7(3):417–425, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i3/p417_s1.

Chakrabarti:1966:BPG

[320] A. Chakrabarti. On the L–S basis of the Poincaré group for the case of nonzero rest mass. Journal of Mathematical Physics, 7(3):426– 430, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p426_s1.

Gilbert:1966:NMT

[321] R. P. Gilbert and S. Y. Shieh. A new method in the theory of potential scattering. Journal of Mathematical Physics, 7(3):431–433, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i3/p431_s1.

Goldberg:1966:CRR

[322] Haim Goldberg. On the coupling and recoupling of relativistic states. Journal of Mathematical Physics, 7(3):434–442, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i3/p434_s1.

Kantowski:1966:SSH

[323] R. Kantowski and R. K. Sachs. Some spatially homogeneous anisotropic relativistic cosmological models. Journal of Mathematical Physics, 7(3): 443–446, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p443_s1.

Kaufman:1966:SFM

[324] Bruria Kaufman. Special functions of mathematical physics from the viewpoint of Lie algebra. Journal of Mathematical Physics, 7(3):447– 457, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p447_s1.

Lamb:1966:MME

[325] D. J. Lamb. Multipole moments in Einstein’s gravitational theory. Journal of Mathematical Physics, 7(3):458–463, March 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i3/p458_s1.

Moyal:1966:GTF

[326] J. E. Moyal. A general theory of first-passage distributions in transport and multiplicative processes. Journal of Mathematical Physics, 7(3): 464–473, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p464_s1.

Payne:1966:ASH

[327] Herbert Payne. Approximate solution of Hill’s equation. Journal of Mathematical Physics, 7(3):474–478, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i3/p474_s1.

Penney:1966:GCS

[328] R. Penney. Geometrization of a complex scalar field. I. Algebra. Journal of Mathematical Physics, 7(3):479–481, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i3/p479_s1.

Rohrlich:1966:FDC

[329] F. Rohrlich and M. Wilner. Functional differential calculus of operators. II. Journal of Mathematical Physics, 7(3):482–493, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i3/p482_s1.

Schmidt:1966:MOC

[330] Helmut Schmidt. Model of an oscillating cosmos which rejuvenates during contraction. Journal of Mathematical Physics, 7(3):494–509, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i3/p494_s1.

Symanzik:1966:EQF

[331] K. Symanzik. Euclidean quantum field theory. I. Equations for a scalar model. Journal of Mathematical Physics, 7(3):510–525, March 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i3/p510_s1.

Tao:1966:VPE

[332] L. N. Tao. On variational principles for electromagnetic theory. Journal of Mathematical Physics, 7(3):526–530, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i3/p526_s1.

Thompson:1966:PTT

[333] Colin J. Thompson. Phase transition of a two-dimensional continuumising model. Journal of Mathematical Physics, 7(3):531–534, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i3/p531_s1.

Unti:1966:TGF

[334] T. W. J. Unti and R. J. Torrence. Theorem on gravitational fields with geodesic rays. Journal of Mathematical Physics, 7(3):535–538, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i3/p535_s1.

Ruedenberg:1966:OIB

[335] Klaus Ruedenberg, Kiyosi O-Ohata, and D. G. Wilson. Overlap integrals between atomic orbitals. Journal of Mathematical Physics, 7(3):539– 546, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p539_s1.

O-Ohata:1966:TCC

[336] Kiyosi O-Ohata and Klaus Ruedenberg. Two-center Coulomb integrals between atomic orbitals. Journal of Mathematical Physics, 7(3):547– 559, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p547_s1.

Tarski:1966:SBN

[337] Jan Tarski. Solution of the Bloch–Nordsieck model. Journal of Mathematical Physics, 7(3):560–573, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i3/p560_s1.

Dubin:1966:IMR

[338] Daniel A. Dubin and Jan Tarski. Indefinite metric resulting from regularization in the infrared region. Journal of Mathematical Physics, 7(3): 574–577, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i3/p574_s1.

Eger:1966:PTH

[339] F. M. Eger and E. P. Gross. Point transformations and the hardsphere Bose gas. Journal of Mathematical Physics, 7(3):578–582, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i3/p578_s1.

Fried:1966:FFL

[340] H. M. Fried. Functional formulation of the Lee model. Journal of Mathematical Physics, 7(3):583–588, March 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i3/p583_s1.

Aronson:1966:PMC

[341] Raphael Aronson. A perturbation method for the classical timedependent pair correlation function. Journal of Mathematical Physics, 7(4):589–602, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i4/p589_s1.

Boyce:1966:ARC

[342] J. F. Boyce. Asymptotic renormalizability conditions on field propagators. Journal of Mathematical Physics, 7(4):603–607, April 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p603_s1.

Collinson:1966:EST

[343] C. D. Collinson. Empty space–times of embedding class two. Journal of Mathematical Physics, 7(4):608–611, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i4/p608_s1.

Derome:1966:SPS

[344] Jean-Robert Derome. Symmetry properties of the 3j-symbols for an arbitrary group. Journal of Mathematical Physics, 7(4):612–615, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p612_s1.

DeWitt:1966:SMH

[345] Hugh E. DeWitt. Statistical mechanics of high-temperature quantum plasmas beyond the ring approximation. Journal of Mathematical Physics, 7(4):616–626, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i4/p616_s1.

Ford:1966:LFS

[346] William F. Ford. Limiting forms of the screened Coulomb T matrix. Journal of Mathematical Physics, 7(4):626–633, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p626_s1.

Granzow:1966:MTT

[347] Kenneth D. Granzow. Multipole theory in the time domain. Journal of Mathematical Physics, 7(4):634–640, April 1966. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i4/p634_s1. See erratum [547].

Gundzik:1966:CRI

[348] M. G. Gundzik. A continuous representation of an indefinite metric space. Journal of Mathematical Physics, 7(4):641–651, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p641_s1.

Islam:1966:LLC

[349] Jamal N. Islam. Leading Landau curves of a class of Feynman diagrams. Journal of Mathematical Physics, 7(4):652–660, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p652_s1.

Keller:1966:EDC

[350] Joseph B. Keller and Frank C. Karal. Effective dielectric constant, permeability, and conductivity of a random medium and the velocity and attenuation coefficient of coherent waves. Journal of Mathematical Physics, 7(4):661–670, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i4/p661_s1.

Kumar:1966:TTU

[351] Kailash Kumar. Talmi transformation for unequal-mass particles and related formulas. Journal of Mathematical Physics, 7(4):671–678, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p671_s1.

Minguzzi:1966:APS

[352] A. Minguzzi. Analyticity properties of the scattering amplitude. Journal of Mathematical Physics, 7(4):679–681, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i4/p679_s1.

Majumdar:1966:CBU

[353] Chanchal K. Majumdar. Continuity of bound and unbound states in a Fermi gas: A soluble example. Journal of Mathematical Physics, 7(4): 682–690, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i4/p682_s1.

Moshinsky:1966:GSI

[354] M. Moshinsky. Gelfand states and the irreducible representations of the symmetric group. Journal of Mathematical Physics, 7(4):691–698, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p691_s1.

Nakanishi:1966:PPV

[355] Noboru Nakanishi. Poles of the proper vertex function in the Bethe– Salpeter formalism. Journal of Mathematical Physics, 7(4):698–701, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p698_s1.

Olsson:1966:ACH

[356] Per O. M. Olsson. Analytic continuations of higher-order hypergeometric functions. Journal of Mathematical Physics, 7(4):702–710, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p702_s1.

Poshusta:1966:PAF

[357] R. D. Poshusta and F. A. Matsen. Permutation-algebraic formulation of spin-free transition density matrices. Journal of Mathematical Physics, 7(4):711–720, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i4/p711_s1.

Sasakawa:1966:SSE

[358] Tatuya Sasakawa. Solution of schrödinger equation involving time. Journal of Mathematical Physics, 7(4):721–741, April 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i4/p721_s1.

Siegert:1966:GVW

[359] Arnold J. F. Siegert and Emmanuel Meeron. Generalizations of the virial and Wall theorems in classical statistical mechanics. Journal of Mathematical Physics, 7(4):741–750, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i4/p741_s1.

Szekeres:1966:PGF

[360] Peter Szekeres. On the propagation of gravitational fields in matter. Journal of Mathematical Physics, 7(4):751–761, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p751_s1.

Ullah:1966:ACH

[361] Nazakat Ullah. Asymptotic correlations of the Hamiltonian matrix elements. Journal of Mathematical Physics, 7(4):762–764, April 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i4/p762_s1.

Zmuidzinas:1966:URL

[362] Jonas Stasys Zmuidzinas. Unitary representations of the Lorentz group on 4-vector manifolds. Journal of Mathematical Physics, 7(4):764–780, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i4/p764_s1.

Cohen:1966:GPS

[363] Leon Cohen. Generalized phase-space distribution functions. Journal of Mathematical Physics, 7(5):781–786, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i5/p781_s1.

Coleman:1966:IVI

[364] Sidney Coleman. The invariance of the vacuum is the invariance of the world. Journal of Mathematical Physics, 7(5):787, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p787_s1.

Contopoulos:1966:AIT

[365] George Contopoulos. Adiabatic invariants and the “third” integral. Journal of Mathematical Physics, 7(5):788–797, May 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i5/p788_s1.

DeWitt:1966:FMI

[366] Hugh De Witt and Paul Fishbane. Failure of the Mayer irreducible cluster theorem with wave mechanics. Journal of Mathematical Physics, 7(5): 798–801, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i5/p798_s1.

Dollard:1966:ASS

[367] John Day Dollard. Adiabatic switching in the schrödinger theory of scattering. Journal of Mathematical Physics, 7(5):802–810, May 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p802_s1.

Fairlie:1966:SCM

[368] D. B. Fairlie. Structure of the crossing matrix for arbitrary internal symmetry groups. Journal of Mathematical Physics, 7(5):811–815, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i5/p811_s1.

Fischer:1966:GSF

[369] J. Fischer, J. Niederle, and R. Raczka. Generalized spherical functions for the noncompact rotation groups. Journal of Mathematical Physics, 7(5): 816–821, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i5/p816_s1.

Klauder:1966:DPR

[370] J. R. Klauder, J. McKenna, and E. J. Woods. Direct-product representations of the canonical commutation relations. Journal of Mathematical Physics, 7(5):822–828, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i5/p822_s1.

Marburger:1966:CEC

[371] J. H. Marburger. Conditions for the existence of closed solutions by the normal ordering method. Journal of Mathematical Physics, 7(5): 829–832, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i5/p829_s1.

Mika:1966:IVPa

[372] Janusz Mika. The initial-value problem for a slab. I. Anisotropic scattering. Journal of Mathematical Physics, 7(5):833–838, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p833_s1.

Mika:1966:IVPb

[373] Janusz Mika. The initial-value problem for a slab. II. Nonuniform slab. Journal of Mathematical Physics, 7(5):839–843, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p839_s1.

Morrow:1966:PRM

[374] Richard A. Morrow. The physical regions of many-particle processes. Journal of Mathematical Physics, 7(5):844–848, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p844_s1.

Mukunda:1966:DOF

[375] N. Mukunda and T. F. Jordan. Determination of optical field correlations from photon counts. Journal of Mathematical Physics, 7(5): 849–853, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i5/p849_s1.

Murthy:1966:RRF

[376] M. V. Murthy. Ray representations of finite nonunitary groups. Journal of Mathematical Physics, 7(5):853–857, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i5/p853_s1.

Muzinich:1966:IUP

[377] I. J. Muzinich. The implications of unitarity on possible higher symmetries. Journal of Mathematical Physics, 7(5):858–862, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p858_s1.

Newman:1966:NBM

[378] E. T. Newman and R. Penrose. Note on the Bondi–Metzner–Sachs group. Journal of Mathematical Physics, 7(5):863–870, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p863_s1.

Remler:1966:SO

[379] E. A. Remler. Spectrum of the S operator. Journal of Mathematical Physics, 7(5):871–874, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i5/p871_s1.

Rosner:1966:EMP

[380] Jonathan L. Rosner. Exchange of massive particles in the Bethe–Salpeter equation. Journal of Mathematical Physics, 7(5):875–884, May 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p875_s1.

Uhlenbrock:1966:PQS

[381] D. A. Uhlenbrock. Properties of quantum statistical expectation values. Journal of Mathematical Physics, 7(5):885–893, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p885_s1.

Wahlquist:1966:RME

[382] Hugo D. Wahlquist and Frank B. Estabrook. Rigid motions in Einstein spaces. Journal of Mathematical Physics, 7(5):894–905, May 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p894_s1.

Fonda:1966:DPM

[383] Luciano Fonda and Gian’ Carlo Ghirardi. Determination of the parameters of model field theories from phase shifts and bound-states energies. Journal of Mathematical Physics, 7(5):906–915, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i5/p906_s1.

Emch:1966:GIF

[384] Gérard Emch and Marcel Guenin. Gauge invariant formulation of the BCS model. Journal of Mathematical Physics, 7(5):915–921, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i5/p915_s1.

Piccirelli:1966:SPL

[385] Robert A. Piccirelli. Some properties of the long-time values of the probability densities for moderately dense gases. Journal of Mathematical Physics, 7(5):922–934, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i5/p922_s1.

Rudra:1966:ITO

[386] P. Rudra. On irreducible tensor operators for finite groups. Journal of Mathematical Physics, 7(5):935–937, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i5/p935_s1.

Eyges:1966:IMQ

[387] Leonard Eyges. Improved method for quantum mechanical threebody problems. II. Repulsive potentials. Journal of Mathematical Physics, 7(5):938–948, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i5/p938_s1.

Chakrabarti:1966:UUR

[388] A. Chakrabarti. Unified unitary representation of the Poincaré group for particles of zero and positive rest mass. Journal of Mathematical Physics, 7(6):949–955, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i6/p949_s1.

Bebb:1966:GDC

[389] H. Barry Bebb. Generalized discrete-continuum radial integrals with Coulomb functions. Journal of Mathematical Physics, 7(6):955–959, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i6/p955_s1.

Cohen:1966:NTEb

[390] H. Cohen and C. N. DeSilva. Nonlinear theory of elastic directed surfaces. Journal of Mathematical Physics, 7(6):960–966, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i6/p960_s1.

Currie:1966:EPH

[391] Douglas G. Currie and Eugene J. Saletan. q-equivalent particle hamiltonians. I. The classical one-dimensional case. Journal of Mathematical Physics, 7(6):967–974, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i6/p967_s1.

Davis:1966:DCF

[392] W. R. Davis and M. K. Moss. On the development of the covariant formulation of the conservation laws of general relativity. Journal of Mathematical Physics, 7(6):975–976, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i6/p975_s1.

Falk:1966:ULB

[393] Harold Falk. Upper and lower bounds for canonical ensemble averages. Journal of Mathematical Physics, 7(6):977–979, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i6/p977_s1.

George:1966:FDR

[394] Claude George and Monique Lévy-Nahas. Finite-dimensional representations of some non-semisimple Lie algebras. Journal of Mathematical Physics, 7(6):980–988, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i6/p980_s1.

Hagen:1966:HOP

[395] C. R. Hagen. Higher-order poles in the S-matrix. Journal of Mathematical Physics, 7(6):989–990, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i6/p989_s1.

Liboff:1966:DCS

[396] Richard L. Liboff. Dynamics of certain spherical charge distributions. Journal of Mathematical Physics, 7(6):991–997, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i6/p991_s1.

Kummer:1966:MGC

[397] Martin Kummer. The most general Clebsch–Gordan coefficients of the universal covering group of the inhomogeneous Lorentz group. Journal of Mathematical Physics, 7(6):997–1015, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i6/p997_s1.

Lieb:1966:QME

[398] Elliott Lieb. Quantum-mechanical extension of the Lebowitz–Penrose theorem on the van der Waals theory. Journal of Mathematical Physics, 7(6):1016–1024, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i6/p1016_s1.

Maitra:1966:SDF

[399] S. C. Maitra. Stationary dust-filled cosmological solution with Λ = 0 and without closed timelike lines. Journal of Mathematical Physics, 7(6):

1025–1030, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i6/p1025_s1.

McGuinness:1966:NDE

[400] D. J. McGuinness. Nonhomogeneous differential equation with a secondorder turning point. Journal of Mathematical Physics, 7(6):1030–1037, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i6/p1030_s1.

Mermin:1966:SSE

[401] N. David Mermin. A short simple evaluation of expressions of the Debye– Waller form. Journal of Mathematical Physics, 7(6):1038, June 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i6/p1038_
s1.

Morita:1966:DGB

[402] Tohru Morita. Derivation of the generalized Boltzmann equation in quantum statistical mechanics. Journal of Mathematical Physics, 7(6):1039– 1045, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i6/p1039_s1.

Nussinov:1966:TCG

[403] S. Nussinov. Theorem on the Clebsch–Gordan series in SU(n). Journal of Mathematical Physics, 7(6):1046–1048, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i6/p1046_s1.

Parks:1966:ACR

[404] W. F. Parks. On alternate commutation relations. Journal of Mathematical Physics, 7(6):1049–1053, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i6/p1049_s1.

Prugovecki:1966:AAFa

[405] Eduard Prugovecki. An axiomatic approach to the formalism of quantum mechanics. I. Journal of Mathematical Physics, 7(6):1054–1069, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i6/p1054_s1.

Prugovecki:1966:AAFb

[406] Eduard Prugovecki. An axiomatic approach to the formalism of quantum mechanics. II. Journal of Mathematical Physics, 7(6):1070–1096, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i6/p1070_s1.

Roberts:1966:DBK

[407] J. E. Roberts. The Dirac bra and ket formalism. Journal of Mathematical Physics, 7(6):1097–1104, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i6/p1097_s1.

Sharp:1966:PBI

[408] R. T. Sharp and Hans Von Baeyer. Polynomial bases and isoscalar factors for SU(3). Journal of Mathematical Physics, 7(6):1105–1122, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i6/p1105_s1.

Stephenson:1966:IMS

[409] John Stephenson. Ising model spin correlations on the triangular lattice. II. Fourth-order correlations. Journal of Mathematical Physics, 7(6): 1123–1132, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i6/p1123_s1.

Yndurain:1966:MFC

[410] F. J. Ynduráin. S-matrix formalism, charge renormalization, and definition of the Hamiltonian in a simple field-theoretical model. Journal of Mathematical Physics, 7(6):1133–1136, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i6/p1133_s1.

Zipoy:1966:TSS

[411] David M. Zipoy. Topology of some spheroidal metrics. Journal of Mathematical Physics, 7(6):1137–1143, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i6/p1137_s1.

Backus:1966:NPI

[412] G. E. Backus and F. Gilbert. Note on product integrals. Journal of Mathematical Physics, 7(6):1144, June 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i6/p1144_s1.

Bradley:1966:SGS

[413] C. J. Bradley. Space groups and selection rules. Journal of Mathematical Physics, 7(7):1145–1152, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1145_s1.

Chappell:1966:KEE

[414] Willard R. Chappell. Kinetic equation for the electron-phonon gas. Journal of Mathematical Physics, 7(7):1153–1157, July 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1153_s1.

Chow:1966:MFG

[415] Yutze Chow. Majorization of Feynman graphs. Journal of Mathematical Physics, 7(7):1158–1170, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1158_s1.

Cocke:1966:SCC

[416] W. J. Cocke. Some collapsing cylinders and their exterior vacuum metrics in general relativity. Journal of Mathematical Physics, 7(7):1171– 1178, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i7/p1171_s1.

Droz-Vincent:1966:CCL

[417] Philippe H. Droz-Vincent. Covariant conservation laws for general relativistic exploding matter. Journal of Mathematical Physics, 7(7):1179– 1184, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i7/p1179_s1.

Ellis:1966:EFM

[418] J. R. Ellis. Electromagnetic fields of moving dipoles and multipoles. Journal of Mathematical Physics, 7(7):1185–1197, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1185_s1.

Emch:1966:NMM

[419] Gérard G. Emch. Non-Markovian model for the approach to equilibrium. Journal of Mathematical Physics, 7(7):1198–1206, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1198_s1.

Ferreira:1966:PML

[420] Erasmo M. Ferreira and Antonio F. F. Teixeira. On the poles of the S matrix for long-range potentials. Journal of Mathematical Physics, 7(7): 1207–1217, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i7/p1207_s1.

Finkelstein:1966:K

[421] David Finkelstein. Kinks. Journal of Mathematical Physics, 7(7):1218– 1225, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i7/p1218_s1.

Halpern:1966:SVS

[422] Martin B. Halpern. Solution of very singular N/D equations — an approach to nonrenormalizable field theory. Journal of Mathematical Physics, 7(7):1226–1250, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1226_s1.

Jaffe:1966:WPF

[423] Arthur Jaffe. Wick polynomials at a fixed time. Journal of Mathematical Physics, 7(7):1250–1255, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1250_s1.

Kruskal:1966:EIC

[424] Martin D. Kruskal and Norman J. Zabusky. Exact invariants for a class of nonlinear wave equations. Journal of Mathematical Physics, 7(7): 1256–1267, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i7/p1256_s1.

Lenard:1966:ODI

[425] A. Lenard. One-dimensional impenetrable bosons in thermal equilibrium. Journal of Mathematical Physics, 7(7):1268–1272, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1268_s1.

Roman:1966:CFT

[426] P. Roman and J. J. Aghassi. Classical field theory and gravitation in a de Sitter world. Journal of Mathematical Physics, 7(7):1273–1283, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i7/p1273_s1.

Rosen:1966:UIR

[427] Gerald Rosen. Unitary irreducible representations of SL(3, R). Journal of Mathematical Physics, 7(7):1284–1294, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i7/p1284_s1.

Schmidt:1966:SAX

[428] Paul W. Schmidt. Small-angle X-ray scattering from rods and platelets. Journal of Mathematical Physics, 7(7):1295–1300, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1295_s1.

Sesma:1966:CT

[429] J. Sesma. On the Chakrabarti transformation. Journal of Mathematical Physics, 7(7):1300–1302, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1300_s1.

Shepelavey:1966:GCCa

[430] Bohdan Shepelavey. Global covariant conservation laws in Riemannian spaces. I. Journal of Mathematical Physics, 7(7):1303–1309, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i7/p1303_s1.

Shepelavey:1966:GCCb

[431] Bohdan Shepelavey. Global covariant conservation laws in Riemannian spaces. II. Journal of Mathematical Physics, 7(7):1310–1320, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i7/p1310_s1.

Stachel:1966:CGN

[432] John J. Stachel. Cylindrical gravitational news. Journal of Mathematical Physics, 7(7):1321–1331, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i7/p1321_s1.

Taylor:1966:MES

[433] C. D. Taylor. On the motion of electrons scattered from an infinite slab. Journal of Mathematical Physics, 7(7):1332–1334, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1332_s1.

Tjon:1966:NQT

[434] J. A. Tjon. A note on the quantum theory of electrical conductivity. Journal of Mathematical Physics, 7(7):1335–1336, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1335_s1.

Wolfson:1966:CFE

[435] Seymour J. Wolfson. A closed form for Elsasser integrals. Journal of Mathematical Physics, 7(7):1337–1339, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i7/p1337_s1.

Glasser:1966:ENE

[436] M. L. Glasser. Erratum: “Note on the Evaluation of Some Fermi Integrals”. Journal of Mathematical Physics, 7(7):1340, July 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1340_s1. See

[3].

Taylor:1966:EPS

[437] John G. Taylor. Erratum: “On Polynomial Systems in a Banach Ring”. Journal of Mathematical Physics, 7(7):1340, July 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i7/p1340_s2. See [142].

Arens:1966:LLN

[438] Richard Arens. Laws like Newton’s. Journal of Mathematical Physics, 7(8):1341–1348, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i8/p1341_s1.

Bertotti:1966:DMS

[439] Bruno Bertotti. Doppler measurement of space–time curvature. Journal of Mathematical Physics, 7(8):1349–1350, August 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i8/p1349_s1.

Bronzan:1966:TMS

[440] J. B. Bronzan. A two-meson solution of the charged scalar static model. Journal of Mathematical Physics, 7(8):1351–1361, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i8/p1351_s1.

Cantoni:1966:CRG

[441] Vittorio Cantoni. A class of representations of the generalized Bondi– Metzner group. Journal of Mathematical Physics, 7(8):1361–1364, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1361_s1.

Chen:1966:RCQ

[442] T. W. Chen, F. Rohrlich, and M. Wilner. Role of causality in quantum field theory and the dynamical postulate. Journal of Mathematical Physics, 7(8):1365–1371, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i8/p1365_s1.

Riccia:1966:WMC

[443] Giacomo Della Riccia and Norbert Wiener. Wave mechanics in classical phase space, Brownian motion, and quantum theory. Journal of Mathematical Physics, 7(8):1372–1383, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i8/p1372_s1.

DesCloizeaux:1966:ALM

[444] Jacques Des Cloizeaux and Michel Gaudin. Anisotropic linear magnetic chain. Journal of Mathematical Physics, 7(8):1384–1400, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1384_s1.

Dulock:1966:DCM

[445] Victor A. Dulock and Harold V. McIntosh. Degeneracy of cyclotron motion. Journal of Mathematical Physics, 7(8):1401–1412, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1401_s1.

Emch:1966:DSQ

[446] Gérard G. Emch. The definition of states in quantum statistical mechanics. Journal of Mathematical Physics, 7(8):1413–1420, August 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i8/p1413_
s1.

Frantz:1966:WPD

[447] Lee M. Frantz. Wave-packet derivation of field-theoretic scattering amplitude. Journal of Mathematical Physics, 7(8):1421–1435, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1421_s1.

Hardy:1966:CDF

[448] Robert J. Hardy. Concise derivation of the formulas for lattice thermal conductivity. Journal of Mathematical Physics, 7(8):1435–1438, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1435_s1.

Hardy:1966:HOC

[449] Robert J. Hardy and William C. Schieve. Higher-order corrections to the lattice thermal conductivity. Journal of Mathematical Physics, 7(8):1439–1452, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i8/p1439_s1.

Johnston:1966:GSH

[450] Tudor Wyatt Johnston. General spherical harmonic tensors in the Boltzmann equation. Journal of Mathematical Physics, 7(8):1453–1458, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1453_s1.

Karayianis:1966:ESS

[451] N. Karayianis, C. A. Morrison, and D. E. Wortman. Exact solutions for a semi-infinite square lattice gas. Journal of Mathematical Physics, 7(8):1458–1463, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i8/p1458_s1.

Landshoff:1966:ESM

[452] P. V. Landshoff and D. I. Olive. Extraction of singularities from the S matrix. Journal of Mathematical Physics, 7(8):1464–1477, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1464_s1.

Lurie:1966:GWI

[453] D. Lurié, Y. Takahashi, and H. Umezawa. Generalized Ward identity and unified treatment of conservation laws. Journal of Mathematical Physics, 7(8):1478–1483, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v7/i8/p1478_s1.

Nagle:1966:LSHa

[454] J. F. Nagle. Lattice statistics of hydrogen bonded crystals. I. The residual entropy of ice. Journal of Mathematical Physics, 7(8):1484–1491, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i8/p1484_s1.

Nagle:1966:LSHb

[455] J. F. Nagle. Lattice statistics of hydrogen bonded crystals. II. The Slater KDP model and the Rys F -model. Journal of Mathematical Physics, 7(8):1492–1496, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i8/p1492_s1.

Nickle:1966:NQT

[456] H. H. Nickle. Nonrelativistic quantum theory of an electron in an arbitrarily intense laser field. Journal of Mathematical Physics, 7(8):1497– 1506, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i8/p1497_s1. See erratum [763].

Ramanathan:1966:NCS

[457] G. V. Ramanathan. A new cluster scheme in statistical physics. Journal of Mathematical Physics, 7(8):1507–1514, August 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i8/p1507_s1.

Sabatier:1966:APP

[458] Pierre C. Sabatier. Asymptotic properties of the potentials in the inversescattering problem at fixed energy. Journal of Mathematical Physics, 7(8):1515–1531, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i8/p1515_s1.

Stell:1966:SIP

[459] G. Stell, J. L. Lebowitz, S. Baer, and W. Theumann. Separation of the interaction potential into two parts in statistical mechanics. II. Graph theory for lattice gases and spin systems with application to systems with long-range potentials. Journal of Mathematical Physics, 7(8):1532– 1547, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i8/p1532_s1.

Zaccaria:1966:TCG

[460] F. Zaccaria. A theorem on the Clebsch–Gordan series in A ( l ), B ( l ), C ( l ), and D ( l ). Journal of Mathematical Physics, 7(8):1548– 1551, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i8/p1548_s1.

Zwanzig:1966:PEM

[461] Robert Zwanzig. Probability and entropy of macroscopic fluctuations. Journal of Mathematical Physics, 7(8):1552–1556, August 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i8/p1552_s1.

Sykes:1966:LCS

[462] M. F. Sykes, J. W. Essam, B. R. Heap, and B. J. Hiley. Lattice constant systems and graph theory. Journal of Mathematical Physics, 7(9):1557– 1572, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i9/p1557_s1.

Essam:1966:PPL

[463] J. W. Essam and M. F. Sykes. Percolation processes. I. Low-density expansion for the mean number of clusters in a random mixture. Journal of Mathematical Physics, 7(9):1573–1581, September 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1573_s1.

Heap:1966:EHI

[464] B. R. Heap. The enumeration of homeomorphically irreducible star graphs. Journal of Mathematical Physics, 7(9):1582–1587, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i9/p1582_s1.

Nagle:1966:OIU

[465] John F. Nagle. On ordering and identifying undirected linear graphs. Journal of Mathematical Physics, 7(9):1588–1592, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1588_
s1.

Landshoff:1966:NTS

[466] P. V. Landshoff, D. I. Olive, and J. C. Polkinghorne. Normal thresholds in subenergy variables. Journal of Mathematical Physics, 7(9):1593–1599, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i9/p1593_s1.

Landshoff:1966:ICS

[467] P. V. Landshoff, D. I. Olive, and J. C. Polkinghorne. The icecreamcone singularity in S-matrix theory. Journal of Mathematical Physics, 7(9):1600–1606, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i9/p1600_s1.

Atkinson:1966:ESE

[468] D. Atkinson. Existence of solutions of N/D equations. Journal of Mathematical Physics, 7(9):1607–1612, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i9/p1607_s1.

Blood:1966:PDF

[469] F. A. Blood. The r-particle distribution function in classical physics. Journal of Mathematical Physics, 7(9):1613–1619, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1613_
s1.

Bruch:1966:DM

[470] L. W. Bruch. N degenerate modes. Journal of Mathematical Physics, 7(9):1620–1624, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i9/p1620_s1.

Farnsworth:1966:HDF

[471] David L. Farnsworth and Roy P. Kerr. Homogeneous dust-filled cosmological solutions. Journal of Mathematical Physics, 7(9):1625–1632, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i9/p1625_s1.

Finkelstein:1966:QST

[472] R. Finkelstein. On quantization in space with torsion. Journal of Mathematical Physics, 7(9):1632–1642, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i9/p1632_s1.

Fonda:1966:APR

[473] L. Fonda, G. C. Ghirardi, T. Weber, and A. Rimini. Analytic properties of resolvents and completeness. Journal of Mathematical Physics, 7(9):1643–1647, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i9/p1643_s1.

Giere:1966:RHW

[474] Albert C. Giere. Relativistic hydromagnetic waves and group velocity. Journal of Mathematical Physics, 7(9):1648–1656, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1648_
s1.

Licht:1966:LS

[475] A. L. Licht. Local states. Journal of Mathematical Physics, 7(9):1656– 1669, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i9/p1656_s1.

Nussinov:1966:APB

[476] S. Nussinov and J. Rosner. Absorptive parts and the Bethe–Salpeter equation for forward scattering. Journal of Mathematical Physics, 7 (9):1670–1679, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i9/p1670_s1.

Prugovecki:1966:FGQ

[477] Eduard Prugovecki. A formalism for generalized quantum mechanics. Journal of Mathematical Physics, 7(9):1680–1696, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1680_
s1.

Rohrlich:1966:OT

[478] F. Rohrlich and J. G. Wray. Ordering theorems. Journal of Mathematical Physics, 7(9):1697–1706, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i9/p1697_s1.

Scharfstein:1966:AAT

[479] H. Scharfstein. Algebraic aspects of trilinear commutation relations. Journal of Mathematical Physics, 7(9):1707–1719, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1707_
s1.

Taylor:1966:EFT

[480] J. G. Taylor. On the existence of field theory. I. The analytic approach. Journal of Mathematical Physics, 7(9):1720–1729, September 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1720_
s1.

Teitler:1966:SS

[481] S. Teitler. The structure of 4-spinors. Journal of Mathematical Physics, 7(9):1730–1738, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i9/p1730_s1.

Teitler:1966:LEU

[482] S. Teitler. Lorentz equivalence, unitary symmetry, and spin unitary symmetry. Journal of Mathematical Physics, 7(9):1739–1743, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i9/p1739_s1.

Ullman:1966:NSS

[483] Nelly Ullman and Robert Ullman. Numerical solutions of singular Fredholm equations. Journal of Mathematical Physics, 7(9):1743–1748, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i9/p1743_s1.

Boyling:1966:CVC

[484] J. B. Boyling. Construction of vanishing cycles for integrals over hyperspheres. Journal of Mathematical Physics, 7(10):1749–1763, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i10/p1749_s1.

Devooght:1966:VGF

[485] J. Devooght. Variation of Green’s functions. Journal of Mathematical Physics, 7(10):1764–1770, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v7/i10/p1764_s1.

Faulkner:1966:ECM

[486] James E. Faulkner. The existence of closed magnetic surfaces. Journal of Mathematical Physics, 7(10):1771–1775, October 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1771_s1.

Fisher:1966:DSP

[487] Michael E. Fisher. On the dimer solution of planar Ising models. Journal of Mathematical Physics, 7(10):1776–1781, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1776_s1.

Greenman:1966:HEB

[488] J. V. Greenman. High-energy behavior of Feynman integrals with spin. I. Journal of Mathematical Physics, 7(10):1782–1796, October 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1782_
s1.

Gruber:1966:AMD

[489] B. Gruber. On an algebraic method for the decomposition of direct products of representations of the groups A_{2}(SU(3)) and B_{2}(SO(5)). Journal of Mathematical Physics, 7(10):1797–1801, October 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1797_s1.

Katz:1966:SCS

[490] Amnon Katz. Scattering cross section for a beam of wave packets. Journal of Mathematical Physics, 7(10):1802–1805, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1802_s1.

Kobe:1966:LCT

[491] Donald H. Kobe. Linked cluster theorem and the Green’s function equations of motion for a many-fermion system. Journal of Mathematical Physics, 7(10):1806–1820, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v7/i10/p1806_s1.

Lopuszanski:1966:RRQ

[492] Jan Lopuszanski and Helmut Reeh. A remark on a reducible quantum field theory with a one-parameter symmetry group. Journal of Mathematical Physics, 7(10):1821–1823, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i10/p1821_s1.

Mehta:1966:CIUa

[493] M. L. Mehta. Classification of irreducible unitary representations of compact simple Lie groups. I. Journal of Mathematical Physics, 7(10):1824– 1832, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i10/p1824_s1.

Mehta:1966:CIUb

[494] M. L. Mehta and P. K. Srivastava. Classification of irreducible unitary representations of compact simple Lie groups. II. Journal of Mathematical Physics, 7(10):1833–1835, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i10/p1833_s1.

Misra:1966:TDF

[495] R. M. Misra and R. A. Singh. Three-dimensional formulation of gravitational null fields. I. Journal of Mathematical Physics, 7(10):1836– 1840, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i10/p1836_s1.

Nozawa:1966:BES

[496] Reikichi Nozawa. Bipolar expansion of screened Coulomb potentials, Helmholtz’ Solid harmonics, and their addition theorems. Journal of Mathematical Physics, 7(10):1841–1860, October 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i10/p1841_s1.

Raczka:1966:DDR

[497] R. Raczka, N. Limić, and J. Niederle. Discrete degenerate representations of noncompact rotation groups. I. Journal of Mathematical Physics, 7(10):1861–1876, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i10/p1861_s1.

Rubin:1966:MAF

[498] Robert J. Rubin and Pieter Ullersma. The momentum autocorrelation function in a Bernoulli chain. Journal of Mathematical Physics, 7(10):1877–1885, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i10/p1877_s1.

Santhanam:1966:SRC

[499] T. S. Santhanam. Some remarks on the construction of invariants of semisimple local Lie groups. Journal of Mathematical Physics, 7 (10):1886–1888, October 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i10/p1886_s1.

Abraham:1966:TPT

[500] C. Abraham. Transformation of Proca-type system of partial differential equations. Journal of Mathematical Physics, 7(11):1889–1890, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p1889_s1.

Armstrong:1966:REA

[501] Lloyd Armstrong. Relativistic effects in atomic fine structure. Journal of Mathematical Physics, 7(11):1891–1899, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i11/p1891_s1.

Baker:1966:VPS

[502] George A. Baker and Roy Chisholm. The validity of perturbation series with zero radius of convergence. Journal of Mathematical Physics, 7 (11):1900–1902, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i11/p1900_s1.

Barut:1966:UAR

[503] A. O. Barut and S. Komy. Unitary and antiunitary ray representations of the product of n commuting parity operators. Journal of Mathematical Physics, 7(11):1903–1907, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i11/p1903_s1.

Barut:1966:IRC

[504] A. O. Barut. On irreducible representation of a class of algebras and related projective representations. Journal of Mathematical Physics, 7 (11):1908–1910, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i11/p1908_s1.

Denman:1966:ICL

[505] Harry H. Denman. Invariance and conservation laws in classical mechanics. II. Journal of Mathematical Physics, 7(11):1910–1915, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p1910_s1.

DeYoung:1966:RSF

[506] David S. De Young. Relativistic spinor formulation of Stokes parameters with application to the inverse Compton effect. Journal of Mathematical Physics, 7(11):1916–1923, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i11/p1916_s1.

Flamand:1966:SSE

[507] G. Flamand. The solution of a Stark-effect model as a dynamical group calculation. Journal of Mathematical Physics, 7(11):1924–1931, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p1924_s1.

Flato:1966:CBE

[508] Moshé Flato and Daniel Sternheimer. On the connection between external and internal symmetries of strongly interacting particles. Journal of Mathematical Physics, 7(11):1932–1958, November 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i11/p1932_s1.

Fleming:1966:MCD

[509] Gordon N. Fleming. A manifestly covariant description of arbitrary dynamical variables in relativistic quantum mechanics. Journal of Mathematical Physics, 7(11):1959–1981, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i11/p1959_s1.

Handelsman:1966:ATC

[510] R. Handelsman and R. M. Lewis. The asymptotic theory of Cerenkov radiation. Journal of Mathematical Physics, 7(11):1982–1999, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p1982_s1.

Jones:1966:CTP

[511] Gerald L. Jones. Complex temperatures and phase transitions. Journal of Mathematical Physics, 7(11):2000–2005, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v7/i11/p2000_s1.

Jordan:1966:CCS

[512] Thomas F. Jordan. Characterizing coherent states of the radiation field. Journal of Mathematical Physics, 7(11):2006–2008, November 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i11/p2006_
s1.

Kacser:1966:DTG

[513] C. Kacser. The discontinuities of the triangle graph as a function of an internal mass. II. Journal of Mathematical Physics, 7(11):2008–2025, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i11/p2008_s1.

Limic:1966:CDR

[514] N. Limić, J. Niederle, and R. Raczka. Continuous degenerate representations of noncompact rotation groups. II. Journal of Mathematical Physics, 7(11):2026–2035, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i11/p2026_s1.

McCormick:1966:BOR

[515] N. J. McCormick and I. Kuscer. Bi-orthogonality relations for solving half-space transport problems. Journal of Mathematical Physics, 7(11):2036–2045, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i11/p2036_s1.

Mattis:1966:TPI

[516] Daniel C. Mattis and Elliott H. Lieb. Theory of paramagnetic impurities in semiconductors. Journal of Mathematical Physics, 7(11):2045– 2052, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i11/p2045_s1.

Mitzner:1966:ASI

[517] K. M. Mitzner. Acoustic scattering from an interface between media of greatly different density. Journal of Mathematical Physics, 7(11):2053– 2060, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i11/p2053_s1.

Robinson:1966:IFQ

[518] Peter D. Robinson. Integral forms for quantum-mechanical momentum operators. Journal of Mathematical Physics, 7(11):2060–2065, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p2060_s1.

Rosen:1966:EPS

[519] Gerald Rosen. Existence of particlelike solutions to nonlinear field theories. Journal of Mathematical Physics, 7(11):2066–2070, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p2066_s1.

Rosen:1966:NLP

[520] Gerald Rosen. Nonexistence of localized periodic solutions to nonlinear field theories. Journal of Mathematical Physics, 7(11):2071, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p2071_s1.

Rosen:1966:SOE

[521] Joe Rosen and Paul Roman. Some observations on enveloping algebras of noncompact groups. Journal of Mathematical Physics, 7(11):2072– 2078, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i11/p2072_s1.

Sabatier:1966:APC

[522] Pierre C. Sabatier. Analytic properties of a class of potentials and the corresponding Jost functions. Journal of Mathematical Physics, 7 (11):2079–2091, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i11/p2079_s1.

Siewert:1966:RTI

[523] C. E. Siewert and P. F. Zweifel. Radiative transfer. II. Journal of Mathematical Physics, 7(11):2092–2102, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i11/p2092_s1.

Spector:1966:TPS

[524] Richard M. Spector. A theorem on peratization of singular potentials and other miscellanea. Journal of Mathematical Physics, 7(11):2103– 2106, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i11/p2103_s1.

Prugovecki:1966:NHS

[525] Eduard Prugovecki. A non-Hilbert-space formulation of quantum field theory. Journal of Mathematical Physics, 7(11):2107–2120, November 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i11/p2107_s1.

Case:1966:SIE

[526] K. M. Case. Singular integral equations. Journal of Mathematical Physics, 7(12):2121–2124, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i12/p2121_s1.

Case:1966:SSC

[527] K. M. Case. Singular solutions of certain integral equations. Journal of Mathematical Physics, 7(12):2125–2135, December 1966. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i12/p2125_s1.

DesCloizeaux:1966:SFG

[528] J. Des Cloizeaux. A soluble Fermi-gas model. validity of transformations of the Bogoliubov type. Journal of Mathematical Physics, 7(12):2136– 2144, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2136_s1.

Compton:1966:TDG

[529] R. T. Compton. The time-dependent Green’s function for electromagnetic waves in moving simple media. Journal of Mathematical Physics, 7(12):2145–2152, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i12/p2145_s1.

Frieman:1966:PCB

[530] Edward A. Frieman and Robert Goldman. Propagation of correlations in a Boltzmann gas. Journal of Mathematical Physics, 7(12):2153–2170, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2153_s1.

Gale:1966:PLS

[531] William A. Gale. Peratization of the logarithmically singular potential. Journal of Mathematical Physics, 7(12):2171–2173, December 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i12/p2171_
s1.

Gotsman:1966:AAR

[532] E. Gotsman and Y. Ne’eman. Algebraic aspects of Regge recurrences. Journal of Mathematical Physics, 7(12):2174–2177, December 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i12/p2174_
s1.

Grover:1966:AED

[533] Richard Grover. Asymptotic expansions of the Dirac density matrix. Journal of Mathematical Physics, 7(12):2178–2186, December 1966. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i12/p2178_
s1.

Johnson:1966:SNQ

[534] B. Robert Johnson and Don Secrest. The solution of the nonrelativistic quantum scattering problem without exchange. Journal of Mathematical Physics, 7(12):2187–2195, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v7/i12/p2187_s1.

Keam:1966:PCG

[535] R. F. Keam. Properties of a causal Green’s function for the Bethe– Salpeter equation. Journal of Mathematical Physics, 7(12):2196–2210, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2196_s1.

Kuszell:1966:NDK

[536] A. Kuszell. A new derivation of the kinetic equation for an inhomogeneous plasma. Journal of Mathematical Physics, 7(12):2211–2216, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2211_s1.

Levy-Leblond:1966:GDM

[537] Jean-Marc Lévy-Leblond. Global and democratic methods for classifying N-particle states. Journal of Mathematical Physics, 7(12):2217–2229, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2217_s1.

Polkinghorne:1966:HAM

[538] J. C. Polkinghorne. Hermitian analyticity and S-matrix singularity structure. Journal of Mathematical Physics, 7(12):2230–2234, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i12/
p2230_s1.

Voisin:1966:RHG

[539] Jacques Voisin. Representations of the homogeneous Galilei group in a spherical basis and the asymptotic behavior of relativistic special functions. Journal of Mathematical Physics, 7(12):2235–2237, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i12/p2235_s1.

Baker:1966:DSI

[540] George A. Baker. Drum shapes and isospectral graphs. Journal of Mathematical Physics, 7(12):2238–2242, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i12/p2238_s1.

Chester:1966:FT

[541] G. V. Chester and J. Sykes. Factorization theorems. Journal of Mathematical Physics, 7(12):2243–2250, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v7/i12/p2243_s1.

Chester:1966:RBL

[542] G. V. Chester and J. Sykes. Relationship between linear response theory and the Boltzmann equation for a weakly interacting system of electrons and phonons. Journal of Mathematical Physics, 7(12):2251–2260, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2251_s1.

Horwitz:1966:SMF

[543] G. Horwitz. Statistical mechanics of finite systems: Asymptotic expansions. I. Two petit canonical ensembles — (N, V, T) and (N, p, T) systems. Journal of Mathematical Physics, 7(12):2261–2270, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v7/i12/p2261_s1.

Landsberg:1966:CSR

[544] P. T. Landsberg and D. J. Morgan. A class of sum rules with application to nondegenerate perturbation theory. Journal of Mathematical Physics, 7(12):2271–2277, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i12/p2271_s1.

Majumdar:1966:MPS

[545] C. K. Majumdar and J. Zittartz. Minimal potentials for schrödinger equation with fixed eigenvalue. Journal of Mathematical Physics, 7 (12):2278–2279, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i12/p2278_s1.

Alcaras:1966:EAG

[546] J. A. Castilho Alcarás, V. C. Aguilera Navarro, and P. Leal Ferreira. Erratum: “Applications of Gel’fand–Moshinsky Bases in Unitary Symmetry and its Breaking”. Journal of Mathematical Physics, 7(12):2280, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i12/p2280_s1. See [292].

Granzow:1966:EMT

[547] Kenneth D. Granzow. Erratum: “Multipole Theory in the Time Domain”. Journal of Mathematical Physics, 7(12):2280, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v7/i12/

p2280_s2. See [347].

Broido:1967:CBR

[548] Michael M. Broido. Certain bootstraps and related equations. Journal of Mathematical Physics, 8(1):1–6, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p1_s1.

Bronzan:1967:TMS

[549] J. B. Bronzan. Two-meson solution of the charged scalar static model with bound states. Journal of Mathematical Physics, 8(1):6–12, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i1/p6_s1.

Emch:1967:CIP

[550] Gérard G. Emch. Common interpretation of phase transitions in various models. Journal of Mathematical Physics, 8(1):13–18, January 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i1/p13_s1.

Emch:1967:VWW

[551] Gérard G. Emch. van der Waals wiggles, Maxwell rule, and temperaturedependent excitations. Journal of Mathematical Physics, 8(1):19–25, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i1/p19_s1.

Greenman:1967:HEB

[552] J. V. Greenman. High-energy behavior of Feynman integrals with spin. II. Journal of Mathematical Physics, 8(1):26–36, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i1/p26_s1.

Helstrom:1967:QCA

[553] Carl W. Helstrom. Quasi-classical analysis of coupled oscillators. Journal of Mathematical Physics, 8(1):37–42, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p37_s1.

Lieb:1967:CES

[554] Elliott H. Lieb. Calculation of exchange second virial coefficient of a hard-sphere gas by path integrals. Journal of Mathematical Physics, 8(1): 43–52, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i1/p43_s1.

Kostin:1967:GVM

[555] Morton D. Kostin and Harvey Brooks. Generalization of the variational method of kahan, rideau, and roussopoulos. II. A variational principle for linear operators and its application to neutron-transport theory. Journal of Mathematical Physics, 8(1):53–56, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p53_s1.

Mirman:1967:RIG

[556] Ronald Mirman. Representations of inhomogeneous SU(n) groups with mixture of homogeneous and inhomogeneous operators diagonal. Journal of Mathematical Physics, 8(1):57–62, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p57_s1.

Resnikoff:1967:GCC

[557] M. Resnikoff. General coupling coefficients for the group SU(3). Journal of Mathematical Physics, 8(1):63–79, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p63_s1.

Resnikoff:1967:RCG

[558] M. Resnikoff. Recoupling coefficients for the group SU(3). Journal of Mathematical Physics, 8(1):79–83, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p79_s1.

Rosenkilde:1967:SETa

[559] Carl E. Rosenkilde. Surface-energy tensors. Journal of Mathematical Physics, 8(1):84–88, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i1/p84_s1.

Rosenkilde:1967:SETb

[560] Carl E. Rosenkilde. Surface-energy tensors for ellipsoids. Journal of Mathematical Physics, 8(1):88–97, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i1/p88_s1.

Rosenkilde:1967:SAF

[561] Carl E. Rosenkilde. Stability of axisymmetric figures of equilibrium of a rotating charged liquid drop. Journal of Mathematical Physics, 8(1): 98–118, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i1/p98_s1.

Tauber:1967:EUC

[562] G. E. Tauber. Expanding universe in conformally flat coordinates. Journal of Mathematical Physics, 8(1):118–123, January 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i1/p118_s1.

Suzuki:1967:ODI

[563] Masuo Suzuki, Bunjiro Tsujiyama, and Shigetoshi Katsura. Onedimensional Ising model with general spin. Journal of Mathematical Physics, 8(1):124–130, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i1/p124_s1.

Wenger:1967:CRV

[564] Daniel L. Wenger. Canonical root vectors of SU_{n}. Journal of Mathematical Physics, 8(1):131–135, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i1/p131_s1.

Wenger:1967:RFG

[565] Daniel L. Wenger. Representation functions of the group of motions of Clifford space. Journal of Mathematical Physics, 8(1):135–140, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i1/p135_s1.

Gilbert:1967:RIO

[566] R. P. Gilbert and H. C. Howard. Role of the integral-operator method in the theory of potential scattering. Journal of Mathematical Physics, 8(1):141–148, January 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i1/p141_s1.

Pomraning:1967:VPE

[567] G. C. Pomraning. Variational principle for eigenvalue equations. Journal of Mathematical Physics, 8(1):149–158, January 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i1/p149_s1.

Bergmann:1967:ISE

[568] Saul M. Bergmann. Induced and spontaneous emission. Journal of Mathematical Physics, 8(2):159–169, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i2/p159_s1.

Evans:1967:DSU

[569] N. T. Evans. Discrete series for the universal covering group of the 3 + 2 de Sitter group. Journal of Mathematical Physics, 8(2):170–184, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p170_s1.

Fairlie:1967:SCM

[570] D. B. Fairlie. Structure of the crossing matrix for arbitrary internal symmetry groups. II. Matrices in SU(n). Journal of Mathematical Physics, 8(2):185–188, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i2/p185_s1.

Foster:1967:NRT

[571] J. Foster and E. T. Newman. Note on the Robinson–Trautman solutions. Journal of Mathematical Physics, 8(2):189–194, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i2/p189_s1.

Frank:1967:PLR

[572] William M. Frank. Point-loop renormalization in regularized field theories. I. Journal of Mathematical Physics, 8(2):194–200, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p194_s1.

Hill:1967:IAD

[573] Robert Nyden Hill. Instantaneous action-at-a-distance in classical relativistic mechanics. Journal of Mathematical Physics, 8(2):201–220, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i2/p201_s1.

Hosokawa:1967:FAC

[574] Iwao Hosokawa. Functional approach to classical non-equilibrium statistical mechanics. Journal of Mathematical Physics, 8(2):221–229, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p221_s1.

Maldonado:1967:SRM

[575] C. D. Maldonado. Series representation method for obtaining the emission coefficient from the integrated intensity distribution for asymmetrical light sources. Journal of Mathematical Physics, 8(2):230–239, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p230_s1.

McLeod:1967:RED

[576] J. B. McLeod and R. S. B. Ong. Relaxation to equilibrium of a dilute electron plasma. Journal of Mathematical Physics, 8(2):240–247, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p240_s1.

Su:1967:TLF

[577] C. H. Su. Theory of the linear Fokker–Planck collision operator. Journal of Mathematical Physics, 8(2):248–254, February 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i2/p248_s1.

Anderson:1967:RBF

[578] James L. Anderson and Samuel Schiminovich. Relations between fieldplus source and Fokker-type action principles. Journal of Mathematical Physics, 8(2):255–264, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i2/p255_s1.

Boyer:1967:MAE

[579] Robert H. Boyer and Richard W. Lindquist. Maximal analytic extension of the Kerr metric. Journal of Mathematical Physics, 8(2):265– 281, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i2/p265_s1.

Dorfman:1967:DKT

[580] J. R. Dorfman and E. G. D. Cohen. Difficulties in the kinetic theory of dense gases. Journal of Mathematical Physics, 8(2):282–297, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i2/p282_s1.

Chappell:1967:MAK

[581] Willard R. Chappell. Microscopic approach to kinetic theory. Journal of Mathematical Physics, 8(2):298–305, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i2/p298_s1.

Leff:1967:TIP

[582] Harvey S. Leff and Michael H. Coopersmith. Translational invariance properties of a finite one-dimensional hard-core fluid. Journal of Mathematical Physics, 8(2):306–314, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i2/p306_s1.

Dobson:1967:TAP

[583] Peter N. Dobson. Temperateness of the absorptive part of the scattering amplitude in the proof of dispersion relations. Journal of Mathematical Physics, 8(2):315–319, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i2/p315_s1.

Eden:1967:CCM

[584] R. J. Eden. Consistency conditions on models for high-energy scattering. Journal of Mathematical Physics, 8(2):320–325, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i2/p320_s1.

Ozsvath:1967:HLU

[585] István Ozsváth. Homogeneous Lichnerowicz universes. Journal of Mathematical Physics, 8(2):326–344, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i2/p326_s1.

Penrose:1967:TA

[586] R. Penrose. Twistor algebra. Journal of Mathematical Physics, 8(2):345– 366, February 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i2/p345_s1.

Aly:1967:SPZ

[587] H. H. Aly and H. J. W. Müller. Singular potentials and the Z = 0 condition. Journal of Mathematical Physics, 8(3):367–372, March 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p367_s1.

Atkinson:1967:NUT

[588] D. Atkinson and M. B. Halpern. Non-usual topologies on space–time and high-energy scattering. Journal of Mathematical Physics, 8(3):373– 387, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p373_s1.

Banta:1967:MEM

[589] Edwin D. Banta. Modification of the Ehrenfest model in statistical mechanics. Journal of Mathematical Physics, 8(3):387–388, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p387_s1.

Girardeau:1967:TEH

[590] M. D. Girardeau. Thermodynamically equivalent Hamiltonian method in nonequilibrium statistical mechanics. Journal of Mathematical Physics, 8(3):389–399, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i3/p389_s1.

Baxter:1967:LWW

[591] Glen Baxter. Local weights which determine area, and the Ising model. Journal of Mathematical Physics, 8(3):399–405, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p399_s1.

Bernstein:1967:NCR

[592] Herbert J. Bernstein. New characterization of the ray representations of the Galilei group. Journal of Mathematical Physics, 8(3):406–408, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p406_s1.

Besieris:1967:TDGa

[593] Ioannis M. Besieris. Time-dependent Green’s function for electromagnetic radiation in a conducting moving medium: Nonrelativistic approximation. Journal of Mathematical Physics, 8(3):409–416, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p409_s1.

Cercignani:1967:LRS

[594] Carlo Cercignani. Linear representation of spinors by tensors. Journal of Mathematical Physics, 8(3):417–422, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p417_s1.

Dyson:1967:SM

[595] Freeman J. Dyson and A. Lenard. Stability of matter. I. Journal of Mathematical Physics, 8(3):423–434, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p423_s1.

Coopersmith:1967:TIP

[596] Michael H. Coopersmith and Harvey S. Leff. Translational invariance properties of a one-dimensional fluid with forces of finite extent. Journal of Mathematical Physics, 8(3):434–442, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p434_s1.

Finkelstein:1967:SGH

[597] R. J. Finkelstein. Symmetry group of the hydrogen atom. Journal of Mathematical Physics, 8(3):443–449, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p443_s1.

Kiang:1967:VTR

[598] Hsiao S. Kiang. Variational theorem for reduced density matrices. Journal of Mathematical Physics, 8(3):450–451, March 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i3/p450_s1.

Fischer:1967:RPW

[599] C. R. Fischer and R. Akerib. Relativistic partial wave analysis of electron scattering. Journal of Mathematical Physics, 8(3):451–453, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p451_s1.

Flores:1967:MIR

[600] Jorge Flores. Mixed irreducible representations for U_{3}. Journal of Mathematical Physics, 8(3):454–459, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i3/p454_s1.

Jones:1967:DFF

[601] Gerald L. Jones and Robert E. Kennedy. Density fluctuations of a fluid. Journal of Mathematical Physics, 8(3):460–466, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p460_s1.

Frank:1967:SCL

[602] William M. Frank. Strong coupling limit in potential theory. I. Journal of Mathematical Physics, 8(3):466–472, March 1967. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p466_s1.

Goldstein:1967:VPW

[603] Rubin Goldstein. Variational principles and weighted averages. Journal of Mathematical Physics, 8(3):473–475, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p473_s1.

Schmidt:1967:CID

[604] Paul W. Schmidt. Calculation of intersect distribution functions in small angle X-ray scattering. Journal of Mathematical Physics, 8(3): 475–477, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p475_s1.

Griffiths:1967:CIFa

[605] Robert B. Griffiths. Correlations in Ising ferromagnets. I. Journal of Mathematical Physics, 8(3):478–483, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p478_s1.

Griffiths:1967:CIFb

[606] Robert B. Griffiths. Correlations in Ising ferromagnets. II. External magnetic fields. Journal of Mathematical Physics, 8(3):484–489, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p484_s1.

Navon:1967:ESL

[607] A. Navon and J. Patera. Embedding of a simple Lie group into a simple Lie group and branching rules. Journal of Mathematical Physics, 8(3): 489–493, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p489_s1.

Hammer:1967:TBS

[608] C. L. Hammer and T. A. Weber. Two-body scattering with wave packets. Journal of Mathematical Physics, 8(3):494–506, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p494_s1.

Jasnow:1967:NDI

[609] David Jasnow and Michael Wortis. New derivation of the Ising model cluster expansion. Journal of Mathematical Physics, 8(3):507–511, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p507_s1.

Masson:1967:PAP

[610] D. Masson. Padé approximant and partial-wave integral equation. Journal of Mathematical Physics, 8(3):512–514, March 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i3/p512_s1.

Strieder:1967:ULB

[611] William Strieder and Stephen Prager. Upper and lower bounds on Knudsen flow rates. Journal of Mathematical Physics, 8(3):514–517, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p514_s1.

Laporte:1967:WMB

[612] O. Laporte and R. G. Fowler. Weber’s mixed boundary-value problem in electrodynamics. Journal of Mathematical Physics, 8(3):518–522, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p518_s1.

Barston:1967:EPLa

[613] E. M. Barston. Eigenvalue problem for Lagrangian systems. Journal of Mathematical Physics, 8(3):523–532, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p523_s1.

Layzer:1967:LOC

[614] Arthur Layzer. Lowest-order correction to many-fermion vertex function. Journal of Mathematical Physics, 8(3):532–535, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p532_s1.

Macfarlane:1967:RIE

[615] A. J. Macfarlane, L. O’Raifeartaigh, and P. S. Rao. Relationship of the internal and external multiplicity structure of compact simple Lie groups.

Journal of Mathematical Physics, 8(3):536–546, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p536_s1.

McKean:1967:CEH

[616] H. P. McKean. Chapman–Enskog–Hilbert expansion for a class of solutions of the telegraph equation. Journal of Mathematical Physics, 8(3): 547–552, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p547_s1.

Nakayama:1967:DHP

[617] Toshio Nakayama and John M. Dawson. Derivations of hierarchies for Nparticle systems and Vlasov systems by means of the functional calculus. Journal of Mathematical Physics, 8(3):553–560, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p553_s1.

Nussenzveig:1967:PPC

[618] H. M. Nussenzveig. Phase problem in coherence theory. Journal of Mathematical Physics, 8(3):561–572, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p561_s1.

Rosen:1967:EMC

[619] Gerald Rosen. Equations of motion in classical nonlinear field theories. Journal of Mathematical Physics, 8(3):573–575, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p573_s1.

Rubin:1967:RWE

[620] Robert J. Rubin. Random walk with an excluded origin. Journal of Mathematical Physics, 8(3):576–581, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p576_s1.

Schultz:1967:SSE

[621] T. D. Schultz. Short simple evaluation of expressions of the Debye–Waller form valid for anharmonic modes not in thermal equilibrium. Journal of Mathematical Physics, 8(3):581, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i3/p581_s1.

Natarajan:1967:QRC

[622] S. Natarajan and K. Viswanath. Quaternionic representations of compact metric groups. Journal of Mathematical Physics, 8(3):582–589, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p582_s1.

Twersky:1967:MSE

[623] Victor Twersky. Multiple scattering of electromagnetic waves by arbitrary configurations. Journal of Mathematical Physics, 8(3):589– 610, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p589_s1.

Voisin:1967:RLA

[624] Jacques Voisin. Representations of the Lie algebra of the homogeneous Galilei group and their relation to the representations of the Lorentz algebra. Journal of Mathematical Physics, 8(3):611–614, March 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p611_s1.

Weinberg:1967:GTR

[625] Steven Weinberg. General theory of resistive beam instabilities. Journal of Mathematical Physics, 8(3):614–641, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i3/p614_s1.

Hostler:1967:RLV

[626] Levere Hostler. Runge–Lenz vector and the Coulomb Green’s function. Journal of Mathematical Physics, 8(3):642–646, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p642_s1.

Tai:1967:TDG

[627] C. T. Tai. Time-dependent Green’s function for a moving isotropic nondispersive medium. Journal of Mathematical Physics, 8(3):646– 647, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i3/p646_s1.

Halpern:1967:RBR

[628] Francis R. Halpern. Relation between the relativistic and nonrelativistic S-matrices. Journal of Mathematical Physics, 8(3):648–650, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p648_s1.

Hartle:1967:EDM

[629] James B. Hartle and John R. Taylor. Existence of N and D matrices. Journal of Mathematical Physics, 8(3):651–656, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p651_s1.

Tanabe:1967:ISL

[630] Kosai Tanabe and Kazuhisa Shima. Isoparity and simple Lie group. Journal of Mathematical Physics, 8(3):657–662, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/p657_s1.

DellAntonio:1967:TNP

[631] G. F. Dell’Antonio and S. Doplicher. Total number of particles and Fock representation. Journal of Mathematical Physics, 8(3):663–666, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i3/p663_s1.

Bartrum:1967:ETN

[632] P. C. Bartrum. Energy tensor of the null electromagnetic field. Journal of Mathematical Physics, 8(4):667–674, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p667_s1.

Boyce:1967:RPW

[633] J. F. Boyce. Relation of the O(2, 1) partial-wave expansion to the Regge representation. Journal of Mathematical Physics, 8(4):675–684, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p675_s1.

Barut:1967:TOM

[634] A. O. Barut and H. Kleinert. Tensor operators and mass formula in the minimal extension of U_{3} by charge conjugation. Journal of Mathematical Physics, 8(4):685–686, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i4/p685_s1.

Campbell:1967:EPS

[635] Graham H. Campbell, Joseph V. Lepore, and Robert J. Riddell. Evaluation of phase-space integrals. Journal of Mathematical Physics, 8(4): 687–691, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p687_s1.

Biedenharn:1967:CDW

[636] L. C. Biedenharn, A. Giovannini, and J. D. Louck. Canonical definition of Wigner coefficients in U_{n}. Journal of Mathematical Physics, 8(4): 691–700, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p691_s1.

Collinson:1967:NTA

[637] C. D. Collinson and D. C. French. Null tetrad approach to motions in empty space–time. Journal of Mathematical Physics, 8(4):701–708, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p701_s1.

Cook:1967:IRS

[638] T. Cook and B. Sakita. Induced representations of strong coupling groups. Journal of Mathematical Physics, 8(4):708–713, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p708_s1.

Derome:1967:SPJ

[639] Jean-Robert Derome. Symmetry properties of the 3 j symbols for SU(3). Journal of Mathematical Physics, 8(4):714–716, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p714_s1.

Cunningham:1967:BPS

[640] Andrew A. Cunningham. Bootstrap prediction of symmetry for a soluble static model. Journal of Mathematical Physics, 8(4):716–729, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p716_s1.

Doebner:1967:GES

[641] H. D. Doebner and G. C. Hegerfeldt. Group embedding of space–time and internal symmetry groups. Journal of Mathematical Physics, 8(4):

730–741, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p730_s1.

Essam:1967:DWF

[642] John W. Essam. Determination of weight factors in linked-cluster expansions for lattice systems. Journal of Mathematical Physics, 8(4): 741–749, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p741_s1.

Epstein:1967:CIM

[643] Henri Epstein. CTP invariance of the S-matrix in a theory of local observables. Journal of Mathematical Physics, 8(4):750–767, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p750_s1.

Galindo:1967:LAE

[644] A. Galindo. Lie algebra extensions of the Poincaré algebra. Journal of Mathematical Physics, 8(4):768–774, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p768_s1.

Gamba:1967:PED

[645] A. Gamba. Peculiarities of the eight-dimensional space. Journal of Mathematical Physics, 8(4):775–781, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i4/p775_s1.

Geroch:1967:TGR

[646] Robert P. Geroch. Topology in general relativity. Journal of Mathematical Physics, 8(4):782–786, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i4/p782_s1.

Gluck:1967:DRS

[647] M. Glück, Y. Gur, and J. Zak. Double representations of space groups. Journal of Mathematical Physics, 8(4):787–790, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p787_s1.

Lewis:1967:ETL

[648] Jordan D. Lewis. Expansion theorem for the linearized Fokker–Planck equation. Journal of Mathematical Physics, 8(4):791–798, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/
p791_s1.

Hestenes:1967:RSF

[649] David Hestenes. Real spinor fields. Journal of Mathematical Physics, 8(4):798–808, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i4/p798_s1.

Hestenes:1967:SI

[650] David Hestenes. Spin and isospin. Journal of Mathematical Physics, 8(4): 809–812, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p809_s1.

Hurley:1967:ODT

[651] James Hurley. One-dimensional three-body problem. Journal of Mathematical Physics, 8(4):813–815, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i4/p813_s1.

Innes:1967:QSM

[652] Frederick R. Innes. Quasi-spin methods and pairs of one-particle CFP in the seniority scheme. Journal of Mathematical Physics, 8(4):816–820, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p816_s1.

Kanwal:1967:TDM

[653] R. P. Kanwal. Theory of diffraction and matched asymptotic expansions. Journal of Mathematical Physics, 8(4):821–822, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p821_s1.

Kuscer:1967:CRE

[654] I. Kuscer and F. Shure. Closure relations for the eigenfunctions of the one-speed transport equation. Journal of Mathematical Physics, 8(4):

823–826, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p823_s1.

Miasek:1967:AHM

[655] Maria Miasek and Helena Dodziuk. Application of Houston’s method to the sum of plane waves over the Brillouin zone. II. Body-centered cubic lattice. Journal of Mathematical Physics, 8(4):827–829, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p827_s1.

Lanczos:1967:EEE

[656] C. Lanczos. Einstein equations and electromagnetism. Journal of Mathematical Physics, 8(4):829–836, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i4/p829_s1.

Lomont:1967:RRR

[657] J. S. Lomont and H. E. Moses. Reduction of reducible representations of the infinitesimal generators of the proper, orthochronous, inhomogeneous Lorentz group. Journal of Mathematical Physics, 8(4):837–850, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p837_s1.

Newton:1967:ECK

[658] Roger G. Newton. Equations with connected kernels for N-particle T operators. Journal of Mathematical Physics, 8(4):851–856, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p851_s1.

Nelson:1967:SHF

[659] T. J. Nelson. A set of harmonic functions for the group SU(3) as specialized matrix elements of a general finite transformation. Journal of Mathematical Physics, 8(4):857–863, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p857_s1.

Wu:1967:NCR

[660] F. Y. Wu, H. T. Tan, and E. Feenberg. Necessary conditions on radial distribution functions. Journal of Mathematical Physics, 8(4):864– 869, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p864_s1.

Schrader:1967:TVP

[661] David M. Schrader. Two variational principles for magnetostatic calculations. Journal of Mathematical Physics, 8(4):870–872, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p870_s1.

Nuttall:1967:ACE

[662] J. Nuttall. Analytic continuation of the off-energy-shell scattering amplitude. Journal of Mathematical Physics, 8(4):873–877, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p873_s1.

Pipkin:1967:PTM

[663] A. C. Pipkin and R. S. Rivlin. Phenomenological theory of magnetic hysteresis. Journal of Mathematical Physics, 8(4):878–883, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p878_s1.

Kadish:1967:ESN

[664] Abraham Kadish. Exponential stability of non-uniform guiding center plasma. Journal of Mathematical Physics, 8(4):884–888, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p884_s1.

Sirovich:1967:PWP

[665] Lawrence Sirovich and James K. Thurber. Plane wave propagation in kinetic theory. Journal of Mathematical Physics, 8(4):888–895, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p888_s1.

Sen:1967:CIR

[666] R. N. Sen. Construction of the irreducible representations of SU_{4}. Journal of Mathematical Physics, 8(4):896–905, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p896_s1.

Sabatier:1967:GMI

[667] Pierre C. Sabatier. General method for the inverse scattering problem at fixed energy. Journal of Mathematical Physics, 8(4):905–918, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p905_s1.

Wahlquist:1967:HNT

[668] Hugo D. Wahlquist and Frank B. Estabrook. Herglotz–Noether theorem in conformal space–time. Journal of Mathematical Physics, 8(4): 919, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i4/p919_s1.

Wait:1967:TDC

[669] James R. Wait. Theory of diffraction by a curved inhomogeneous body. Journal of Mathematical Physics, 8(4):920–925, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p920_s1.

Terwiel:1967:EED

[670] R. H. Terwiel. Evaluation of expressions of the Debye–Waller form. Journal of Mathematical Physics, 8(4):926, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p926_s1.

Wertheim:1967:IET

[671] M. S. Wertheim. Integral equations in the theory of classical fluids. Journal of Mathematical Physics, 8(4):927–938, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p927_s1.

Trollope:1967:NFE

[672] J. R. Trollope. Null fields in Einstein–Maxwell field theory. II. Journal of Mathematical Physics, 8(4):938–941, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p938_s1.

West:1967:SCL

[673] Geoffrey B. West. Scattering from Coulomb-like potentials without divergences or cutoffs. I. Formalism: Generalized Lippmann–Schwinger equations. Journal of Mathematical Physics, 8(4):942–953, April 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p942_s1.

Zaccaria:1967:CGS

[674] F. Zaccaria. Clebsch–Gordan series in the exceptional groups. Journal of Mathematical Physics, 8(4):953–954, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i4/p953_s1.

Poshusta:1967:PSR

[675] R. D. Poshusta. Permutational symmetry of reduced density matrices. Journal of Mathematical Physics, 8(4):955–961, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i4/p955_s1.

Wilcox:1967:EOP

[676] R. M. Wilcox. Exponential operators and parameter differentiation in quantum physics. Journal of Mathematical Physics, 8(4):962–982, April 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i4/p962_s1.

Blackman:1967:NAU

[677] Jerome Blackman. New approach to unified field theory. Journal of Mathematical Physics, 8(5):983–987, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i5/p983_s1.

Andersen:1967:CGS

[678] C. M. Andersen. Clebsch–Gordan series for symmetrized tensor products. Journal of Mathematical Physics, 8(5):988–997, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/p988_s1.

Anderson:1967:CES

[679] J. T. Anderson. Complete excitation spectra. Journal of Mathematical Physics, 8(5):998–1003, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i5/p998_s1.

Beyer:1967:ACL

[680] W. A. Beyer and L. Heller. Analytic continuation of Laplace transforms by means of asymptotic series. Journal of Mathematical Physics, 8(5): 1004–1018, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1004_s1.

Arai:1967:CEI

[681] Tadashi Arai and John R. Gabriel. Cluster expansion of an inverse overlap matrix for solids. Journal of Mathematical Physics, 8(5):1018–1025, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i5/p1018_s1.

Balaban:1967:SFM

[682] M. M. Balaban, A. E. Green, and P. M. Naghdi. Simple force multipoles in the theory of deformable surfaces. Journal of Mathematical Physics, 8(5):1026–1036, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i5/p1026_s1.

Critchfield:1967:BWE

[683] Charles L. Critchfield. Bargmann–Wigner equations. Journal of Mathematical Physics, 8(5):1037–1039, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i5/p1037_s1.

Erdmann:1967:NTP

[684] R. C. Erdmann. Neutron transport from a point source in two adjacent, dissimilar, semi-infinite media. Journal of Mathematical Physics, 8(5): 1040–1044, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1040_s1.

Harris:1967:BCT

[685] A. Brooks Harris. Bounds for certain thermodynamic averages. Journal of Mathematical Physics, 8(5):1044–1045, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i5/p1044_s1.

Hestenes:1967:SFD

[686] David Hestenes. Spinor fields as distortions of space–time. Journal of Mathematical Physics, 8(5):1046–1050, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i5/p1046_s1.

Gale:1967:RPS

[687] William A. Gale. Regularization and peratization of singular potentials. Journal of Mathematical Physics, 8(5):1050–1052, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/p1050_s1.

Lam:1967:RPC

[688] John Lam. Radiation of a point charge moving uniformly over an infinite array of conducting half-planes. Journal of Mathematical Physics, 8(5): 1053–1060, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1053_s1.

Mermin:1967:AOC

[689] N. D. Mermin. Absence of ordering in certain classical systems. Journal of Mathematical Physics, 8(5):1061–1064, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i5/p1061_s1.

Misra:1967:TDF

[690] R. M. Misra and R. A. Singh. Three-dimensional formulation of gravitational null fields. II. Journal of Mathematical Physics, 8(5):1065–1068, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i5/p1065_s1.

Mukunda:1967:RLA

[691] N. Mukunda. Realizations of Lie algebras in classical mechanics. Journal of Mathematical Physics, 8(5):1069–1072, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i5/p1069_s1.

Rieder:1967:PHC

[692] Z. Rieder, J. L. Lebowitz, and E. Lieb. Properties of a harmonic crystal in a stationary nonequilibrium state. Journal of Mathematical Physics, 8(5): 1073–1078, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1073_s1.

Limic:1967:EEA

[693] N. Limić, J. Niederle, and R. Raczka. Eigenfunction expansions associated with the second-order invariant operator on hyperboloids and cones. III. Journal of Mathematical Physics, 8(5):1079–1093, May 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/p1079_
s1.

Sack:1967:RES

[694] R. A. Sack, C. C. J. Roothaan, and W. Kołos. Recursive evaluation of some atomic integrals. Journal of Mathematical Physics, 8(5):1093– 1094, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1093_s1.

Ullah:1967:EDR

[695] Nazakat Ullah. Exact distributions of the reduced-width amplitude. Journal of Mathematical Physics, 8(5):1095–1096, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/p1095_s1.

Imre:1967:WMQ

[696] Kaya Imre, Ercüment Özizmir, Marcos Rosenbaum, and P. F. Zweifel. Wigner method in quantum statistical mechanics. Journal of Mathematical Physics, 8(5):1097–1108, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i5/p1097_s1.

Biswas:1967:SBS

[697] S. N. Biswas. Symmetries of the Bethe–Salpeter equation for relativistic bound-state problem. Journal of Mathematical Physics, 8(5):1109– 1111, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1109_s1.

Brandt:1967:BPH

[698] Richard A. Brandt. Bogoliubov–Parasiuk–Hepp renormalization theorem and spacelike regularization. Journal of Mathematical Physics, 8(5): 1112–1120, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i5/p1112_s1.

Frank:1967:CNP

[699] William M. Frank. Class of nonanalytic perturbations in quantum mechanics. Journal of Mathematical Physics, 8(5):1121–1127, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i5/p1121_s1.

Fleischman:1967:DSC

[700] Owen Fleischman and John G. Nagel. Double structure of a combined internal and space–time symmetry group and mass splitting. Journal of Mathematical Physics, 8(5):1128–1134, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i5/p1128_s1.

Moses:1967:RRW

[701] H. E. Moses. Reduction of relativistic wavefunctions to the irreducible representations of the inhomogeneous Lorentz group. I. Nonzero mass components. Journal of Mathematical Physics, 8(5):1134–1154, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i5/p1134_s1.

Pattnaik:1967:LLC

[702] Mamata Pattnaik. Leading Landau curves for a larger class of Feynman diagrams. Journal of Mathematical Physics, 8(5):1155–1162, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i5/p1155_s1.

Lezuo:1967:WCS

[703] Klaus J. Lezuo. Weyl coefficients in SU(3). Journal of Mathematical Physics, 8(5):1163–1170, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i5/p1163_s1.

Ellis:1967:DPF

[704] G. F. R. Ellis. Dynamics of pressure-free matter in general relativity. Journal of Mathematical Physics, 8(5):1171–1194, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i5/p1171_s1.

Hegerfeldt:1967:BTS

[705] G. C. Hegerfeldt. Branching theorem for the symplectic groups. Journal of Mathematical Physics, 8(6):1195–1196, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i6/p1195_s1.

Kremer:1967:FEG

[706] Hugo F. Kremer. Fresnel’s equation in general relativity. Journal of Mathematical Physics, 8(6):1197–1199, June 1967. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i6/p1197_s1.

Kobe:1967:DPC

[707] Donald H. Kobe. Derivation of the principle of compensation of dangerous diagrams. Journal of Mathematical Physics, 8(6):1200–1210, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i6/p1200_s1.

Levy-Nahas:1967:DCL

[708] Monique Levy-Nahas. Deformation and contraction of Lie algebras. Journal of Mathematical Physics, 8(6):1211–1222, June 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i6/p1211_s1.

Hong:1967:ATE

[709] Soonsung Hong. Asymptotic theory of electromagnetic and acoustic diffraction by smooth convex surfaces of variable curvature. Journal of Mathematical Physics, 8(6):1223–1232, June 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i6/p1223_s1.

Pang:1967:LRO

[710] Sing Chin Pang and K. T. Hecht. Lowering and raising operators for the orthogonal group in the chain O(n) ⊃ O(n − 1) ⊃ …, and their graphs. Journal of Mathematical Physics, 8(6):1233–1251, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i6/p1233_s1.

Sciarrino:1967:DUI

[711] A. Sciarrino and M. Toller. Decomposition of the unitary irreducible representations of the group SL(2C) restricted to the subgroup SU(1, 1). Journal of Mathematical Physics, 8(6):1252–1265, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i6/p1252_s1.

Zimering:1967:MME

[712] Shimshon Zimering. Mathematical methods for evaluating second-order three-body interactions between atoms or ions with Gaussian wavefunctions. Journal of Mathematical Physics, 8(6):1266–1270, June 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i6/p1266_
s1.

Woolcock:1967:ABS

[713] W. S. Woolcock. Asymptotic behavior of Stieltjes transforms. I. Journal of Mathematical Physics, 8(6):1270–1275, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i6/p1270_s1.

Coriell:1967:PDR

[714] S. R. Coriell and J. L. Jackson. Probability distribution of the radius of gyration of a flexible polymer. Journal of Mathematical Physics, 8(6): 1276–1284, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i6/p1276_s1.

Ezawa:1967:CBE

[715] Hiroshi Ezawa and Marshall Luban. Criterion for Bose–Einstein condensation and representation of canonical commutation relations. Journal of Mathematical Physics, 8(6):1285–1311, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i6/p1285_s1.

Fonda:1967:RCC

[716] Luciano Fonda and Gian Carlo Ghirardi. Renormalized coupling constants and the inversion problem. Journal of Mathematical Physics, 8(6): 1312–1314, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i6/p1312_s1.

Hakim:1967:RRSa

[717] Rémi Hakim. Remarks on relativistic statistical mechanics. I. Journal of Mathematical Physics, 8(6):1315–1344, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i6/p1315_s1.

Herbut:1967:BAA

[718] Fedor Herbut and Milan Vujicić. Basic algebra of antilinear operators and some applications. I. Journal of Mathematical Physics, 8(6):1345– 1354, June 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i6/p1345_s1.

Torrence:1967:AGR

[719] R. J. Torrence and A. I. Janis. Approach to gravitational radiation scattering. Journal of Mathematical Physics, 8(7):1355–1366, July 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1355_
s1.

Chakrabarti:1967:CFF

[720] A. Chakrabarti. Canonical form factors and current commutation relations. Journal of Mathematical Physics, 8(7):1367–1378, July 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1367_
s1.

Hakim:1967:RRSb

[721] Rémi Hakim. Remarks on relativistic statistical mechanics. II. Hierarchies for the reduced densities. Journal of Mathematical Physics, 8(7): 1379–1400, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i7/p1379_s1.

Hejtmanek:1967:TBR

[722] Hans Hejtmanek. Time behavior of a reactor and ergodic theory of semigroups. Journal of Mathematical Physics, 8(7):1401–1405, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1401_s1.

Nordtvedt:1967:UBE

[723] K. Nordtvedt. Upper bounds to eigenvalues of the one-dimensional Sturm–Liouville equation. Journal of Mathematical Physics, 8(7):1406– 1409, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i7/p1406_s1.

Goldman:1967:LDB

[724] Robert Goldman and Edward A. Frieman. Logarithmic density behavior of a nonequilibrium Boltzmann gas. Journal of Mathematical Physics, 8(7):1410–1426, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i7/p1410_s1.

Gibberd:1967:NAI

[725] R. W. Gibberd and C. A. Hurst. New approach to the Ising model. II. Journal of Mathematical Physics, 8(7):1427–1435, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1427_s1.

Gruber:1967:RBO

[726] B. Gruber. Relation between “outer” and “inner” multiplicity for SU(2) and SU(3). Journal of Mathematical Physics, 8(7):1436–1439, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1436_s1.

Rafanelli:1967:URL

[727] Kenneth Rafanelli. Unitary representations of the Lorentz group and particle dynamics. Journal of Mathematical Physics, 8(7):1440–1442, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1440_s1.

Wong:1967:EBS

[728] M. K. F. Wong. Electromagnetic behavior of superconductors with impurity scattering. Journal of Mathematical Physics, 8(7):1443–1448, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1443_s1.

Lewis:1967:UPS

[729] Robert M. Lewis. A unifying principle in statistical mechanics. Journal of Mathematical Physics, 8(7):1448–1459, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i7/p1448_s1.

Lanford:1967:IRI

[730] O. Lanford and D. Ruelle. Integral representations of invariant states on B^{∗} algebras. Journal of Mathematical Physics, 8(7):1460–1463, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1460_s1.

Bartrum:1967:NEF

[731] Peter C. Bartrum. Null electromagnetic field in the form of spherical radiation. Journal of Mathematical Physics, 8(7):1464–1467, July 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1464_
s1.

Boulware:1967:STC

[732] David G. Boulware and S. Deser. Stress-tensor commutators and Schwinger terms. Journal of Mathematical Physics, 8(7):1468–1477, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1468_s1.

Cohen:1967:NKM

[733] Jeffrey M. Cohen. Note on the Kerr metric and rotating masses. Journal of Mathematical Physics, 8(7):1477–1478, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i7/p1477_s1.

Darewych:1967:PSC

[734] G. W. Darewych and H. Schiff. Particlelike solutions of a class of nonlinear field equations. Journal of Mathematical Physics, 8(7):1479–1482, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1479_s1.

Jaen:1967:QCF

[735] Jong K. Jaen. Quantum correlation function for the noninteracting particle system. Journal of Mathematical Physics, 8(7):1483–1491, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i7/p1483_s1.

Dautcourt:1967:CHG

[736] G. Dautcourt. Characteristic hypersurfaces in general relativity. I. Journal of Mathematical Physics, 8(7):1492–1501, July 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1492_s1.

Tompkins:1967:DTC

[737] Donald R. Tompkins. Decomposition of tensors of the classical groups. Journal of Mathematical Physics, 8(7):1502–1514, July 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i7/p1502_s1.

Mavroyannis:1967:DEW

[738] C. Mavroyannis. Dispersion of electromagnetic waves in molecular crystals. Journal of Mathematical Physics, 8(8):1515–1522, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1515_s1.

Mavroyannis:1967:SEW

[739] C. Mavroyannis. Scattering of electromagnetic waves in molecular crystals. Journal of Mathematical Physics, 8(8):1522–1537, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1522_s1.

Dyson:1967:GSE

[740] Freeman J. Dyson. Ground-state energy of a finite system of charged particles. Journal of Mathematical Physics, 8(8):1538–1545, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1538_s1.

Mazo:1967:LBM

[741] R. M. Mazo. Lower bound on the mean kinetic energy of a system of particles. Journal of Mathematical Physics, 8(8):1546–1547, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1546_s1.

Collinson:1967:EST

[742] C. D. Collinson. Empty space–times algebraically special on a given world line or hypersurface. Journal of Mathematical Physics, 8(8):1547– 1550, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i8/p1547_s1.

Bohm:1967:AAP

[743] Arno Böhm. Associative algebra in the problem of mass formulas. Journal of Mathematical Physics, 8(8):1551–1558, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i8/p1551_s1.

Marx:1967:KGE

[744] Egon Marx. Klein–Gordon equation for spinors. Journal of Mathematical Physics, 8(8):1559–1566, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i8/p1559_s1.

Newton:1967:CBC

[745] R. G. Newton. Connection between complex angular momenta and the inverse scattering problem at fixed energy. Journal of Mathematical Physics, 8(8):1566–1570, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i8/p1566_s1.

Maduemezia:1967:MEI

[746] Awele Maduemezia. Matrix elements for irreducible representations of the U(6, 6) algebra in harmonic function space. Journal of Mathematical Physics, 8(8):1571–1580, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i8/p1571_s1.

Rice:1967:SOT

[747] T. M. Rice. Superconductivity in one and two dimensions. II. Charged systems. Journal of Mathematical Physics, 8(8):1581–1591, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1581_s1.

Dixon:1967:DEB

[748] W. G. Dixon. Description of extended bodies by multipole moments in special relativity. Journal of Mathematical Physics, 8(8):1591–1605, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1591_s1.

Stapp:1967:FNP

[749] Henry P. Stapp. Finiteness of the number of positive-α Landau surfaces in bounded portions of the physical region. Journal of Mathematical Physics, 8(8):1606–1610, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i8/p1606_s1.

Berk:1967:CBS

[750] H. L. Berk, D. L. Book, and D. Pfirsch. Convergence of the Bremmer series for the spatially inhomogeneous Helmholtz equation. Journal of Mathematical Physics, 8(8):1611–1619, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i8/p1611_s1.

Hendry:1967:SNT

[751] W. L. Hendry. Solutions to the neutron transport equation for a critical slab by perturbation theory. Journal of Mathematical Physics, 8(8):1619– 1630, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i8/p1619_s1.

Brysk:1967:DTD

[752] Henry Brysk. Drukarev transformation of Dirac equation. Journal of Mathematical Physics, 8(8):1631–1632, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i8/p1631_s1.

Blood:1967:RCT

[753] F. A. Blood. Relativistic and causal theories with four-fermion interaction Hamiltonians. Journal of Mathematical Physics, 8(8):1633– 1641, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i8/p1633_s1.

Dialetis:1967:STC

[754] D. Dialetis. Some theorems concerning the phase problem of coherence theory. Journal of Mathematical Physics, 8(8):1641–1649, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1641_s1.

Herzel:1967:GFD

[755] John C. Herzel. Green’s functions and double-time distribution functions in classical statistical mechanics. Journal of Mathematical Physics, 8(8):1650–1652, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i8/p1650_s1.

Cercignani:1967:EUL

[756] Carlo Cercignani. Existence and uniqueness in the large for boundary value problems in kinetic theory. Journal of Mathematical Physics, 8(8):1653–1656, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i8/p1653_s1.

Ruelle:1967:SCS

[757] D. Ruelle. States of classical statistical mechanics. Journal of Mathematical Physics, 8(8):1657–1668, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i8/p1657_s1.

Common:1967:ATP

[758] A. K. Common. Application of the theory of padé approximants to the solution of the N/D equations. Journal of Mathematical Physics, 8(8):1669–1675, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i8/p1669_s1.

Meyer:1967:NST

[759] Richard E. Meyer. Near-steady transition waves of a collisionless plasma. Journal of Mathematical Physics, 8(8):1676–1684, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i8/p1676_s1.

Streater:1967:UHR

[760] R. F. Streater. Uniqueness of the Haag–Ruelle scattering states. Journal of Mathematical Physics, 8(8):1685–1693, August 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i8/p1685_s1.

Kursunoglu:1967:URM

[761] Behram Kursunoğlu. Unitary representations of U(2, 2) and massless fields. Journal of Mathematical Physics, 8(8):1694–1699, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i8/p1694_s1.

Cantoni:1967:RSR

[762] Vittorio Cantoni. Reduction of some representations of the generalized Bondi–Metzner group. Journal of Mathematical Physics, 8(8):1700– 1706, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i8/p1700_s1.

Nickle:1967:ENQ

[763] H. H. Nickle. Erratum: “Nonrelativistic Quantum Theory of an Electron in an Arbitrarily Intense Laser Field”. Journal of Mathematical Physics, 8(8):1706, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i8/p1706_s1. See [456].

Richardson:1967:LCA

[764] John M. Richardson and Leo C. Levitt. Linear closure approximation method for classical statistical mechanics. Journal of Mathematical Physics, 8(9):1707–1715, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i9/p1707_s1.

Micha:1967:EMR

[765] David A. Micha. Expansion of the T matrix for resonance collisions. Journal of Mathematical Physics, 8(9):1716–1719, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1716_s1.

Das:1967:CEF

[766] A. Das and C. V. Coffman. A class of eigenvalues of the fine-structure constant and internal energy obtained from a class of exact solutions of the combined Klein–Gordon–Maxwell–Einstein field equations. Journal of Mathematical Physics, 8(9):1720–1735, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1720_s1.

Bahar:1967:GWM

[767] E. Bahar. Generalized WKB method with applications to problems of propagation in nonhomogeneous media. Journal of Mathematical Physics, 8(9):1735–1746, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i9/p1735_s1.

Epstein:1967:RRC

[768] Saul T. Epstein, Jean H. Epstein, and Barbara Kennedy. Recursion relations for Coulomb matrix elements. Journal of Mathematical Physics, 8(9):1747, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i9/p1747_s1.

Gorge:1967:AFQ

[769] V. Gorgé and F. Rohrlich. Analytic functionals in quantum field theory. Journal of Mathematical Physics, 8(9):1748–1756, September 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1748_
s1.

Hill:1967:CFR

[770] Robert Nyden Hill. Canonical formulation of relativistic mechanics. Journal of Mathematical Physics, 8(9):1756–1773, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1756_s1.

Sack:1967:ESH

[771] Robert A. Sack. Expansions in spherical harmonics. IV. Integral form of the radial dependence. Journal of Mathematical Physics, 8(9):1774–1784, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1774_s1.

Rulf:1967:RBC

[772] Benjamin Rulf. Relation between creeping waves and lateral waves on a curved interface. Journal of Mathematical Physics, 8(9):1785–1793, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1785_s1.

Mann:1967:EFS

[773] James E. Mann. Electric fields in a semi-infinite medium whose conductivity varies laterally. Journal of Mathematical Physics, 8(9):1794–1797, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1794_s1.

Mehta:1967:DHO

[774] C. L. Mehta. Degree of higher-order optical coherence. Journal of Mathematical Physics, 8(9):1798–1807, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i9/p1798_s1.

Williams:1967:EMC

[775] David N. Williams. Energy-momentum conservation implies translation invariance: Some didactic remarks. Journal of Mathematical Physics, 8(9):1807–1816, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i9/p1807_s1.

Heintz:1967:ACA

[776] W. H. Heintz and R. L. Gibbs. Angular coefficients of atomic matrix elements involving interelectronic coordinates. Journal of Mathematical Physics, 8(9):1817–1828, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i9/p1817_s1.

Paliov:1967:HEP

[777] A. Paliov and S. Rosendorff. High-energy phase shifts produced by repulsive singular potentials. Journal of Mathematical Physics, 8(9):1829– 1837, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1829_s1.

Lerche:1967:IVP

[778] Ian Lerche. Initial-value problem for relativistic plasma oscillations. Journal of Mathematical Physics, 8(9):1838–1847, September 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1838_
s1.

Gudder:1967:CMO

[779] Stanley Gudder. Coordinate and momentum observables in axiomatic quantum mechanics. Journal of Mathematical Physics, 8(9):1848–1858, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1848_s1.

Zwanziger:1967:ACN

[780] Daniel Zwanziger. Algebraic calculation of nonrelativistic Coulomb phase shifts. Journal of Mathematical Physics, 8(9):1858–1860, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i9/p1858_s1.

Vezzetti:1967:RIA

[781] David J. Vezzetti and Joseph B. Keller. Refractive index, attenuation, dielectric constant, and permeability for waves in a polarizable medium. Journal of Mathematical Physics, 8(9):1861–1870, September 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1861_
s1.

Loos:1967:GFP

[782] Hendricus G. Loos. Gauge field of a point charge. Journal of Mathematical Physics, 8(9):1870–1877, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i9/p1870_s1.

Pasquier:1967:TPS

[783] Jeanne-Yvonne Pasquier. Three-particle states and Green’s function. Journal of Mathematical Physics, 8(9):1878–1885, September 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1878_
s1.

Barston:1967:EPLb

[784] E. M. Barston. Eigenvalue problem for Lagrangian systems. II. Journal of Mathematical Physics, 8(9):1886–1892, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1886_s1.

Pao:1967:BVP

[785] Young ping Pao. Boundary-value problems for the linearized and weakly nonlinear Boltzmann equation. Journal of Mathematical Physics, 8 (9):1893–1898, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i9/p1893_s1.

Wong:1967:ROG

[786] M. K. F. Wong. Representations of the orthogonal group. I. Lowering and raising operators of the orthogonal group and matrix elements of the generators. Journal of Mathematical Physics, 8(9):1899–1911, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1899_s1.

Gaskill:1967:GVL

[787] J. R. Gaskill and M. Arenstein. Geometrical view of Lagrange multipliers in mechanics. Journal of Mathematical Physics, 8(9):1912–1915, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i9/p1912_s1.

Kapuy:1967:TBO

[788] E. Kapuy and N. H. March. Two-body orbitals for one-dimensional fermion gas with application to repulsive δ-function interactions. Journal of Mathematical Physics, 8(9):1915–1920, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i9/p1915_s1.

Niederle:1967:DDM

[789] J. Niederle. Decompositions of discrete most degenerate representations of SO_{0}(p, q) when restricted to representations of SO_{0}(p, q − 1) or SO_{0}(p−1, q). Journal of Mathematical Physics, 8(9):1921–1930, September 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i9/p1921_s1.

Yao:1967:UIR

[790] Tsu Yao. Unitary irreducible representations of SU(2, 2). I. Journal of Mathematical Physics, 8(10):1931–1954, October 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i10/p1931_s1.

Roffman:1967:RTL

[791] Eric H. Roffman. Representation theory of local current algebras. Journal of Mathematical Physics, 8(10):1954–1956, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p1954_s1.

Sabatier:1967:IFA

[792] Pierre C. Sabatier. Interpolation formulas in the angular momentum plane. Journal of Mathematical Physics, 8(10):1957–1972, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/
p1957_s1.

Rosenbaum:1967:GSM

[793] David M. Rosenbaum. General space-momentum commutation relation. Journal of Mathematical Physics, 8(10):1973–1978, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p1973_
s1.

Gutzwiller:1967:PIA

[794] Martin C. Gutzwiller. Phase-integral approximation in momentum space and the bound states of an atom. Journal of Mathematical Physics, 8(10):1979–2000, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p1979_s1.

Liboff:1967:CRB

[795] Richard L. Liboff and Giovanni E. Perona. Compatibility requirements in BBGKY expansion. Journal of Mathematical Physics, 8(10):2001– 2012, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i10/p2001_s1.

Almstrom:1967:APC

[796] H. Almström and P. Olsson. Analytic properties of certain radial matrix elements. Journal of Mathematical Physics, 8(10):2013–2021, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i10/p2013_s1.

Dilley:1967:EPW

[797] James Dilley. Evaluation of a partial-wave unitarity integral. Journal of Mathematical Physics, 8(10):2022–2025, October 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2022_s1.

Manoharan:1967:UBH

[798] A. C. Manoharan. Upper bounds on holomorphy envelopes for Wightman functions. Journal of Mathematical Physics, 8(10):2026–2029, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i10/p2026_s1.

McNamara:1967:INP

[799] B. McNamara and K. J. Whiteman. Invariants of nearly periodic Hamiltonian systems. Journal of Mathematical Physics, 8(10):2029–2038, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i10/p2029_s1.

Wong:1967:RCN

[800] Jack Wong. Results on certain non-Hermitian Hamiltonians. Journal of Mathematical Physics, 8(10):2039–2042, October 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2039_s1.

Stanciu:1967:FES

[801] George N. Stanciu. Further exact solutions of the Dirac equation. Journal of Mathematical Physics, 8(10):2043–2047, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2043_s1.

Chand:1967:RLA

[802] P. Chand, C. L. Mehta, N. Mukunda, and E. C. G. Sudarshan. Realization of Lie algebras by analytic functions of generators of a given Lie algebra. Journal of Mathematical Physics, 8(10):2048–2059, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i10/p2048_s1.

Rund:1967:SFO

[803] H. Rund. Systems of first-order linear ordinary differential equations as canonical or Euler–Lagrange equations. Journal of Mathematical Physics, 8(10):2059–2062, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i10/p2059_s1.

Kummer:1967:RPR

[804] Hans Kummer. n-representability problem for reduced density matrices. Journal of Mathematical Physics, 8(10):2063–2081, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2063_
s1.

Runnels:1967:PTB

[805] L. K. Runnels. Phase transition of a Bethe lattice gas of hard molecules. Journal of Mathematical Physics, 8(10):2081–2087, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2081_
s1.

Carroll:1967:EOE

[806] M. M. Carroll and R. S. Rivlin. Electro-optical effects. I. Journal of Mathematical Physics, 8(10):2088–2091, October 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2088_s1.

Cawley:1967:FPS

[807] Robert G. Cawley. Formal properties of the solution to the radiationless relativistic Kepler problem. Journal of Mathematical Physics, 8 (10):2092–2096, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p2092_s1.

Pomraning:1967:CVP

[808] G. C. Pomraning. Complementary variational principles and their application to neutron transport problems. Journal of Mathematical Physics, 8(10):2096–2108, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p2096_s1.

Gudder:1967:SOA

[809] Stanley Gudder. Systems of observables in axiomatic quantum mechanics. Journal of Mathematical Physics, 8(10):2109–2113, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/
p2109_s1.

Loos:1967:IHG

[810] Hendricus G. Loos. Internal holonomy groups of Yang–Mills fields. Journal of Mathematical Physics, 8(10):2114–2124, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2114_s1.

Judd:1967:CNS

[811] B. R. Judd and H. T. Wadzinski. A class of null spectroscopic coefficients. Journal of Mathematical Physics, 8(10):2125–2130, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2125_
s1.

Taylor:1967:TCP

[812] John R. Taylor. Timelike cluster properties in nonrelativistic scattering. Journal of Mathematical Physics, 8(10):2131–2137, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2131_
s1.

Bellman:1967:IIE

[813] R. Bellman, H. Kagiwada, R. Kalaba, and S. Ueno. Invariant imbedding equations for the dissipation functions of an inhomogeneous finite slab with anisotropic scattering. Journal of Mathematical Physics, 8 (10):2137–2142, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p2137_s1.

DeDominicis:1967:PCF

[814] C. De Dominicis and F. Englert. Potential-correlation function duality in statistical mechanics. Journal of Mathematical Physics, 8(10):2143– 2146, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i10/p2143_s1.

Finkelstein:1967:GFP

[815] R. Finkelstein and D. Levy. Group formulation of potential scattering. Journal of Mathematical Physics, 8(10):2147–2153, October 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i10/p2147_
s1.

Post:1967:NRC

[816] H. R. Post. A necessary restriction on correlation functions arising from 3- (or more) particle systems in one dimension. Journal of Mathematical Physics, 8(10):2154, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i10/p2154_s1.

Goldberg:1967:SSH

[817] J. N. Goldberg, A. J. Macfarlane, E. T. Newman, F. Rohrlich, and E. C. G. Sudarshan. Spin-s spherical harmonics and – D. Journal of Mathematical Physics, 8(11):2155–2161, November 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2155_s1.

Goldberg:1967:ITN

[818] Joshua N. Goldberg. Invariant transformations and Newman–Penrose constants. Journal of Mathematical Physics, 8(11):2161–2166, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i11/p2161_s1.

Lambropoulos:1967:SDE

[819] P. Lambropoulos. Solution of the differential equation (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0. Journal of Mathematical Physics, 8(11):2167–2169, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2167_s1.

Scott:1967:NFC

[820] W. T. Scott. New form of characteristic functional for cascade processes. Journal of Mathematical Physics, 8(11):2170–2179, November 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2170_
s1.

Canosa:1967:EMN

[821] José Canosa. Expansion method for nonlinear boundary-value problems. Journal of Mathematical Physics, 8(11):2180–2183, November 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2180_
s1.

Kijewski:1967:IMM

[822] L. Kijewski and J. K. Percus. Interpolation method for the many-body problem. Journal of Mathematical Physics, 8(11):2184–2193, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i11/p2184_s1.

Kaplan:1967:MPE

[823] L. M. Kaplan and M. Resnikoff. Matrix products and the explicit 3, 6, 9, and 12-j coefficients of the regular representation of SU(n). Journal of Mathematical Physics, 8(11):2194–2205, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2194_s1.

Radhakrishnan:1967:IMS

[824] D. Radhakrishnan and T. S. Santhanam. Internal multiplicity structure and Clebsch–Gordan series for the exceptional group G (2). Journal of Mathematical Physics, 8(11):2206–2209, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2206_s1.

Mukunda:1967:URG

[825] N. Mukunda. Unitary representations of the group O(2, 1) in an O(1, 1) basis. Journal of Mathematical Physics, 8(11):2210–2220, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i11/p2210_s1.

Staruszkiewicz:1967:GTS

[826] A. Staruszkiewicz. Geometric theory of the spin-weighted functions. Journal of Mathematical Physics, 8(11):2221–2222, November 1967. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2221_
s1.

Sawaguri:1967:FRE

[827] T. Sawaguri and W. Tobocman. Finite-range effects in distorted-wave born-approximation calculations of nucleon transfer reactions. Journal of Mathematical Physics, 8(11):2223–2230, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2223_s1.

Bacry:1967:PGT

[828] H. Bacry and J. L. Richard. Partial group-theoretical treatment for the relativistic hydrogen atom. Journal of Mathematical Physics, 8(11):2230– 2233, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i11/p2230_s1.

Friedel:1967:DNM

[829] Helmut Friedel. Discontinuities in nonideal magnetogasdynamics. Journal of Mathematical Physics, 8(11):2234–2236, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2234_s1.

Liu:1967:EDT

[830] C. H. Liu. Effective dielectric tensor and propagation constant of plane waves in a random anisotropic medium. Journal of Mathematical Physics, 8(11):2236–2242, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2236_s1.

Ferreira:1967:PVD

[831] Erasmo M. Ferreira, Nilo Guillén, and Javier Sesma. Properties of velocity-dependent potentials. Journal of Mathematical Physics, 8(11): 2243–2249, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2243_s1.

Alisauskas:1967:SGM

[832] S. J. Alisauskas and A. P. Jucys. Substitution group and mirror reflection symmetry in special unitary groups. Journal of Mathematical Physics, 8(11):2250–2255, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2250_s1.

Pauri:1967:CRR

[833] M. Pauri and G. M. Prosperi. Canonical realizations of the rotation group. Journal of Mathematical Physics, 8(11):2256–2267, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i11/p2256_s1.

Cornille:1967:CBM

[834] Henri Cornille. Connection between the Marchenko formalism and N/D equations: Regular interactions. I. Journal of Mathematical Physics, 8 (11):2268–2281, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2268_s1.

Atkinson:1967:ESC

[835] David Atkinson. Existence of solutions of crossing-symmetric N/D equations. Journal of Mathematical Physics, 8(11):2281–2296, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i11/p2281_s1.

Penney:1967:GCS

[836] R. Penney. Geometrization of a complex scalar field. II. Analysis. Journal of Mathematical Physics, 8(11):2297–2301, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i11/p2297_s1.

Estabrook:1967:TFJ

[837] Frank B. Estabrook and Hugo D. Wahlquist. Tetrad formulation of junction conditions in general relativity. Journal of Mathematical Physics, 8(11):2302–2308, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i11/p2302_s1.

Masson:1967:APS

[838] D. Masson. Analyticity in the potential strength. Journal of Mathematical Physics, 8(11):2308–2314, November 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i11/p2308_s1.

Farnsworth:1967:SNG

[839] David L. Farnsworth. Some new general relativistic dust metrics possessing isometries. Journal of Mathematical Physics, 8(12):2315–2317, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2315_s1.

Mockel:1967:IIG

[840] André Mockel. Invariant imbedding as a generalization of the resolvent equation. Journal of Mathematical Physics, 8(12):2318–2321, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i12/p2318_s1.

Ranada:1967:CM

[841] Antonio F. Rañada. Causality and the S matrix. Journal of Mathematical Physics, 8(12):2321–2326, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i12/p2321_s1.

Cox:1967:DMC

[842] Joseph R. Cox. On the determination of the many-channel potential matrix and S matrix from a single function. Journal of Mathematical Physics, 8(12):2327–2331, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i12/p2327_s1.

Ferdinand:1967:SMD

[843] Arthur E. Ferdinand. Statistical mechanics of dimers on a quadratic lattice. Journal of Mathematical Physics, 8(12):2332–2339, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i12/p2332_s1.

Lieb:1967:SDP

[844] Elliott H. Lieb. Solution of the dimer problem by the transfer matrix method. Journal of Mathematical Physics, 8(12):2339–2341, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i12/p2339_s1.

Chen:1967:FDV

[845] Tuan Wu Chen. Functional derivatives and vector meson fields. Journal of Mathematical Physics, 8(12):2342–2346, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2342_s1.

Newton:1967:GDM

[846] Roger G. Newton. Generalization of the determinantal method to continuous channels. Journal of Mathematical Physics, 8(12):2347–2350, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2347_s1.

Odeh:1967:EBT

[847] F. Odeh. Existence and bifurcation theorems for the Ginzburg–Landau equations. Journal of Mathematical Physics, 8(12):2351–2356, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i12/p2351_s1.

Spector:1967:PWS

[848] Richard M. Spector. Properties of the wavefunction for singular potentials. Journal of Mathematical Physics, 8(12):2357–2365, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v8/i12/p2357_s1.

Saenger:1967:BSE

[849] Robert M. Saenger. Bethe–Salpeter equation in momentum space. Journal of Mathematical Physics, 8(12):2366–2375, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2366_s1.

Meckler:1967:DCD

[850] Alvin Meckler. Dynamical content of differential geometry. Journal of Mathematical Physics, 8(12):2376–2392, December 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2376_s1.

Klauder:1967:WCP

[851] John R. Klauder. Weak correspondence principle. Journal of Mathematical Physics, 8(12):2392–2399, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i12/p2392_s1.

Rosen:1967:IDP

[852] Gerald Rosen. Internal dynamics of particlelike solutions to nonlinear field theories. Journal of Mathematical Physics, 8(12):2400–2406, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2400_s1.

Harris:1967:PSD

[853] Stewart Harris. Proof that successive derivatives of Boltzmann’s H function for a discrete velocity gas alternate in sign. Journal of Mathematical Physics, 8(12):2407–2410, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i12/p2407_s1.

Chen:1967:ERN

[854] Robert L. W. Chen. Energy requirement for nonlinear density fluctuations in a Vlasov plasma. Journal of Mathematical Physics, 8(12):2410– 2414, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2410_s1.

Federbush:1967:RAM

[855] Paul Federbush. Results on the analyticity of many-body scattering amplitudes in perturbation theory. Journal of Mathematical Physics, 8 (12):2415–2419, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i12/p2415_s1.

Swift:1967:DRT

[856] Arthur R. Swift. Daughter Regge trajectory in a field theory model. Journal of Mathematical Physics, 8(12):2420–2433, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2420_s1.

Smith:1967:SRD

[857] P. R. Smith and B. G. Wybourne. Selection rules and the decomposition of the Kronecker square of irreducible representations. Journal of Mathematical Physics, 8(12):2434–2440, December 1967. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2434_s1.

Lichtman:1967:EOS

[858] D. Lichtman and R. B. McQuistan. Exact occupation statistics for onedimensional arrays of dumbbells. Journal of Mathematical Physics, 8 (12):2441–2444, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v8/i12/p2441_s1.

Besieris:1967:TDGb

[859] Ioannis M. Besieris and Ralph T. Compton. Time-dependent Green’s function for electromagnetic waves in moving conducting media. Journal of Mathematical Physics, 8(12):2445–2451, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v8/i12/p2445_s1.

Katz:1967:ISC

[860] Amnon Katz. Infinite systems of classical particles. Journal of Mathematical Physics, 8(12):2451–2459, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v8/i12/p2451_s1.

Theumann:1967:SPG

[861] Alba Theumann. Single-particle Green’s function for a one-dimensional many-fermion system. Journal of Mathematical Physics, 8(12):2460– 2467, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2460_s1.

Smith:1967:HSG

[862] O. J. Smith and C. E. Siewert. The half-space Green’s function for an atmosphere with a polarized radiation field. Journal of Mathematical Physics, 8(12):2467–2474, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v8/i12/p2467_s1.

Tompkins:1968:STT

[863] Donald R. Tompkins. Symmetrized tensor “trace” operations. Journal of Mathematical Physics, 9(1):1–8, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i1/p1_s1.

Katzin:1968:RFI

[864] Gerald H. Katzin and Jack Levine. Related first integral theorem: A method for obtaining conservation laws of dynamical systems with geodesic trajectories in Riemannian spaces admitting symmetries. Journal of Mathematical Physics, 9(1):8–15, January 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i1/p8_s1.

Moses:1968:RRW

[865] H. E. Moses. Reduction of relativistic wavefunctions to the irreducible representations of the inhomogeneous Lorentz group. Part II. Zero-mass components. Journal of Mathematical Physics, 9(1):16–31, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i1/p16_s1.

Common:1968:PAB

[866] A. K. Common. Padé approximants and bounds to series of Stieltjes. Journal of Mathematical Physics, 9(1):32–38, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i1/p32_s1.

Mirman:1968:ISC

[867] Ronald Mirman. Invariants and scalars of compact inhomogeneous unitary algebras. Journal of Mathematical Physics, 9(1):39–46, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i1/p39_s1.

Mirman:1968:NPI

[868] Ronald Mirman. Number of polynomial invariants of adjoint and fundamental compact inhomogeneous unitary algebras. Journal of Mathematical Physics, 9(1):47–49, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i1/p47_s1.

Mukunda:1968:URH

[869] N. Mukunda. Unitary representations of the homogeneous Lorentz group in an O(2, 1) basis. Journal of Mathematical Physics, 9(1):50–61, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i1/p50_s1.

Roffman:1968:PK

[870] Eric H. Roffman. n-particle kinematics. Journal of Mathematical Physics, 9(1):62–65, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i1/p62_s1.

Leaf:1968:WTC

[871] Boris Leaf. Weyl transformation and the classical limit of quantum mechanics. Journal of Mathematical Physics, 9(1):65–72, January 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i1/p65_s1.

Fitzgerald:1968:GFN

[872] R. Fitzgerald and S. Schiminovich. Gauge fields with noninvariant interactions. Journal of Mathematical Physics, 9(1):73–80, January 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i1/p73_s1.

Erdmann:1968:GFO

[873] R. C. Erdmann and C. E. Siewert. Green’s functions for the onespeed transport equation in spherical geometry. Journal of Mathematical Physics, 9(1):81–89, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i1/p81_s1.

Calogero:1968:VPD

[874] F. Calogero and A. Degasperis. Values of the potential and its derivatives at the origin in terms of the S-wave phase shift and bound-state parameters. Journal of Mathematical Physics, 9(1):90–116, January 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i1/p90_s1.

Hakim:1968:RKE

[875] Rémi Hakim and André Mangeney. Relativistic kinetic equations including radiation effects. I. Vlasov approximation. Journal of Mathematical Physics, 9(1):116–130, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i1/p116_s1.

Clark:1968:CEM

[876] John W. Clark and Paul Westhaus. Cluster expansions in many-fermion theory. I. “Factor-cluster” formalisms. Journal of Mathematical Physics, 9(1):131–148, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i1/p131_s1.

Westhaus:1968:CEM

[877] Paul Westhaus and John W. Clark. Cluster expansions in many-fermion theory. II. Rearrangements of primitive decomposition equations. Journal of Mathematical Physics, 9(1):149–154, January 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i1/p149_s1.

Neuringer:1968:ESI

[878] Joseph L. Neuringer and James J. Stekert. Electromagnetic scattering from an inhomogeneous, collisionless plasma cylinder. Journal of Mathematical Physics, 9(1):155–163, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i1/p155_s1.

Matzner:1968:SMS

[879] Richard A. Matzner. Scattering of massless scalar waves by a Schwarzschild “singularity”. Journal of Mathematical Physics, 9(1):163– 170, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i1/p163_s1.

Flicker:1968:PDF

[880] Michael Flicker. Pair distribution function of a one-dimensional hard rod gas. Journal of Mathematical Physics, 9(1):171–173, January 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i1/p171_s1.

Carmi:1968:FGB

[881] Gideon Carmi. Further generalized Bose condensation, anisotropic ODLRO and thin^{4} he films. Journal of Mathematical Physics, 9(1):174– 180, January 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i1/p174_s1.

TenKate:1968:DAS

[882] A. Ten Kate. Dirac algebra and the six-dimensional Lorentz group. Journal of Mathematical Physics, 9(2):181–185, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p181_s1.

Gamba:1968:RCG

[883] A. Gamba. Representations and classes in groups of finite order. Journal of Mathematical Physics, 9(2):186–192, February 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i2/p186_s1.

Fleming:1968:SPG

[884] Gordon N. Fleming. Structure of the Poincaré generators. Journal of Mathematical Physics, 9(2):193–204, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i2/p193_s1.

Nesbet:1968:NWM

[885] R. K. Nesbet. Note on the WKB method. Journal of Mathematical Physics, 9(2):205–206, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i2/p205_s1.

Aslaksen:1968:URA

[886] Erik W. Aslaksen and John R. Klauder. Unitary representations of the affine group. Journal of Mathematical Physics, 9(2):206–211, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i2/p206_s1.

Litvin:1968:CGC

[887] D. B. Litvin and J. Zak. Clebsch–Gordan coefficients for space groups. Journal of Mathematical Physics, 9(2):212–221, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p212_s1.

Kerner:1968:RNR

[888] Edward H. Kerner. Remarks on the nature of relativistic particle orbits. Journal of Mathematical Physics, 9(2):222–232, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p222_s1.

Derrick:1968:PMI

[889] G. H. Derrick and Wan Kay-Kong. Particle motion and interaction in nonlinear field theories. Journal of Mathematical Physics, 9(2):232– 240, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i2/p232_s1.

Thompson:1968:ISI

[890] Colin J. Thompson. Infinite-spin Ising model in one dimension. Journal of Mathematical Physics, 9(2):241–245, February 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i2/p241_s1.

Scarfone:1968:SLM

[891] Leonard M. Scarfone. 2V sector of the Lee model. Journal of Mathematical Physics, 9(2):246–254, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i2/p246_s1.

Muller:1968:HES

[892] H. J. W. Müller and K. Schilcher. High-energy scattering for Yukawa potentials. Journal of Mathematical Physics, 9(2):255–259, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i2/p255_s1.

Handelsman:1968:ATC

[893] R. A. Handelsman and R. M. Lewis. Asymptotic theory of Cerenkov radiation in inhomogeneous media. Journal of Mathematical Physics, 9(2):260–269, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i2/p260_s1.

Plebanski:1968:ETS

[894] Jerzy Plebanski and John Stachel. Einstein tensor and spherical symmetry. Journal of Mathematical Physics, 9(2):269–283, February 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p269_s1.

Luehr:1968:IVT

[895] C. P. Luehr and M. Rosenbaum. Intrinsic vector and tensor techniques in Minkowski space with applications to special relativity. Journal of Mathematical Physics, 9(2):284–298, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i2/p284_s1.

Misra:1968:SQD

[896] S. P. Misra and T. S. Shankara. Semiclassical and quantum descriptions. Journal of Mathematical Physics, 9(2):299–304, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p299_s1.

Thiess:1968:AFQ

[897] Frank B. Thiess. Analytic functionals in quantum field theory and the regularization of ultraviolet divergences. Journal of Mathematical Physics, 9(2):305–315, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i2/p305_s1.

Kibble:1968:CSP

[898] T. W. B. Kibble. Coherent soft-photon states and infrared divergences. I. Classical currents. Journal of Mathematical Physics, 9(2):315–324, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i2/p315_s1.

Kritz:1968:EPD

[899] Arnold H. Kritz, S. Radin, and G. Sandri. Evolution of the probability distributions and of the correlation functions in a Bogoliubov plasma. Journal of Mathematical Physics, 9(2):325–335, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i2/p325_s1.

Kantowski:1968:AIO

[900] R. Kantowski. Another interpretation of the optical scalars. Journal of Mathematical Physics, 9(2):336–338, February 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i2/p336_s1.

Jackiw:1968:MUP

[901] Roman Jackiw. Minimum uncertainty product, number-phase uncertainty product, and coherent states. Journal of Mathematical Physics, 9(3):339–346, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i3/p339_s1.

Luban:1968:OOP

[902] Marshall Luban and Michael Revzen. Onset of ODLRO and the phase transition in the ideal boson gas. Journal of Mathematical Physics, 9(3): 347–349, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p347_s1.

Richmond:1968:ECM

[903] Peter Richmond and Geoffrey L. Sewell. Electron correlations, magnetic ordering, and Mott insulation in solids. Journal of Mathematical Physics, 9(3):349–356, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i3/p349_s1.

Giles:1968:FQS

[904] Robin Giles. Foundations for quantum statistics. Journal of Mathematical Physics, 9(3):357–371, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i3/p357_s1.

Baraff:1968:TEW

[905] G. A. Baraff. Transmission of electromagnetic waves through a conducting slab. I. The two-sided Wiener–Hopf solution. Journal of Mathematical Physics, 9(3):372–384, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i3/p372_s1.

Devi:1968:SLI

[906] V. Syamala Devi and L. S. R. K. Prasad. State labeling of the irreducible representations of SU_{n}. Journal of Mathematical Physics, 9(3):

384–385, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p384_s1.

Chadam:1968:UDP

[907] John M. Chadam. Unitarity of dynamical propagators of perturbed Klein–Gordon equations. Journal of Mathematical Physics, 9(3):386– 396, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p386_s1.

Jang:1968:RAW

[908] S. Jang. Relationships among the Wigner 9-j symbols. Journal of Mathematical Physics, 9(3):397–402, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i3/p397_s1.

Collinson:1968:EPF

[909] C. D. Collinson. Embeddings of the plane-fronted waves and other space– times. Journal of Mathematical Physics, 9(3):403–410, March 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i3/p403_s1.

Mukunda:1968:SDB

[910] N. Mukunda and E. C. G. Sudarshan. Structure of the Dirac bracket in classical mechanics. Journal of Mathematical Physics, 9(3):411– 417, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p411_s1.

Mukunda:1968:URL

[911] N. Mukunda. Unitary representations of the Lorentz groups: Reduction of the supplementary series under a noncompact subgroup. Journal of Mathematical Physics, 9(3):417–431, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i3/p417_s1.

Krieger:1968:APP

[912] Joseph B. Krieger. Asymptotic properties of perturbation theory. Journal of Mathematical Physics, 9(3):432–435, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i3/p432_s1.

Schlitt:1968:NSS

[913] D. W. Schlitt. Numerical solution of a singular integral equation encountered in polymer physics. Journal of Mathematical Physics, 9(3): 436–439, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p436_s1.

Chen:1968:SSW

[914] Yung Ming Chen. Scattering of scalar waves by a convex transparent object with statistical surface irregularities. Journal of Mathematical Physics, 9(3):439–450, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i3/p439_s1.

Geroch:1968:LCS

[915] Robert Geroch. Local characterization of singularities in general relativity. Journal of Mathematical Physics, 9(3):450–465, March 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i3/p450_s1.

Kelly:1968:GGI

[916] D. G. Kelly and S. Sherman. General Griffiths’ inequalities on correlations in Ising ferromagnets. Journal of Mathematical Physics, 9(3): 466–484, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p466_s1.

Couch:1968:TGW

[917] W. E. Couch, R. J. Torrence, A. I. Janis, and E. T. Newman. Tail of a gravitational wave. Journal of Mathematical Physics, 9(3):484– 496, March 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i3/p484_s1.

Estabrook:1968:DAS

[918] F. B. Estabrook, H. D. Wahlquist, and C. G. Behr. Dyadic analysis of spatially homogeneous world models. Journal of Mathematical Physics, 9(4):497–504, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i4/p497_s1.

Sharma:1968:AFT

[919] R. R. Sharma. Alpha-function technique for two-center integrals. Journal of Mathematical Physics, 9(4):505–509, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i4/p505_s1.

Broido:1968:RCE

[920] Michael M. Broido. Renormalization and composite effects in the Lee model. Journal of Mathematical Physics, 9(4):510–514, April 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i4/p510_s1.

Benioff:1968:SPC

[921] Paul A. Benioff. Some properties of the contact between theory and experiment. Journal of Mathematical Physics, 9(4):514–524, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i4/p514_s1.

Maradudin:1968:SPP

[922] A. A. Maradudin and I. P. Ipatova. Structure of the phonon propagator for an anharmonic crystal. Journal of Mathematical Physics, 9(4):525– 532, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i4/p525_s1.

Mukunda:1968:ZMR

[923] N. Mukunda. Zero-mass representations of the Poincaré group in an O(3, 1) basis. Journal of Mathematical Physics, 9(4):532–536, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i4/p532_s1.

Wray:1968:CFO

[924] J. G. Wray. Current formalism. I. Ordering theorems for currents. Journal of Mathematical Physics, 9(4):537–552, April 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i4/p537_s1.

Wray:1968:CFI

[925] J. G. Wray. Current formalism. II. The S matrix in perturbation theory. Journal of Mathematical Physics, 9(4):552–570, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i4/p552_s1.

Cooper:1968:PTC

[926] Martin J. Cooper. Point transformation of classical hard-core potential. Journal of Mathematical Physics, 9(4):571–578, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i4/p571_s1.

Dutta:1968:MBA

[927] S. K. Dutta, W. A. Wilson, and A. E. S. Green. Modified Born approximation and elastic scattering by weak central potentials. Journal of Mathematical Physics, 9(4):578–585, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i4/p578_s1.

Eftimiu:1968:GFP

[928] C. Eftimiu and K. W. Thompson. Green’s function of a particle in a uniform external field. Journal of Mathematical Physics, 9(4):586–587, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i4/p586_s1.

Gutkowski:1968:APC

[929] D. Gutkowski and A. Scalia. Analytic properties of a class of nonlocal interactions. I. Journal of Mathematical Physics, 9(4):588–592, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i4/p588_s1.

Martin:1968:AGF

[930] A. W. Martin and P. L. Pritchett. Asymptotic gravitational field of the “electron”. Journal of Mathematical Physics, 9(4):593–597, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i4/p593_s1.

Hawking:1968:GRE

[931] S. W. Hawking. Gravitational radiation in an expanding universe. Journal of Mathematical Physics, 9(4):598–604, April 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i4/p598_s1.

Craig:1968:PER

[932] R. A. Craig. Perturbation expansion for real-time Green’s functions. Journal of Mathematical Physics, 9(4):605–611, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i4/p605_s1.

Sankaranarayanan:1968:ETR

[933] A. Sankaranarayanan. Extension theorem and representations. Journal of Mathematical Physics, 9(4):611–620, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i4/p611_s1.

Dollard:1968:SST

[934] John D. Dollard. Screening in the Schrödinger theory of scattering. Journal of Mathematical Physics, 9(4):620–624, April 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i4/p620_s1.

Collinson:1968:NGT

[935] C. D. Collinson. Note on the geometric theory of neutrinos. Journal of Mathematical Physics, 9(4):625–627, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i4/p625_s1.

Kiefer:1968:CSO

[936] H. M. Kiefer and D. M. Fradkin. Comments on separability operators, invariance ladder operators, and quantization of the Kepler problem in prolate-spheroidal coordinates. Journal of Mathematical Physics, 9(4): 627–632, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i4/p627_s1.

Cercignani:1968:EUC

[937] Carlo Cercignani. Existence, uniqueness, and convergence of the solutions of models in kinetic theory. Journal of Mathematical Physics, 9(4): 633–639, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i4/p633_s1.

Kramer:1968:IRS

[938] Peter Kramer. Irreducible representations of the semidirect-product group K_{n} = A_{n}: S_{n} and the harmonic-oscillator shell model. Journal of Mathematical Physics, 9(4):639–649, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i4/p639_s1.

Baxter:1968:DRL

[939] R. J. Baxter. Dimers on a rectangular lattice. Journal of Mathematical Physics, 9(4):650–654, April 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i4/p650_s1.

Gordon:1968:EBE

[940] Roy G. Gordon. Error bounds in equilibrium statistical mechanics. Journal of Mathematical Physics, 9(5):655–663, May 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i5/p655_s1.

Sharp:1968:PB

[941] R. T. Sharp and S. C. Pieper. O(5) Polynomial bases. Journal of Mathematical Physics, 9(5):663–667, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i5/p663_s1.

delaPena-Auerbach:1968:QMD

[942] L. de la Peña-Auerbach, E. Braun, and L. S. García Colín. Quantummechanical description of a Brownian particle. Journal of Mathematical Physics, 9(5):668–674, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i5/p668_s1.

Goldberg:1968:GTI

[943] Joshua N. Goldberg. Green’s theorem and invariant transformations. Journal of Mathematical Physics, 9(5):674–679, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p674_s1.

Zwanzig:1968:FSH

[944] Robert Zwanzig. Frequency spectra of harmonic lattices with weak longrange interactions. Journal of Mathematical Physics, 9(5):679–685, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i5/p679_s1.

Roth:1968:EIC

[945] Charles Roth. Electrostatic interactions of the configuration d^{n−1}sp. Journal of Mathematical Physics, 9(5):686–692, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p686_s1.

Mehta:1968:SII

[946] C. L. Mehta. Some inequalities involving traces of operators. Journal of Mathematical Physics, 9(5):693–697, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p693_s1.

Lenard:1968:SMI

[947] A. Lenard and Freeman J. Dyson. Stability of matter. II. Journal of Mathematical Physics, 9(5):698–711, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p698_s1.

Johnson:1968:OBP

[948] Porter Johnson. Orear behavior in potential scattering. Journal of Mathematical Physics, 9(5):712–718, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i5/p712_s1.

Wilner:1968:IES

[949] M. Wilner. An integral equation for the scattering operator. Journal of Mathematical Physics, 9(5):719–724, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p719_s1.

Alexanian:1968:FTF

[950] Moorad Alexanian. Field-theoretic formulation of quantum statistical mechanics. Journal of Mathematical Physics, 9(5):725–733, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i5/p725_s1.

Alexanian:1968:IRS

[951] Moorad Alexanian. Integral representation for systems of interacting particles. Journal of Mathematical Physics, 9(5):734–737, May 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p734_s1.

Komy:1968:SPL

[952] S. Komy and L. O’Raifeartaigh. Some properties of ladder operators. Journal of Mathematical Physics, 9(5):738–744, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p738_s1.

Diesendorf:1968:QCP

[953] Mark Diesendorf and Barry W. Ninham. Quantum corrections to the pair distribution function of a classical plasma. Journal of Mathematical Physics, 9(5):745–752, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i5/p745_s1.

Bowden:1968:STE

[954] Robert L. Bowden, F. Joseph McCrosson, and Edgar A. Rhodes. Solution of the transport equation with anisotropic scattering in slab geometry. Journal of Mathematical Physics, 9(5):753–759, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p753_s1.

Greene:1968:TDM

[955] John M. Greene. Two-dimensional measure-preserving mappings. Journal of Mathematical Physics, 9(5):760–768, May 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i5/p760_s1.

Leaf:1968:WTN

[956] Boris Leaf. Weyl transform in nonrelativistic quantum dynamics. Journal of Mathematical Physics, 9(5):769–781, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p769_s1.

Aly:1968:NFT

[957] H. H. Aly and J. G. Taylor. New field-theory approach to singular potentials. Journal of Mathematical Physics, 9(5):781–785, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p781_s1.

Peres:1968:RPG

[958] Asher Peres. Realization of Poincaré-group generators on a light cone. Journal of Mathematical Physics, 9(5):785–789, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/p785_s1.

Eriksen:1968:PHO

[959] Erik Eriksen. Properties of higher-order commutator products and the Baker–Hausdorff formula. Journal of Mathematical Physics, 9(5):790– 796, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i5/p790_s1.

Nutt:1968:TTP

[960] Gerald L. Nutt. Two- and three-particle Coulomb systems. Journal of Mathematical Physics, 9(5):796–803, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p796_s1.

Rosen:1968:NFE

[961] Gerald Rosen. Nonexistence of finite-energy stationary quantum states in nonlinear field theories. Journal of Mathematical Physics, 9(5):804–805, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i5/p804_s1.

Eyges:1968:IMQ

[962] Leonard Eyges and John R. Jasperse. Improved method for quantummechanical three-body problem III. Use of Sturmian functions. Journal of Mathematical Physics, 9(5):805–812, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i5/p805_s1.

Bichteler:1968:GES

[963] Klaus Bichteler. Global existence of spin structures for gravitational fields. Journal of Mathematical Physics, 9(6):813–815, June 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i6/p813_s1.

Berzi:1968:LGL

[964] Vittorio Berzi and Vittorio Gorini. Lie-group and Lie-algebra inhomogenizations. Journal of Mathematical Physics, 9(6):816–828, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i6/p816_s1.

Berzi:1968:ICR

[965] Vittorio Berzi and Vittorio Gorini. Inhomogenizations and complex representations. Journal of Mathematical Physics, 9(6):829–836, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i6/p829_s1.

Hartwig:1968:IMS

[966] Robert E. Hartwig and John Stephenson. Ising-model spin correlations. Journal of Mathematical Physics, 9(6):836–848, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i6/p836_s1.

Rosenzweig:1968:EQC

[967] C. Rosenzweig and J. B. Krieger. Exact quantization conditions. Journal of Mathematical Physics, 9(6):849–860, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i6/p849_s1.

Winicour:1968:STI

[968] J. Winicour. Some total invariants of asymptotically flat space–times. Journal of Mathematical Physics, 9(6):861–867, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i6/p861_s1.

Mason:1968:EFP

[969] Rodney J. Mason. Electric-field penetration into a plasma with a fractionally accommodating boundary. Journal of Mathematical Physics, 9(6):868–874, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i6/p868_s1.

Green:1968:GOZ

[970] Melville S. Green. Generalized Ornstein–Zernike approach to critical phenomena. Journal of Mathematical Physics, 9(6):875–890, June 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i6/p875_s1.

Corbett:1968:CBS

[971] John V. Corbett. Convergence of the Born series. Journal of Mathematical Physics, 9(6):891–898, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i6/p891_s1.

Wannier:1968:BBF

[972] G. H. Wannier and J. P. Van Dyke. Are Bloch bands at finite field adiabatically connected to those at zero field? Journal of Mathematical Physics, 9(6):899–904, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i6/p899_s1.

Cohen:1968:AMK

[973] Jeffrey M. Cohen. Angular momentum and the Kerr metric. Journal of Mathematical Physics, 9(6):905–906, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i6/p905_s1.

Bellman:1968:NDI

[974] R. Bellman, H. Kagiwada, R. Kalaba, and S. Ueno. New derivation of the integro-differential equations for Chandrasekhar’s X and Y functions. Journal of Mathematical Physics, 9(6):906–908, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i6/p906_s1.

Bellman:1968:DTL

[975] R. E. Bellman, H. H. Kagiwada, R. E. Kalaba, and S. Ueno. Diffuse transmission of light from a central source through an inhomogeneous spherical shell with isotropic scattering. Journal of Mathematical Physics, 9(6): 909–912, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i6/p909_s1.

Schild:1968:ETB

[976] A. Schild and J. A. Schlosser. Electromagnetic two-body problem for particles with spin. Journal of Mathematical Physics, 9(6):913–915, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i6/p913_s1.

delaPena-Auerbach:1968:PIQ

[977] L. de la Peña-Auerbach and Leopoldo S. García-Colín. Possible interpretation of quantum mechanics. Journal of Mathematical Physics, 9(6):

916–921, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i6/p916_s1.

delaPena-Auerbach:1968:SGS

[978] L. de la Peña-Auerbach and Leopoldo S. García-Colín. Simple generalization of Schrödinger’s equation. Journal of Mathematical Physics, 9(6): 922–927, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i6/p922_s1.

Carruthers:1968:LFTa

[979] Peter Carruthers. Local field theory and isospin invariance. I. Free-field theory of spinless bosons. Journal of Mathematical Physics, 9(6):928– 945, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i6/p928_s1.

Emch:1968:NSM

[980] Gerard G. Emch and Geoffrey L. Sewell. Nonequilibrium statistical mechanics of open systems. Journal of Mathematical Physics, 9(6):946–958, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i6/p946_s1.

Fronsdal:1968:MEO

[981] C. Fronsdal and Harun Ar-Rashid. Matrix elements of operators of class I representations of SL(n, C). Journal of Mathematical Physics, 9(6): 959–964, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i6/p959_s1.

Storer:1968:PIC

[982] R. G. Storer. Path-integral calculation of the quantum-statistical density matrix for attractive Coulomb forces. Journal of Mathematical Physics, 9(6):964–970, June 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i6/p964_s1.

Cocke:1968:CPS

[983] W. J. Cocke and Jeffrey M. Cohen. Cauchy problem in the scalar-tensor gravitational theory. Journal of Mathematical Physics, 9(7):971–973, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p971_s1.

Gabriel:1968:NMR

[984] John Gabriel. New methods for reduction of group representations. II. Journal of Mathematical Physics, 9(7):973–976, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p973_s1.

Scarfone:1968:TDM

[985] Leonard M. Scarfone. Tamm–Dancoff method in the 2V sector of the Lee model. Journal of Mathematical Physics, 9(7):977–983, July 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p977_s1.

Katz:1968:DNL

[986] Amnon Katz. Derivation of Newton’s law of gravitation from general relativity. Journal of Mathematical Physics, 9(7):983–985, July 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p983_s1.

Dombey:1968:APS

[987] Norman Dombey and R. H. Jones. Analyticity properties of the scattering amplitude for singular potentials. Journal of Mathematical Physics, 9(7):986–995, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i7/p986_s1.

Rosen:1968:PSN

[988] Gerald Rosen. Particlelike solutions to nonlinear complex scalar field theories with positive-definite energy densities. Journal of Mathematical Physics, 9(7):996–998, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i7/p996_s1.

Rosen:1968:CPS

[989] Gerald Rosen. Charged particlelike solutions to nonlinear complex scalar field theories. Journal of Mathematical Physics, 9(7):999–1002, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p999_s1.

Calogero:1968:VBP

[990] F. Calogero, O. D. Corbella, A. Degasperis, and M. B. De Stefano. Variational bounds for the potential in terms of the S-wave phase shift. Journal of Mathematical Physics, 9(7):1002–1006, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i7/p1002_s1.

Nagle:1968:WGM

[991] John F. Nagle. Weak-graph method for obtaining formal series expansions for lattice statistical problems. Journal of Mathematical Physics, 9(7):1007–1019, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i7/p1007_s1.

Nagle:1968:CTG

[992] J. F. Nagle and H. N. V. Temperley. Combinatorial theorem for graphs on a lattice. Journal of Mathematical Physics, 9(7):1020–1026, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p1020_s1.

Beltrametti:1968:RLG

[993] E. G. Beltrametti and A. A. Blasi. Rotation and Lorentz groups in a finite geometry. Journal of Mathematical Physics, 9(7):1027–1035, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p1027_s1.

Conger:1968:MCD

[994] Robert L. Conger. Model for converging detonations in solid explosives. Journal of Mathematical Physics, 9(7):1036–1040, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p1036_s1.

Smith:1968:PTC

[995] P. R. Smith and B. G. Wybourne. Plethysm and the theory of complex spectra. Journal of Mathematical Physics, 9(7):1040–1051, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p1040_s1.

Misra:1968:TNV

[996] M. Misra. Type-null vacuum solutions in general relativity. Journal of Mathematical Physics, 9(7):1052, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i7/p1052_s1.

Kaufmann:1968:RSS

[997] William J. Kaufmann III. Radiating spheres in the scalar-tensor theory of general relativity. Journal of Mathematical Physics, 9(7):1053– 1057, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1053_s1.

Devi:1968:BRC

[998] V. Syamala Devi and T. Venkatarayudu. Bases for the representations of u_{4} in the chain U_{4} ⊃ U_{2} + U_{2}. Journal of Mathematical Physics, 9(7): 1057–1058, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1057_s1.

Thompson:1968:SMI

[999] Colin J. Thompson. Spherical model as an instance of eigenvalue degeneracy. Journal of Mathematical Physics, 9(7):1059–1062, July 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p1059_
s1.

Matzner:1968:SBS

[1000] Richard A. Matzner. 3-sphere “backgrounds” for the space sections of the Taub cosmological solution. Journal of Mathematical Physics, 9(7): 1063–1067, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1063_s1.

Corio:1968:COA

[1001] P. L. Corio. Construction of orthonormal angular-momentum operators. Journal of Mathematical Physics, 9(7):1067–1071, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p1067_s1.

Stewart:1968:SEE

[1002] J. M. Stewart and G. F. R. Ellis. Solutions of Einstein’s equations for a fluid which exhibit local rotational symmetry. Journal of Mathematical Physics, 9(7):1072–1082, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i7/p1072_s1.

Armstrong:1968:REA

[1003] Lloyd Armstrong, Jr. Relativistic effects in atomic fine structure. II. Journal of Mathematical Physics, 9(7):1083–1086, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i7/p1083_s1.

Gordon:1968:EBS

[1004] R. G. Gordon. Error bounds in spectroscopy and nonequilibrium statistical mechanics. Journal of Mathematical Physics, 9(7):1087–1092, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i7/p1087_s1.

Pearle:1968:CIC

[1005] Philip Pearle. Construction of an invariance from a conservation law, and vice versa, in classical mechanics. Journal of Mathematical Physics, 9(7): 1092–1102, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1092_s1.

Whitten:1968:SRT

[1006] R. C. Whitten and F. T. Smith. Symmetric representation for three-body problems. II. Motion in space. Journal of Mathematical Physics, 9(7): 1103–1113, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1103_s1.

McDonald:1968:NNS

[1007] J. F. McDonald and L. D. Favro. Nearest-neighbor spacing distribution for an ensemble of random matrices with a nonrandom bias. Journal of Mathematical Physics, 9(7):1114–1119, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i7/p1114_s1.

Lanford:1968:MES

[1008] Oscar E. Lanford III and Derek W. Robinson. Mean entropy of states in quantum-statistical mechanics. Journal of Mathematical Physics, 9(7):

1120–1125, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i7/p1120_s1.

Lutzky:1968:PDE

[1009] M. Lutzky. Parameter differentiation of exponential operators and the Baker–Campbell–Hausdorff formula. Journal of Mathematical Physics, 9(7):1125–1128, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i7/p1125_s1.

Rosenblatt:1968:RBE

[1010] M. Rosenblatt. Remarks on the Burgers equation. Journal of Mathematical Physics, 9(7):1129–1136, July 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i7/p1129_s1.

Challifour:1968:TOP

[1011] John L. Challifour. Time-ordered products in two-dimensional field theories. Journal of Mathematical Physics, 9(8):1137–1145, August 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i8/p1137_
s1.

Pauri:1968:CRG

[1012] M. Pauri and G. M. Prosperi. Canonical realizations of the Galilei group. Journal of Mathematical Physics, 9(8):1146–1162, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i8/p1146_s1.

Miller:1968:SFCa

[1013] Willard Miller, Jr. Special functions and the complex Euclidean group in 3-space. I. Journal of Mathematical Physics, 9(8):1163–1175, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1163_s1.

Miller:1968:SFCb

[1014] Willard Miller, Jr. Special functions and the complex Euclidean group in 3-space. II. Journal of Mathematical Physics, 9(8):1175–1187, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1175_s1.

Pajas:1968:DRS

[1015] P. Pajas and R. Raczka. Degenerate representations of the symplectic groups. I. The compact group sp(n). Journal of Mathematical Physics, 9(8):1188–1201, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1188_s1.

Miura:1968:KVEa

[1016] Robert M. Miura. Korteweg–de Vries equation and generalizations. I. A remarkable explicit nonlinear transformation. Journal of Mathematical Physics, 9(8):1202–1204, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i8/p1202_s1.

Miura:1968:KVEb

[1017] Robert M. Miura, Clifford S. Gardner, and Martin D. Kruskal. Korteweg–de Vries equation and generalizations. II. Existence of conservation laws and constants of motion. Journal of Mathematical Physics, 9(8):1204–1209, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1204_s1.

Ferreira:1968:PMV

[1018] Erasmo M. Ferreira, Nilo Guillén, and Javier Sesma. Properties of the S matrix for velocity-dependent potentials. Journal of Mathematical Physics, 9(8):1210–1223, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i8/p1210_s1.

Kemmer:1968:IRF

[1019] N. Kemmer, D. L. Pursey, and S. A. Williams. Irreducible representations of the five-dimensional rotation group. I. Journal of Mathematical Physics, 9(8):1224–1230, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i8/p1224_s1.

Williams:1968:IRF

[1020] S. A. Williams and D. L. Pursey. Irreducible representations of the five-dimensional rotation group. II. Journal of Mathematical Physics, 9(8):1230–1240, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1230_s1.

Sabatier:1968:ASP

[1021] Pierre C. Sabatier. Approach to scattering problems through Interpolation formulas and application to spin-orbit potentials. Journal of Mathematical Physics, 9(8):1241–1258, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i8/p1241_s1.

Humi:1968:EFM

[1022] Mayer Humi. Extension of the factorization method. Journal of Mathematical Physics, 9(8):1258–1265, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i8/p1258_s1.

Kadish:1968:EEI

[1023] Abraham Kadish. Extended energy-integral technique for linear differential equations. Journal of Mathematical Physics, 9(8):1266–1269, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1266_s1.

Miller:1968:CSE

[1024] Marvin M. Miller. Convergence of the Sudarshan expansion for the diagonal coherent-state weight functional. Journal of Mathematical Physics, 9(8):1270–1274, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1270_s1.

Chakrabarti:1968:LBP

[1025] Amitabha Chakrabarti, Monique Levy-Nahas, and Roland Seneor. ‘Lorentz basis’ of the Poincaré group. Journal of Mathematical Physics, 9(8):1274–1283, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1274_s1.

Griffiths:1968:RSS

[1026] Robert B. Griffiths and J. L. Lebowitz. Random spin systems: Some rigorous results. Journal of Mathematical Physics, 9(8):1284–1292, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1284_s1.

Miller:1968:TCC

[1027] Kenneth J. Miller. Two-center Coulomb and hybrid integrals. Journal of Mathematical Physics, 9(8):1292–1294, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i8/p1292_s1.

Ansari:1968:QRCa

[1028] S. M. Razaullah Ansari. Quasibinomial representations of Clebsch– Gordan coefficients. I. Square symbol. Journal of Mathematical Physics, 9(8):1295–1299, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1295_s1.

Ansari:1968:QRCb

[1029] S. M. Razaullah Ansari. Quasibinomial representations of Clebsch– Gordan coefficients. II. “negative” representations. Journal of Mathematical Physics, 9(8):1299–1305, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i8/p1299_s1.

Rosen:1968:CIL

[1030] Joe Rosen. Construction of invariants for Lie algebras of inhomogeneous pseudo-orthogonal and pseudo-unitary groups. Journal of Mathematical Physics, 9(8):1305–1307, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v9/i8/p1305_s1.

Swan:1968:CDT

[1031] G. W. Swan. Current distribution in a thin superconducting strip. Journal of Mathematical Physics, 9(8):1308–1312, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i8/p1308_s1.

Casher:1968:FIM

[1032] A. Casher, D. Lurié, and M. Revzen. Functional integrals for manyboson systems. Journal of Mathematical Physics, 9(8):1312–1318, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1312_s1.

Vishveshwara:1968:GSS

[1033] C. V. Vishveshwara. Generalization of the “schwarzschild surface” to arbitrary static and stationary metrics. Journal of Mathematical Physics, 9(8):1319–1322, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i8/p1319_s1.

Pinski:1968:IPS

[1034] G. Pinski. Interaction of particlelike solutions in a classical nonlinear model field theory. Journal of Mathematical Physics, 9(8):1323–1326, August 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i8/p1323_s1.

Richardson:1968:ESM

[1035] R. W. Richardson. Exactly solvable many-boson model. Journal of Mathematical Physics, 9(9):1327–1343, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i9/p1327_s1.

Ehlers:1968:ISE

[1036] J. Ehlers, P. Geren, and R. K. Sachs. Isotropic solutions of the Einstein– Liouville equations. Journal of Mathematical Physics, 9(9):1344–1349, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1344_s1.

Woolcock:1968:ABS

[1037] W. S. Woolcock. Asymptotic behavior of Stieltjes transforms. II. Journal of Mathematical Physics, 9(9):1350–1356, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i9/p1350_s1.

Fontana:1968:OTC

[1038] Peter R. Fontana and William J. Meath. One- and two-center expansions of the Breit–Pauli Hamiltonian. Journal of Mathematical Physics, 9(9):1357–1364, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i9/p1357_s1.

Robinson:1968:SVP

[1039] P. D. Robinson and A. M. Arthurs. Some variational principles for integral equations. Journal of Mathematical Physics, 9(9):1364–1368, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1364_s1.

Sullivan:1968:APA

[1040] John J. Sullivan. Antisymmetric projection in the approximation of no spin-orbit coupling. Journal of Mathematical Physics, 9(9):1369–1374, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1369_s1.

Mehta:1968:ICG

[1041] M. L. Mehta and P. K. Srivastava. Irreducible corepresentations of groups having a compact simple Lie group as a subgroup of index 2. Journal of Mathematical Physics, 9(9):1375–1385, September 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i9/p1375_s1.

Whiteman:1968:INP

[1042] K. J. Whiteman and B. McNamara. Invariants of nearly periodic Hamiltonian systems. II. Journal of Mathematical Physics, 9(9):1385–1389, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1385_s1.

Fink:1968:AEF

[1043] James P. Fink. Asymptotic estimates of Feynman integrals. Journal of Mathematical Physics, 9(9):1389–1400, September 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i9/p1389_s1.

Pandres:1968:SPH

[1044] Dave Pandres and Donald A. Jacobson. Scalar product for harmonic functions of the group SU(2). Journal of Mathematical Physics, 9 (9):1401–1403, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i9/p1401_s1.

Speer:1968:AR

[1045] Eugene R. Speer. Analytic renormalization. Journal of Mathematical Physics, 9(9):1404–1410, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i9/p1404_s1.

Turner:1968:VQO

[1046] James E. Turner. Violation of the quantum ordering of propositions in hidden-variable theories. Journal of Mathematical Physics, 9(9):1411– 1415, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1411_s1.

Yndurain:1968:LCD

[1047] F. J. Ynduráin. Lorentz covariant distributions. Journal of Mathematical Physics, 9(9):1415–1422, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i9/p1415_s1.

Yndurain:1968:HTP

[1048] F. J. Ynduráin. “Haag Theorem” for the point-coupling relativistic Lee model. Journal of Mathematical Physics, 9(9):1423–1424, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1423_s1.

Rafanelli:1968:GSO

[1049] Kenneth Rafanelli. Generalized second-order relativistic wave equations. I. Journal of Mathematical Physics, 9(9):1425–1428, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1425_s1.

Favro:1968:ERM

[1050] L. D. Favro and J. F. McDonald. Ensemble of random matrices with a random bias. Journal of Mathematical Physics, 9(9):1429–1434, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1429_s1.

Miller:1968:SFCc

[1051] Willard Miller, Jr. Special functions and the complex Euclidean group in 3-space. III. Journal of Mathematical Physics, 9(9):1434–1444, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1434_s1.

Tiersten:1968:NBI

[1052] H. F. Tiersten. Natural boundary and initial conditions from a modification of Hamilton’s principle. Journal of Mathematical Physics, 9 (9):1445–1451, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i9/p1445_s1.

Sposito:1968:QEP

[1053] Garrison Sposito. Quantum-ergodic problem. Journal of Mathematical Physics, 9(9):1452–1456, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i9/p1452_s1.

Goldstein:1968:ISM

[1054] Rubin Goldstein. Iterative solutions by means of trial operators. Journal of Mathematical Physics, 9(9):1456–1461, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i9/p1456_s1.

Keam:1968:RBS

[1055] R. F. Keam. Reduction of the Bethe–Salpeter equation for a zero-mass system. Journal of Mathematical Physics, 9(9):1462–1467, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i9/p1462_s1.

Licht:1968:BDL

[1056] A. L. Licht. Background dependence of local states. Journal of Mathematical Physics, 9(9):1468–1475, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i9/p1468_s1.

Licht:1968:CSL

[1057] A. L. Licht. Coherent superposition of local states. Journal of Mathematical Physics, 9(9):1476–1480, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i9/p1476_s1.

Kraut:1968:EWP

[1058] Edgar A. Kraut. Elastic-wave propagation and the factorization of matrix-valued functions of several complex variables. Journal of Mathematical Physics, 9(9):1481–1487, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i9/p1481_s1.

Klauder:1968:NZM

[1059] L. T. Klauder, Jr. and J. G. Gay. Note on Zak’s method for constructing representations of space groups. Journal of Mathematical Physics, 9 (9):1488–1489, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i9/p1488_s1.

Atkinson:1968:DSV

[1060] D. Atkinson and A. P. Contogouris. Diffraction scattering and vectormeson resonances. Journal of Mathematical Physics, 9(9):1489–1500, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1489_s1.

Cornille:1968:TBR

[1061] H. Cornille and G. Rubinstein. Threshold-behavior requirements for P - wave N/D equations. Journal of Mathematical Physics, 9(9):1501–1516, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i9/p1501_s1.

Novozhilov:1968:URG

[1062] Yuri V. Novozhilov and Igor A. Terentjev. Unitary representations of the generalized Poincaré groups. Journal of Mathematical Physics, 9 (10):1517–1526, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i10/p1517_s1.

Burke:1968:NSM

[1063] T. Burke and J. L. Lebowitz. Note on statistical mechanics of random systems. Journal of Mathematical Physics, 9(10):1526–1534, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i10/p1526_s1.

Hill:1968:QCS

[1064] Robert Nyden Hill. Quantum corrections to the second virial coefficient at high temperatures. Journal of Mathematical Physics, 9(10):1534– 1547, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i10/p1534_s1.

Stapp:1968:CHA

[1065] Henry P. Stapp. Crossing, Hermitian analyticity, and the connection between spin and statistics. Journal of Mathematical Physics, 9(10):1548– 1592, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i10/p1548_s1.

Rosen:1968:TID

[1066] S. P. Rosen. TCP invariance and the dimensionality of space–time. Journal of Mathematical Physics, 9(10):1593–1594, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1593_s1.

Grgin:1968:NPD

[1067] Emile Grgin. Null and pseudonull data for scalar fields. Journal of Mathematical Physics, 9(10):1595–1601, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1595_s1.

Bialynicki-Birula:1968:GFI

[1068] I. Bialynicki-Birula and J. Piasecki. Green’s functions for the Ising chain. Journal of Mathematical Physics, 9(10):1602–1605, October 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1602_
s1.

Bacry:1968:PK

[1069] Henri Bacry and Jean-Marc Lévy-Leblond. Possible kinematics. Journal of Mathematical Physics, 9(10):1605–1614, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1605_s1.

Yao:1968:UIR

[1070] Tsu Yao. Unitary irreducible representations of SU(2, 2). II. Journal of Mathematical Physics, 9(10):1615–1626, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1615_s1.

Mavromatis:1968:IFP

[1071] H. A. Mavromatis and K. Schilcher. Inverse functions of the products of two Bessel functions and applications to potential scattering. Journal of Mathematical Physics, 9(10):1627–1632, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1627_s1.

Johnson:1968:HEF

[1072] Porter Johnson. High-energy fixed-angle potential scattering. Journal of Mathematical Physics, 9(10):1633–1638, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1633_s1.

Doebner:1968:LDG

[1073] H. D. Doebner and O. Melsheimer. Limitable dynamical groups in quantum mechanics. I. General theory and a spinless model. Journal of Mathematical Physics, 9(10):1638–1656, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i10/p1638_s1.

Matzner:1968:ASS

[1074] Richard A. Matzner. Almost symmetric spaces and gravitational radiation. Journal of Mathematical Physics, 9(10):1657–1668, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i10/p1657_s1.

Klink:1968:MPC

[1075] William H. Klink. Multiplicity problem for compact subgroups of noncompact groups. Journal of Mathematical Physics, 9(10):1669–1679, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i10/p1669_s1.

McQuistan:1968:EOK

[1076] R. B. McQuistan and D. Lichtman. Exact occupation kinetics for onedimensional arrays of dumbbells. Journal of Mathematical Physics, 9 (10):1680–1684, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i10/p1680_s1.

Ehrhardt:1968:CUM

[1077] P. O. G. Ehrhardt and D. B. Fairlie. Crossing and unitarity in a multichannel static model. I. Journal of Mathematical Physics, 9(10):1685– 1689, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i10/p1685_s1.

deAzcarraga:1968:PGI

[1078] José A. de Azcárraga and Luis J. Boya. Poincaré group and the invariant relativistic equations for massive particles of any spin. Journal of Mathematical Physics, 9(10):1689–1693, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1689_s1.

Katz:1968:RAE

[1079] Amnon Katz. Retarded and advanced electromagnetic fields in Friedmann universes. Journal of Mathematical Physics, 9(10):1694–1700, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i10/p1694_s1.

Carroll:1968:EOEa

[1080] M. M. Carroll and R. S. Rivlin. Electro-optical effects. II. Journal of Mathematical Physics, 9(10):1701–1704, October 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i10/p1701_s1.

Kim:1968:MAP

[1081] Shoon Kyung Kim. Modified admittance, perpendicular susceptibilities, and transformation of correlations of the Ising model. Journal of Mathematical Physics, 9(10):1705–1714, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i10/p1705_s1.

Datta:1968:SEF

[1082] B. K. Datta and A. K. Raychaudhuri. Stationary electromagnetic fields in general relativity. Journal of Mathematical Physics, 9(10):1715–1718, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i10/p1715_s1.

Federbush:1968:LOA

[1083] Paul Federbush. Local operator algebras in the presence of superselection rules. Journal of Mathematical Physics, 9(10):1718–1721, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i10/p1718_s1.

Abu-Shumays:1968:GFEa

[1084] I. K. Abu-Shumays and E. H. Bareiss. Generating functions for the exact solution of the transport equation. I. Journal of Mathematical Physics, 9(10):1722–1731, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i10/p1722_s1.

Kahane:1968:ARC

[1085] Joseph Kahane. Algorithm for reducing contracted products of γ matrices. Journal of Mathematical Physics, 9(10):1732–1738, October 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i10/
p1732_s1.

Geroch:1968:SSS

[1086] Robert Geroch. Spinor structure of space–times in general relativity. I. Journal of Mathematical Physics, 9(11):1739–1744, November 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1739_
s1.

Harrison:1968:NSE

[1087] B. Kent Harrison. New solutions of the Einstein–Maxwell equations from old. Journal of Mathematical Physics, 9(11):1744–1752, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i11/p1744_s1.

Bowden:1968:SER

[1088] Robert L. Bowden and Norman R. Richardson. Solution of the equations of Radiative Transfer in a free-electron atmosphere. Journal of Mathematical Physics, 9(11):1753–1761, November 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1753_s1.

Finkelstein:1968:CBS

[1089] David Finkelstein and Julio Rubinstein. Connection between spin, statistics, and kinks. Journal of Mathematical Physics, 9(11):1762–1779, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1762_s1.

Kobe:1968:PCDa

[1090] Donald H. Kobe. Principle of compensation of dangerous diagrams for boson systems. I. Maximum overlap. Journal of Mathematical Physics, 9(11):1779–1794, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i11/p1779_s1.

Kobe:1968:PCDb

[1091] Donald H. Kobe. Principle of compensation of dangerous diagrams for boson systems. II. Minimum quasiparticle number. Journal of Mathematical Physics, 9(11):1795–1805, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i11/p1795_s1.

Hakim:1968:RSP

[1092] Rémi Hakim. Relativistic stochastic processes. Journal of Mathematical Physics, 9(11):1805–1818, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i11/p1805_s1.

Dunnett:1968:IMC

[1093] D. A. Dunnett, E. W. Laing, and J. B. Taylor. Invariants in the motion of a charged particle in a spatially modulated magnetic field. Journal of Mathematical Physics, 9(11):1819–1823, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1819_s1.

McKerrell:1968:KSH

[1094] A. McKerrell. Kinematic singularities of helicity and transversity amplitudes and asymptotic Regge-pole contributions. Journal of Mathematical Physics, 9(11):1824–1832, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i11/p1824_s1.

Roth:1968:SOI

[1095] Charles Roth. Spin-orbit interactions of the configuration ℓ^{n−1}ℓ′ℓ′′. Journal of Mathematical Physics, 9(11):1832–1834, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1832_s1.

Carruthers:1968:LFT

[1096] P. Carruthers. Local field theory and isospin invariance. II. Free field theory of arbitrary spin particles. Journal of Mathematical Physics, 9 (11):1835–1845, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i11/p1835_s1.

Levine:1968:VAM

[1097] A. D. Levine and C. F. Hayes. Variational approach to multiple scattering. Journal of Mathematical Physics, 9(11):1845–1847, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i11/p1845_s1.

Atkinson:1968:NRS

[1098] D. Atkinson. Narrow resonance saturation of sum rules: A nonuniqueness theorem. Journal of Mathematical Physics, 9(11):1848–1864, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i11/p1848_s1.

Compton:1968:OTD

[1099] R. T. Compton, Jr. One- and two-dimensional Green’s functions for electromagnetic waves in moving simple media. Journal of Mathematical Physics, 9(11):1865–1872, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i11/p1865_s1.

Williams:1968:DLCa

[1100] M. M. R. Williams. Diffusion length and criticality problems in twoand three-dimensional, one-speed neutron transport theory. I. Rectangular coordinates. Journal of Mathematical Physics, 9(11):1873–1885, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1873_s1.

Williams:1968:DLCb

[1101] M. M. R. Williams. Diffusion length and criticality problems in twoand three-dimensional, one-speed neutron transport theory. II. Circular cylindrical coordinates. Journal of Mathematical Physics, 9(11):1885– 1889, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1885_s1.

Frank:1968:SCL

[1102] William M. Frank. Strong-coupling limit in potential theory. II. Journal of Mathematical Physics, 9(11):1890–1898, November 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1890_

s1. See erratum [1405].

Chadan:1968:DNS

[1103] K. Chadan and A. Montes. Derivation of nonrelativistic sum rules from the causality condition of Wigner and Van Kampen. Journal of Mathematical Physics, 9(11):1898–1914, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i11/p1898_s1.

Canosa:1968:ABC

[1104] José Canosa and Julian Cole. Asymptotic behavior of certain nonlinear boundary-value problems. Journal of Mathematical Physics, 9(11):1915– 1921, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1915_s1.

Chappell:1968:MAK

[1105] Willard R. Chappell. Microscopic approach to kinetic theory: Inhomogeneous systems. Journal of Mathematical Physics, 9(11):1921–1926, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1921_s1.

Pinski:1968:GTC

[1106] G. Pinski. Galilean tensor calculus. Journal of Mathematical Physics, 9(11):1927–1930, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i11/p1927_s1.

Jasperse:1968:ATB

[1107] J. R. Jasperse. Approach to the three-body scattering problem. Journal of Mathematical Physics, 9(11):1931–1936, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i11/p1931_s1.

Delbourgo:1968:RFD

[1108] R. Delbourgo. Representation functions for the degenerate baryon series in U(6) ⊗ U(6). Journal of Mathematical Physics, 9(11):1936–1939, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i11/p1936_s1.

Tross:1968:CMFa

[1109] Ralph G. Tross and Louis H. Lund. Cell model of a fluid. I. Evaluation of the partition function and series expansions. Journal of Mathematical Physics, 9(11):1940–1956, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i11/p1940_s1.

Tross:1968:CMFb

[1110] Ralph G. Tross and Louis H. Lund. Cell model of a fluid. II. Thermodynamic properties of the system. Journal of Mathematical Physics, 9 (11):1957–1975, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i11/p1957_s1.

Lewis:1968:CEI

[1111] H. R. Lewis, Jr. Class of exact invariants for classical and quantum time-dependent harmonic oscillators. Journal of Mathematical Physics, 9(11):1976–1986, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i11/p1976_s1.

Carmeli:1968:RTD

[1112] Moshe Carmeli. Representations of the three-dimensional rotation group in terms of direction and angle of rotation. Journal of Mathematical Physics, 9(11):1987–1992, November 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i11/p1987_s1.

Abu-Shumays:1968:GFEb

[1113] I. K. Abu-Shumays and E. H. Bareiss. Generating functions for the exact solution of the transport equation. II. Time-dependent with anisotropic scattering. Journal of Mathematical Physics, 9(12):1993–2001, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p1993_s1.

Lalor:1968:IWP

[1114] Éamon Lalor. Inverse wave propagator. Journal of Mathematical Physics, 9(12):2001–2006, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i12/p2001_s1.

Gillis:1968:RMP

[1115] N. S. Gillis. Response of a many-particle system to quasistatic changes in volume. Journal of Mathematical Physics, 9(12):2007–2016, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p2007_s1.

Mitchell:1968:RDP

[1116] Thomas P. Mitchell and J. A. Burns. Relativistic dynamics of a point charge in a magnetic-monopole field. Journal of Mathematical Physics, 9(12):2016–2017, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i12/p2016_s1.

Ozizmir:1968:TRE

[1117] Ercüment Özizmir. Transmission and reflection of electromagnetic waves at the boundary of a relativistic collisionless plasma. Journal of Mathematical Physics, 9(12):2018–2028, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i12/p2018_s1.

Hegerfeldt:1968:PFS

[1118] G. C. Hegerfeldt, K. Kraus, and E. P. Wigner. Proof of the Fermion superselection rule without the assumption of time-reversal invariance. Journal of Mathematical Physics, 9(12):2029–2031, December 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2029_
s1.

Canosa:1968:NBV

[1119] José Canosa. Nonlinear boundary-value problems in one- and twodimensional composite domains. Journal of Mathematical Physics, 9 (12):2032–2036, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i12/p2032_s1.

Anderson:1968:VSN

[1120] N. Anderson and A. M. Arthurs. Variational solutions of nonlinear Poisson–Boltzmann boundary-value problems. Journal of Mathematical Physics, 9(12):2037–2038, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i12/p2037_s1.

Moses:1968:RRR

[1121] Harry E. Moses. Reduction of reducible representations of the Poincaré group to standard helicity representations. Journal of Mathematical Physics, 9(12):2039–2049, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i12/p2039_s1.

Newton:1968:DAD

[1122] Roger G. Newton. Determination of the amplitude from the differential cross section by unitarity. Journal of Mathematical Physics, 9(12):2050– 2055, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2050_s1. See addendum [1437].

Muraskin:1968:IHM

[1123] Murray Muraskin. Interpretation of Γ_{[jk]}^{i} in a Huygens model. Journal of Mathematical Physics, 9(12):2056–2060, December 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2056_s1.

Radhakrishnan:1968:CGS

[1124] D. Radhakrishnan. On the Clebsch–Gordan series of a semisimple Lie algebra. Journal of Mathematical Physics, 9(12):2061–2063, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p2061_s1.

Suzuki:1968:TIM

[1125] Masuo Suzuki. Theorems on the Ising model with general spin and phase transition. Journal of Mathematical Physics, 9(12):2064–2068, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p2064_s1.

Barston:1968:EPL

[1126] E. M. Barston. Eigenvalue problem for Lagrangian systems. III. Journal of Mathematical Physics, 9(12):2069–2075, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2069_s1.

Hoegh-Krohn:1968:AFS

[1127] Raphael Høegh-Krohn. Asymptotic fields in some models of quantum field theory. I. Journal of Mathematical Physics, 9(12):2075–2080, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2075_s1.

Chen:1968:GFE

[1128] K. C. Chen and J. L. Yen. Green’s functions for electromagnetic waves in moving lossy media. Journal of Mathematical Physics, 9(12):2081– 2086, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2081_s1.

Chakrabarti:1968:CRA

[1129] A. Chakrabarti. Class of representations of the IU(n) and IO(n) algebras and respective deformations to U(n, 1), O(n, 1). Journal of Mathematical Physics, 9(12):2087–2100, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i12/p2087_s1.

Kuriyan:1968:MAR

[1130] Jacob G. Kuriyan, N. Mukunda, and E. C. G. Sudarshan. Master analytic representation: Reduction of O(2, 1) in an O(1, 1) basis. Journal of Mathematical Physics, 9(12):2100–2108, December 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2100_s1.

Speisman:1968:AVS

[1131] Gerald Speisman. Auxiliary variables in statistical mechanics: Variational principle. Journal of Mathematical Physics, 9(12):2109–2120, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2109_s1.

Carmi:1968:FOP

[1132] Gideon Carmi. First-order phase transitions in quantum-Coulomb plasmas. Journal of Mathematical Physics, 9(12):2120–2131, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i12/
p2120_s1.

Sobel:1968:DEH

[1133] M. I. Sobel. Differential equations for Half-off-shell matrix elements. Journal of Mathematical Physics, 9(12):2132–2137, December 1968. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2132_
s1.

Aviran:1968:MES

[1134] A. Aviran and J. Zak. Matrix elements in systems with nonunitary symmetry. Journal of Mathematical Physics, 9(12):2138–2145, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p2138_s1.

Stoyanov:1968:MRL

[1135] D. Tz. Stoyanov and I. T. Todorov. Majorana representations of the Lorentz group and infinite-component fields. Journal of Mathematical Physics, 9(12):2146–2167, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i12/p2146_s1.

Risk:1968:AED

[1136] Clifford Risk. Analyticity of the envelope diagrams. Journal of Mathematical Physics, 9(12):2168–2172, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i12/p2168_s1.

Siegert:1968:IML

[1137] A. J. F. Siegert and David J. Vezzetti. On the Ising model with longrange interaction. Journal of Mathematical Physics, 9(12):2173–2193, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v9/i12/p2173_s1.

Kundt:1968:HDE

[1138] W. Kundt and E. T. Newman. Hyperbolic differential equations in two dimensions. Journal of Mathematical Physics, 9(12):2193–2210, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v9/i12/p2193_s1.

Lebrun:1968:CSP

[1139] Jean-Pierre Lebrun. Class of solutions of the physical 3-particle unitarity equations in static S-matrix theory. Journal of Mathematical Physics, 9(12):2211–2225, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i12/p2211_s1.

Luehr:1968:IST

[1140] C. P. Luehr and M. Rosenbaum. Intrinsic spinor techniques with applications to the Lorentz group and the Dirac equation. Journal of Mathematical Physics, 9(12):2225–2236, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i12/p2225_s1.

Wu:1968:TSA

[1141] Hsin i Wu and Paul W. Schmidt. Theory of the small-angle X-ray scattering from randomly oriented plane laminas. Journal of Mathematical Physics, 9(12):2237–2252, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v9/i12/p2237_s1.

Rubin:1968:TPI

[1142] Robert J. Rubin. Transmission properties of an isotopically disordered one-dimensional harmonic crystal. Journal of Mathematical Physics, 9 (12):2252–2266, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i12/p2252_s1.

Carroll:1968:EOEb

[1143] M. M. Carroll and R. S. Rivlin. Electro-optical effects. III. Journal of Mathematical Physics, 9(12):2267–2278, December 1968. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v9/i12/p2267_s1.

Landshoff:1968:CTM

[1144] P. V. Landshoff. Cayley transform of the S matrix. Journal of Mathematical Physics, 9(12):2279–2282, December 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v9/i12/p2279_s1.

Thorpe:1969:CPC

[1145] John A. Thorpe. Curvature and the Petrov canonical forms. Journal of Mathematical Physics, 10(1):1–7, January 1969. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i1/p1_s1.

Millard:1969:TIP

[1146] Kenneth Millard. Translational invariance properties of a finite onedimensional hard-core fluid using the grand canonical ensemble. Journal of Mathematical Physics, 10(1):7–13, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i1/p7_s1.

Lerche:1969:ITW

[1147] I. Lerche. Instability of transverse waves in a relativistic plasma. Journal of Mathematical Physics, 10(1):13–20, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i1/p13_s1.

Chang:1969:URB

[1148] Shau-Jin Chang and L. O’Raifeartaigh. Unitary representations of SL(2, C) in an E(2) basis. Journal of Mathematical Physics, 10(1):21– 29, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i1/p21_s1.

Zickendraht:1969:CSA

[1149] W. Zickendraht. Configuration-space approach to the four-particle problem. Journal of Mathematical Physics, 10(1):30–37, January 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p30_s1.

Penrose:1969:SZR

[1150] Roger Penrose. Solutions of the zero-rest-mass equations. Journal of Mathematical Physics, 10(1):38–39, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i1/p38_s1.

Dobson:1969:MNI

[1151] J. F. Dobson. Many-neighbored Ising chain. Journal of Mathematical Physics, 10(1):40–45, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i1/p40_s1.

Gamblin:1969:CCT

[1152] Rodger L. Gamblin. Comments on the classical theory of magnetic monopoles. Journal of Mathematical Physics, 10(1):46–48, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/
p46_s1.

Gould:1969:NTB

[1153] H. W. Gould. Note on two binomial coefficient identities of Rosenbaum. Journal of Mathematical Physics, 10(1):49–50, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p49_s1.

Federbush:1969:PAD

[1154] Paul Federbush. Partially alternate derivation of a result of Nelson. Journal of Mathematical Physics, 10(1):50–52, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p50_s1.

Antoine:1969:DFSa

[1155] J. P. Antoine. Dirac formalism and symmetry problems in quantum mechanics. I. General Dirac formalism. Journal of Mathematical Physics, 10(1):53–69, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i1/p53_s1.

Carter:1969:KHO

[1156] Brandon Carter. Killing horizons and orthogonally transitive groups in space–time. Journal of Mathematical Physics, 10(1):70–81, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i1/p70_s1.

Nussenzveig:1969:HFSa

[1157] H. M. Nussenzveig. High-frequency scattering by a transparent sphere. I. Direct reflection and transmission. Journal of Mathematical Physics, 10(1):82–124, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i1/p82_s1.

Nussenzveig:1969:HFSb

[1158] H. M. Nussenzveig. High-frequency scattering by a transparent sphere. II. Theory of the Rainbow and the Glory. Journal of Mathematical Physics, 10(1):125–176, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i1/p125_s1.

Majumdar:1969:PTS

[1159] Chanchal K. Majumdar. Problem of two spin deviations in a linear chain with next-nearest-neighbor interactions. Journal of Mathematical Physics, 10(1):177–180, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i1/p177_s1.

Zimering:1969:SAB

[1160] Shimshon Zimering. Some asymptotic behavior of Stieltjes transforms. Journal of Mathematical Physics, 10(1):181–183, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p181_s1.

Broido:1969:BSE

[1161] Michael M. Broido and John G. Taylor. Bethe–Salpeter equation. Journal of Mathematical Physics, 10(1):184–209, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p184_s1.

Kuszell:1969:KDI

[1162] A. Kuszell. Kinetic description of an inhomogeneous plasma. Journal of Mathematical Physics, 10(1):210–220, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i1/p210_s1.

Ciftan:1969:CSSa

[1163] Mikael Ciftan and L. C. Biedenharn. Combinatorial structure of state vectors in U_{n}. I. Hook patterns for maximal and semimaximal states in U_{n}. Journal of Mathematical Physics, 10(1):221–232, January 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i1/p221_
s1.

Summerfield:1969:RAG

[1164] G. C. Summerfield and P. F. Zweifel. Relationships among generalized phase-space distributions. Journal of Mathematical Physics, 10(2):233– 235, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i2/p233_s1.

Kleiner:1969:SGL

[1165] W. H. Kleiner and T. A. Kaplan. Symmetry of the ground level of a Hamiltonian. Journal of Mathematical Physics, 10(2):236–238, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i2/p236_s1.

Sirovich:1969:WPO

[1166] Lawrence Sirovich and James K. Thurber. Wave propagation and other spectral problems in kinetic theory. Journal of Mathematical Physics, 10(2):239–249, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i2/p239_s1.

Neuringer:1969:CFS

[1167] J. L. Neuringer. Closed-form solution of the differential equation (∂^{2}/(∂x∂y) + ax∂/∂x + by∂/∂y + cxy + ∂/∂t)P = 0 subject to the initial condition P (x, y, t = 0) = Φ(x, y). Journal of Mathematical Physics, 10(2):250–251, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i2/p250_s1.

Delaney:1969:IRM

[1168] R. M. Delaney and B. Gruber. Inner and restriction multiplicity for classical groups. Journal of Mathematical Physics, 10(2):252–265, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i2/p252_s1.

Duderstadt:1969:TPP

[1169] James J. Duderstadt. Theory of the propagation of the plane-wave disturbances in a distribution of thermal neutrons. Journal of Mathematical Physics, 10(2):266–276, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i2/p266_s1.

Cisneros:1969:STD

[1170] Arturo Cisneros and Harold V. McIntosh. Symmetry of the twodimensional hydrogen atom. Journal of Mathematical Physics, 10 (2):277–286, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i2/p277_s1.

Kaper:1969:ESR

[1171] Hans G. Kaper. Elementary solutions of the reduced three-dimensional transport equation. Journal of Mathematical Physics, 10(2):286–297, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i2/p286_s1.

Bakri:1969:SSF

[1172] Mahmoud M. Bakri. de Sitter symmetric field theory. I. One-particle theory. Journal of Mathematical Physics, 10(2):298–320, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i2/
p298_s1.

Mills:1969:IPS

[1173] Robert L. Mills and John F. Reading. Inversion problem with separable potentials. Journal of Mathematical Physics, 10(2):321–331, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i2/p321_s1.

Mirman:1969:CCR

[1174] R. Mirman. Classification and complete reducibility of representations of compact inhomogeneous groups and algebras. Journal of Mathematical Physics, 10(2):331–338, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i2/p331_s1.

Ehrhardt:1969:CUM

[1175] P. O. G. Ehrhardt and D. B. Fairlie. Crossing and unitarity in a multichannel static model. II. Journal of Mathematical Physics, 10(2):339– 341, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i2/p339_s1.

Millman:1969:PTN

[1176] Martin H. Millman and Joseph B. Keller. Perturbation theory of nonlinear boundary-value problems. Journal of Mathematical Physics, 10(2):342–361, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i2/p342_s1.

Morton:1969:ECA

[1177] J. B. Morton and S. Corrsin. Experimental confirmation of the applicability of the Fokker–Planck equation to a nonlinear oscillator. Journal of Mathematical Physics, 10(2):361–368, February 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i2/p361_s1.

Goldberg:1969:GGT

[1178] Joshua N. Goldberg and Ezra T. Newman. Generalization of Green’s theorem. Journal of Mathematical Physics, 10(2):369–370, February 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i2/p369_s1.

Coster:1969:PRD

[1179] Joseph Coster and Henry P. Stapp. Physical-region discontinuity equations for many-particle scattering amplitudes. I. Journal of Mathematical Physics, 10(3):371–396, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i3/p371_s1.

Hauge:1969:NAD

[1180] E. H. Hauge and E. G. D. Cohen. Normal and abnormal diffusion in Ehrenfests’ wind-tree model. Journal of Mathematical Physics, 10(3): 397–414, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p397_s1.

Rosen:1969:FIT

[1181] Gerald Rosen, J. A. Okolowski, and Gene Eckstut. Functional integration theory for incompressible fluid turbulence. II. Journal of Mathematical Physics, 10(3):415–421, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i3/p415_s1.

Deliyannis:1969:STD

[1182] Platon C. Deliyannis and Evangelos K. Ifantis. Spectral theory of the difference equation f(n + 1) + f(n − 1) = [E − ϕ(n)]f(n). Journal of Mathematical Physics, 10(3):421–425, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i3/p421_s1.

Coffey:1969:IAO

[1183] Timothy P. Coffey, Jr. Invariants to all orders in classical perturbation theory. Journal of Mathematical Physics, 10(3):426–438, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i3/p426_s1.

Chen:1969:ISR

[1184] Robert L. W. Chen and Thomas H. Neighbors. Isoperimetric solutions related to the thermodynamics of plasmas. Journal of Mathematical Physics, 10(3):439–442, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i3/p439_s1.

Hammersley:1969:NFT

[1185] J. M. Hammersley, A. Feuerverger, A. Izenman, and K. Makani. Negative finding for the three-dimensional dimer problem. Journal of Mathematical Physics, 10(3):443–446, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i3/p443_s1.

Wait:1969:AMC

[1186] James R. Wait. Approach to the mode conversion problem in nonuniform acoustic waveguides. Journal of Mathematical Physics, 10(3):447– 451, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p447_s1.

Boillat:1969:RVE

[1187] Guy Boillat. Ray velocity and exceptional waves: A covariant formulation. Journal of Mathematical Physics, 10(3):452–454, March 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i3/p452_
s1.

Moshinsky:1969:GAF

[1188] M. Moshinsky and V. Syamala Devi. General approach to fractional parentage coefficients. Journal of Mathematical Physics, 10(3):455– 466, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p455_s1.

Wybourne:1969:HRL

[1189] B. G. Wybourne. Hermite’s reciprocity law and the angular-momentum states of equivalent particle configurations. Journal of Mathematical Physics, 10(3):467–471, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i3/p467_s1.

Robinson:1969:CBG

[1190] Peter D. Robinson. Complementary bounds for ground-state energies. Journal of Mathematical Physics, 10(3):472–474, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i3/p472_s1.

Baroody:1969:FSM

[1191] E. M. Baroody. Frequency spectrum and momentum autocorrelation function for a simple lattice. Journal of Mathematical Physics, 10(3): 475–479, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p475_s1.

Ting:1969:TSF

[1192] Chung-Yu Ting. Theoretical study of finite dielectric-coated cylindrical antenna. Journal of Mathematical Physics, 10(3):480–493, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i3/p480_s1.

Bloxham:1969:MSSa

[1193] M. J. W. Bloxham, D. I. Olive, and J. C. Polkinghorne. S-matrix singularity structure in the physical region. I. Properties of multiple integrals. Journal of Mathematical Physics, 10(3):494–502, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i3/p494_s1.

Spector:1969:HES

[1194] Richard M. Spector. High-energy scattering cross section for singular potentials at large angles. Journal of Mathematical Physics, 10(3):503– 505, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p503_s1.

Moyal:1969:MET

[1195] J. E. Moyal. Mean ergodic theorems in quantum mechanics. Journal of Mathematical Physics, 10(3):506–509, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i3/p506_s1.

Zimmerman:1969:BSN

[1196] Robert L. Zimmerman. Bound states of a new kind in quantum field theory. Journal of Mathematical Physics, 10(3):509–518, March 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i3/p509_
s1.

Maduemezia:1969:PWA

[1197] Awele Maduemezia. Partial-wave analysis in terms of SL(2, C)/SU(2) harmonic functions. Journal of Mathematical Physics, 10(3):519–523, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i3/p519_s1.

Hahne:1969:ACA

[1198] G. E. Hahne. Analytic continuation of Appell’s hypergeometric series F_{2} to the vicinity of the singular point x = 1, y = 1. Journal of Mathematical Physics, 10(3):524–531, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i3/p524_s1.

Holland:1969:FT

[1199] Douglas Francis Holland. Finite transformations of SU_{3}. Journal of Mathematical Physics, 10(3):531–535, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i3/p531_s1.

Su:1969:KVE

[1200] C. H. Su and C. S. Gardner. Korteweg–de Vries equation and generalizations. III. Derivation of the Korteweg–de Vries equation and Burgers equation. Journal of Mathematical Physics, 10(3):536–539, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i3/p536_s1.

vanderMerwe:1969:SOD

[1201] Johannes H. van der Merwe and Alwyn J. van der Merwe. Secondorder dispersion-energy series for axially symmetric molecules. Journal of Mathematical Physics, 10(3):539–544, March 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i3/p539_s1.

Bloxham:1969:MSSb

[1202] M. J. W. Bloxham, D. I. Olive, and J. C. Polkinghorne. S-matrix singularity structure in the physical region. II. Unitarity integrals. Journal of Mathematical Physics, 10(4):545–552, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i4/p545_s1.

Bloxham:1969:MSSc

[1203] M. J. W. Bloxham, D. I. Olive, and J. C. Polkinghorne. S-matrix singularity structure in the physical region. III. General discussion of simple Landau singularities. Journal of Mathematical Physics, 10(4):553–561, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p553_s1.

Calogero:1969:LBG

[1204] F. Calogero and C. Marchioro. Lower bounds to the ground-state energy of systems containing identical particles. Journal of Mathematical Physics, 10(4):562–569, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p562_s1. See comment [1436].

Carmeli:1969:GTA

[1205] Moshe Carmeli. Group-theoretic approach to the new conserved quantities in general relativity. Journal of Mathematical Physics, 10(4):569– 574, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p569_s1.

Capri:1969:EGT

[1206] Anton Z. Capri. Electron in a given time-dependent electromagnetic field. Journal of Mathematical Physics, 10(4):575–580, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p575_s1.

Pahor:1969:IIC

[1207] S. Pahor and P. F. Zweifel. Invariant imbedding and Case eigenfunctions. Journal of Mathematical Physics, 10(4):581–589, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p581_s1.

Meetz:1969:RCS

[1208] Kurt Meetz. Realization of chiral symmetry in a curved isospin space. Journal of Mathematical Physics, 10(4):589–593, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p589_s1.

Keam:1969:SBS

[1209] R. F. Keam. Solution of the Bethe–Salpeter equation with a squarewell potential. Journal of Mathematical Physics, 10(4):594–603, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p594_s1.

Howgate:1969:MEC

[1210] David W. Howgate. Method for the evaluation of certain time-dependent thermal averages. Journal of Mathematical Physics, 10(4):604–606, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p604_s1.

Klink:1969:URP

[1211] William H. Klink. Unitary representations of the Poincaré Group contained in a class of representations of the conformal group. Journal of Mathematical Physics, 10(4):606–613, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i4/p606_s1.

Gomatam:1969:PF

[1212] Jagannathan Gomatam and F. Rohrlich. Powers of the D functions. Journal of Mathematical Physics, 10(4):614–616, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p614_s1.

Katzin:1969:CCF

[1213] Gerald H. Katzin, Jack Levine, and William R. Davis. Curvature collineations: A fundamental symmetry property of the space–times of general relativity defined by the vanishing Lie derivative of the Riemann curvature tensor. Journal of Mathematical Physics, 10(4):617–629, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p617_s1.

Peres:1969:RFM

[1214] Asher Peres. Renormalization of a finite matrix Hamiltonian. Journal of Mathematical Physics, 10(4):629–632, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i4/p629_s1.

Tauber:1969:MVM

[1215] G. E. Tauber. Massive vector meson interacting with the gravitational field. I. General formalism. Journal of Mathematical Physics, 10(4):633– 638, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p633_s1.

Hoegh-Krohn:1969:AFS

[1216] Raphael Høegh-Krohn. Asymptotic fields in some models of quantum field theory. II. Journal of Mathematical Physics, 10(4):639–643, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p639_s1.

Eyges:1969:GSM

[1217] Leonard Eyges. Ground-state many-boson problem with repulsive potentials. Journal of Mathematical Physics, 10(4):644–654, April 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p644_
s1.

Ross:1969:SAP

[1218] William R. Ross and John R. Taylor. Some analytic properties of scattering amplitudes for long-range forces. Journal of Mathematical Physics, 10(4):654–660, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p654_s1.

Brittin:1969:FMS

[1219] W. E. Brittin and W. R. Chappell. Functional methods in statistical mechanics. I. Classical theory. Journal of Mathematical Physics, 10(4):

661–674, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p661_s1.

Blomberg:1969:FIR

[1220] Clas Blomberg. Functional integrals representing distribution functions in statistical mechanics. Journal of Mathematical Physics, 10(4):675– 682, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p675_s1.

Ishidzu:1969:AMS

[1221] Takehiko Ishidzu, Hideo Kawarada, and Masachiyo Sato. SU_{3} algebra in the mixed-shell space (0p, 0d, 1s). I. Journal of Mathematical Physics, 10(4):683–688, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p683_s1.

Watson:1969:MSE

[1222] Kenneth M. Watson. Multiple scattering of electromagnetic waves in an underdense plasma. Journal of Mathematical Physics, 10(4):688– 702, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p688_s1.

Yoseloff:1969:CAR

[1223] Mark L. Yoseloff and Harold W. Kuhn. Combinatorial approach to the N-representability of P -density matrices. Journal of Mathematical Physics, 10(4):703–706, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p703_s1.

Kahn:1969:LDF

[1224] Peter B. Kahn and N. Rosenzweig. Level density of a Fermi system: Nonperiodic perturbations of the energy-level scheme. Journal of Mathematical Physics, 10(4):707–715, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i4/p707_s1.

Northover:1969:SCV

[1225] F. H. Northover. Sufficiency condition for the validity of the WKB approximation. Journal of Mathematical Physics, 10(4):715–724, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p715_s1.

Davidson:1969:LID

[1226] Ernest R. Davidson. Linear inequalities for density matrices. Journal of Mathematical Physics, 10(4):725–734, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i4/p725_s1.

Stabell:1969:FOA

[1227] R. Stabell. First-order approximation to a spherically symmetric solution of Einstein’s equations. Journal of Mathematical Physics, 10(4):735– 739, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p735_s1.

Harter:1969:ATR

[1228] William G. Harter. Algebraic theory of ray representations of finite groups. Journal of Mathematical Physics, 10(4):739–752, April 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p739_
s1.

Montroll:1969:RWL

[1229] Elliott W. Montroll. Random walks on lattices. III. Calculation of firstpassage times with application to exciton trapping on photosynthetic units. Journal of Mathematical Physics, 10(4):753–765, April 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i4/p753_
s1.

Thompson:1969:ACF

[1230] James Thompson and Glen Schrank. Algebraic classification of fourdimensional Riemann spaces. Journal of Mathematical Physics, 10(4): 766–770, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i4/p766_s1.

Smith:1969:PES

[1231] Vedene H. Smith, Jr. and Frank E. Harris. Projection of exact spin eigenfunctions. Journal of Mathematical Physics, 10(4):771–778, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i4/p771_s1.

Lee:1969:ICM

[1232] Huan Lee. On the isospin crossing matrix. Journal of Mathematical Physics, 10(4):779–787, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p779_s1.

Carter:1969:PSS

[1233] Benjamin P. Carter. Partial separation of the Schrödinger equation for two charged particles in a magnetic field. Journal of Mathematical Physics, 10(4):788–791, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p788_s1.

Ruhl:1969:CFS

[1234] W. Rühl. Classical fields on spacelike mass shells. Journal of Mathematical Physics, 10(4):791–798, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i4/p791_s1.

Cohn:1969:CCS

[1235] Jack Cohn. Considerations on the classical spinning electron. Journal of Mathematical Physics, 10(5):799–805, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i5/p799_s1.

Levy-Leblond:1969:NGF

[1236] Jean-Marc Lévy-Leblond. Nonsaturation of gravitational forces. Journal of Mathematical Physics, 10(5):806–812, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i5/p806_s1.

Bose:1969:ZMR

[1237] S. K. Bose and R. Parker. Zero-mass representation of Poincaré group and conformal invariance. Journal of Mathematical Physics, 10(5):812– 813, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i5/p812_s1.

Baker:1969:BEB

[1238] George A. Baker, Jr. Best error bounds for Padé approximants to convergent series of Stieltjes. Journal of Mathematical Physics, 10(5):814–820, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i5/p814_s1.

Navon:1969:CSD

[1239] A. Navon. Criteria for Slater determinants and quasiparticle vacuum states. Journal of Mathematical Physics, 10(5):821–826, May 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p821_
s1.

Chandler:1969:MCC

[1240] Colston Chandler and Henry P. Stapp. Macroscopic causality conditions and properties of scattering amplitudes. Journal of Mathematical Physics, 10(5):826–859, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i5/p826_s1.

Dunne:1969:AKT

[1241] S. A. Dunne. Application of the Kirillov theory to the representations of O(2, 1). Journal of Mathematical Physics, 10(5):860–868, May 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p860_
s1.

Fan:1969:SDP

[1242] Chungpeng Fan. Solution of the dimer problem by the S-matrix method. Journal of Mathematical Physics, 10(5):868–871, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p868_s1.

Gamba:1969:LLA

[1243] A. Gamba. Lie-like approach to the theory of representations of finite groups. Journal of Mathematical Physics, 10(5):872–874, May 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p872_
s1.

Gibbs:1969:ASN

[1244] Alan G. Gibbs. Analytical solutions of the Neutron transport equation in arbitrary convex geometry. Journal of Mathematical Physics, 10(5): 875–890, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i5/p875_s1.

Casey:1969:WPS

[1245] K. F. Casey, J. R. Matthes, and C. Yeh. Wave propagation in sinusoidally stratified plasma media. Journal of Mathematical Physics, 10(5): 891–897, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i5/p891_s1.

Mukunda:1969:ELG

[1246] N. Mukunda. Expansion of Lie groups and representations of SL(3, C). Journal of Mathematical Physics, 10(5):897–911, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p897_s1.

Molyneux:1969:SPP

[1247] John E. Molyneux. Statistical properties of polycrystalline dielectrics. Journal of Mathematical Physics, 10(5):912–917, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i5/p912_s1.

Lukierski:1969:ART

[1248] Jerzy Lukierski. Analytic representations of two-point functions with noncanonical light-cone singularities. I. Journal of Mathematical Physics, 10(5):918–924, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i5/p918_s1.

Moravcsik:1969:DIA

[1249] Michael J. Moravcsik and Wing-Yin Yu. Determination of invariant amplitudes from experimental observables. Journal of Mathematical Physics, 10(5):925–928, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i5/p925_s1.

Cawley:1969:HDT

[1250] Robert G. Cawley. Hamilton–Dirac theory of Hamilton’s equations. Journal of Mathematical Physics, 10(5):928–932, May 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i5/p928_s1.

Prugovecki:1969:MTD

[1251] Eduard Prugovecki. Measure-theoretical description of Klein–Gordon multiparticle states. Journal of Mathematical Physics, 10(5):933–945, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i5/p933_s1.

Johnson:1969:OBP

[1252] Porter W. Johnson. Orear behavior in potential scattering. II. Journal of Mathematical Physics, 10(5):946–953, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i5/p946_s1.

Hoffman:1969:SNS

[1253] Richard B. Hoffman. Stationary “noncanonical” solutions of the Einstein vacuum field equations. Journal of Mathematical Physics, 10(5): 953–956, May 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i5/p953_s1.

Delbourgo:1969:DRF

[1254] R. Delbourgo, K. Koller, and Ruth M. Williams. Degenerate representation functions for SO(ν), SU(ν), and SU(ν) ⊗ SU(ν) and their analytical reductions. Journal of Mathematical Physics, 10(6):957–963, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i6/p957_s1.

Resibois:1969:CKE

[1255] P. Résibois, J. Brocas, and G. Decan. Comments on kinetic equation for autocorrelation functions. Journal of Mathematical Physics, 10(6): 964–974, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i6/p964_s1.

Chow:1969:GKC

[1256] Yutze Chow. Gel’fand–Kirillov conjecture on the Lie field of an algebraic Lie algebra. Journal of Mathematical Physics, 10(6):975–992, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i6/p975_s1.

Girardeau:1969:BSG

[1257] M. D. Girardeau. Broken symmetry and generalized Bose condensation in restricted geometries. Journal of Mathematical Physics, 10(6): 993–998, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i6/p993_s1.

Coffey:1969:NP

[1258] Timothy P. Coffey and G. W. Ford. Nonlinear perturbations. Journal of Mathematical Physics, 10(6):998–1003, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i6/p998_s1.

Gutzwiller:1969:PIA

[1259] Martin C. Gutzwiller. Phase-integral approximation in momentum space and the bound states of an atom. II. Journal of Mathematical Physics, 10(6):1004–1020, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i6/p1004_s1.

Felderhof:1969:DNE

[1260] B. U. Felderhof. Dynamics near equilibrium of systems described in thermal Hartree–Fock approximation. Journal of Mathematical Physics, 10(6):1021–1025, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i6/p1021_s1.

Gibberd:1969:NNN

[1261] R. W. Gibberd. Next-nearest-neighbor Ising model. Journal of Mathematical Physics, 10(6):1026–1029, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i6/p1026_s1.

Peters:1969:GBD

[1262] P. C. Peters. Geometrization of the Brans–Dicke scalar field. Journal of Mathematical Physics, 10(6):1029–1031, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i6/p1029_s1.

Shibuya:1969:EDM

[1263] Tai ichi Shibuya and Oktay Sinanoglu. Expansion of the density matrix of an N-fermion system in terms of the correlation densities of fermion clusters. Journal of Mathematical Physics, 10(6):1032–1037, June 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i6/p1032_
s1.

Roth:1969:AMC

[1264] C. Roth. Algebraic matrices for the configurations (d + s)^{n}p. Journal of Mathematical Physics, 10(6):1038–1044, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i6/p1038_s1.

Garrido:1969:GIP

[1265] L. M. Garrido. General interaction picture from action principle for mechanics. Journal of Mathematical Physics, 10(6):1045–1056, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i6/p1045_s1.

Lanczos:1969:QAP

[1266] C. Lanczos. Quadratic action principle of relativity. Journal of Mathematical Physics, 10(6):1057–1065, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i6/p1057_s1.

Wong:1969:ROG

[1267] M. K. F. Wong. Representations of the orthogonal group. II. Polynomial bases for the irreducible representations of the orthogonal group. Journal of Mathematical Physics, 10(6):1065–1068, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i6/p1065_s1.

Green:1969:CDP

[1268] H. S. Green and T. Triffet. Co-diagonal perturbations. Journal of Mathematical Physics, 10(6):1069–1089, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i6/p1069_s1.

Davies:1969:LDS

[1269] John A. Davies and Peter D. Yedinak. Lattice dynamics of simple cubic lattices with long-range interactions. Journal of Mathematical Physics, 10(6):1090–1097, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i6/p1090_s1.

Zachary:1969:SCS

[1270] W. W. Zachary. Sufficient conditions for stability of the Faddeev equations. Journal of Mathematical Physics, 10(6):1098–1105, June 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i6/p1098_
s1.

Glasser:1969:EKI

[1271] M. L. Glasser. Evaluation of “kondo” integrals. Journal of Mathematical Physics, 10(6):1105–1106, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i6/p1105_s1.

Mattis:1969:MIM

[1272] Daniel C. Mattis and Michael Plischke. Magnetization of Ising model in nonzero magnetic field. Journal of Mathematical Physics, 10(6):1107– 1108, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i6/p1107_s1.

Itzykson:1969:GRC

[1273] C. Itzykson. Group representation in a continuous basis: An example. Journal of Mathematical Physics, 10(6):1109–1114, June 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i6/p1109_s1.

Yang:1969:TOD

[1274] C. N. Yang and C. P. Yang. Thermodynamics of a one-dimensional system of bosons with repulsive delta-function interaction. Journal of Mathematical Physics, 10(7):1115–1122, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i7/p1115_s1.

Simon:1969:GNB

[1275] Barry Simon. On the growth of the number of bound states with increase in potential strength. Journal of Mathematical Physics, 10(7): 1123–1126, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i7/p1123_s1.

Rosenbaum:1969:SHS

[1276] David M. Rosenbaum. Super Hilbert space and the quantum-mechanical time operators. Journal of Mathematical Physics, 10(7):1127–1144, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1127_s1.

Vajk:1969:ERW

[1277] J. Peter Vajk. Exact Robertson–Walker cosmological solutions containing relativistic fluids. Journal of Mathematical Physics, 10(7):1145–1151, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1145_s1.

Kulhanek:1969:EBW

[1278] J. Kulhanek. Equations of the de Broglie wavefield in the case of spherical symmetry. Journal of Mathematical Physics, 10(7):1152–1156, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1152_s1.

Besieris:1969:AER

[1279] Ioannis M. Besieris. Analysis of electromagnetic radiation in the presence of a uniformly moving, uniaxially anisotropic medium. Journal of Mathematical Physics, 10(7):1156–1167, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i7/p1156_s1.

Brandt:1969:GBO

[1280] Richard A. Brandt and O. W. Greenberg. Generalized Bose operators in the Fock space of a single Bose operator. Journal of Mathematical Physics, 10(7):1168–1176, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i7/p1168_s1.

Shampine:1969:ESC

[1281] L. F. Shampine. Existence of solutions for certain nonlinear boundaryvalue problems. Journal of Mathematical Physics, 10(7):1177–1178, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1177_s1.

Harrison:1969:ACC

[1282] Don E. Harrison, Jr., Warren L. Gay, and H. M. Effron. Algorithm for the calculation of the classical equations of motion of an N-body system. Journal of Mathematical Physics, 10(7):1179–1184, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i7/p1179_s1.

Hajj:1969:WEM

[1283] Jack P. Hajj and David L. Weaver. Wave equations, multipole solutions, and characteristic hypersurfaces for massless particles. Journal of Mathematical Physics, 10(7):1185–1190, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i7/p1185_s1.

McDonald:1969:WDC

[1284] J. F. McDonald. On the width distribution for a complex system whose Hamiltonian contains a small interaction that is odd under time-reversal. Journal of Mathematical Physics, 10(7):1191–1194, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i7/p1191_s1.

Kinnersley:1969:TVM

[1285] William Kinnersley. Type D vacuum metrics. Journal of Mathematical Physics, 10(7):1195–1203, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i7/p1195_s1.

Carter:1969:CGC

[1286] J. C. Carter and J. J. Coyne. SU(6) Clebsch–Gordan coefficients for the product 35 ⊗ 70. Journal of Mathematical Physics, 10(7):1204–1210, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1204_s1.

Baxter:1969:MTL

[1287] R. J. Baxter. F model on a triangular lattice. Journal of Mathematical Physics, 10(7):1211–1216, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i7/p1211_s1.

Peters:1969:CEP

[1288] P. C. Peters. Covariant electromagnetic potentials and fields in Friedmann universes. Journal of Mathematical Physics, 10(7):1216–1224, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1216_s1.

Kim:1969:IMB

[1289] Shoon Kyung Kim. Involutional matrices based on the representation theory of GL (2). Journal of Mathematical Physics, 10(7):1225–1234, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i7/p1225_s1.

McDonald:1969:HHS

[1290] S. C. McDonald. Hyperplane helicity states. Journal of Mathematical Physics, 10(7):1234–1240, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i7/p1234_s1.

Iagolnitzer:1969:MCD

[1291] Daniel Iagolnitzer. S-matrix and classical description of interactions. Journal of Mathematical Physics, 10(7):1241–1264, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i7/p1241_s1.

Gervais:1969:PWE

[1292] J. L. Gervais and G. Velo. Partial-wave expansion in the crossed channel for scattering amplitudes invariant under the Galilei group. Journal of Mathematical Physics, 10(7):1265–1270, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i7/p1265_s1.

Dalton:1969:CPB

[1293] N. W. Dalton and D. W. Wood. Critical point behavior of the Ising model with higher-neighbor interactions present. Journal of Mathematical Physics, 10(7):1271–1302, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i7/p1271_s1.

Girardeau:1969:PSP

[1294] M. D. Girardeau. Proof of the symmetrization postulate. Journal of Mathematical Physics, 10(7):1302–1304, July 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i7/p1302_s1.

Arunasalam:1969:TDT

[1295] V. Arunasalam. Transverse dielectric tensor for a free-electron gas in a uniform magnetic field. Journal of Mathematical Physics, 10(8):1305– 1313, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i8/p1305_s1.

Coon:1969:SCI

[1296] Darryl D. Coon. Singularities of conformal-invariant scattering amplitudes. Journal of Mathematical Physics, 10(8):1314–1318, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/
p1314_s1.

Cushing:1969:ESM

[1297] James T. Cushing. Exact static-model bootstrap solutions for arbitrary 2 × 2 crossing matrices. Journal of Mathematical Physics, 10(8):1319– 1326, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i8/p1319_s1.

Jernow:1969:AFL

[1298] Stanley Jernow and Emil Kazes. Asymptotic fields of the Lee model. Journal of Mathematical Physics, 10(8):1326–1339, August 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1326_
s1.

Kraut:1969:DEW

[1299] Edgar A. Kraut and Guy W. Lehman. Diffraction of electromagnetic waves by a right-angle dielectric wedge. Journal of Mathematical Physics, 10(8):1340–1348, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1340_s1.

Kennedy:1969:IAD

[1300] Frederick James Kennedy. Instantaneous action-at-a-distance formulation of classical electrodynamics. Journal of Mathematical Physics, 10(8):1349–1362, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1349_s1.

Coffey:1969:MCP

[1301] Timothy P. Coffey. Motion of a charged particle in a spatially periodic magnetic field. Journal of Mathematical Physics, 10(8):1362–1368, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1362_s1.

Taniuti:1969:PMN

[1302] Tosiya Taniuti and Nobuo Yajima. Perturbation method for a nonlinear wave modulation. I. Journal of Mathematical Physics, 10(8):1369– 1372, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i8/p1369_s1.

Kac:1969:CBS

[1303] Mark Kac and Colin J. Thompson. Critical behavior of several lattice models with long-range interaction. Journal of Mathematical Physics, 10(8):1373–1386, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1373_s1.

Korenman:1969:PER

[1304] Victor Korenman. On perturbation expansions for real-time Green’s functions. Journal of Mathematical Physics, 10(8):1387–1388, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1387_s1.

Majumdar:1969:NNNa

[1305] Chanchal K. Majumdar and Dipan K. Ghosh. On next-nearest-neighbor interaction in linear chain. I. Journal of Mathematical Physics, 10 (8):1388–1398, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1388_s1.

Majumdar:1969:NNNb

[1306] Chanchal K. Majumdar and Dipan K. Ghosh. On next-nearest-neighbor interaction in linear chain. II. Journal of Mathematical Physics, 10 (8):1399–1402, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1399_s1.

Lieb:1969:PTZ

[1307] Elliott H. Lieb and Colin J. Thompson. Phase transition in zero dimensions: A remark on the spherical model. Journal of Mathematical Physics, 10(8):1403–1406, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i8/p1403_s1.

Miller:1969:DDC

[1308] Marvin M. Miller. Dynamics in the diagonal coherent-state representation. Journal of Mathematical Physics, 10(8):1406–1411, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/
p1406_s1.

Roth:1969:PSA

[1309] J. Reece Roth. Periodic small-amplitude solutions to Volterra’s problem of two conflicting populations and their application to the plasma continuity equations. Journal of Mathematical Physics, 10(8):1412–1414, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1412_s1.

Simon:1969:GGS

[1310] Barry Simon. On the growth of the ground-state binding energy with increase in potential strength. Journal of Mathematical Physics, 10 (8):1415–1421, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1415_s1.

Peak:1969:SFH

[1311] D. Peak and A. Inomata. Summation over Feynman histories in polar coordinates. Journal of Mathematical Physics, 10(8):1422–1428, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1422_s1.

Sollfrey:1969:IFP

[1312] William Sollfrey. Inverse functions of the products of two Bessel functions. Journal of Mathematical Physics, 10(8):1429–1430, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/
p1429_s1.

Morrison:1969:GTM

[1313] J. C. Morrison. Group theory and mixed atomic configurations. Journal of Mathematical Physics, 10(8):1431–1437, August 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1431_s1.

Wong:1969:CC

[1314] Jack Wong. On a condition for completeness. Journal of Mathematical Physics, 10(8):1438–1440, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i8/p1438_s1.

Fink:1969:PSV

[1315] James P. Fink. Partial sums of vacuum polarization graphs. Journal of Mathematical Physics, 10(8):1440–1441, August 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1440_s1.

Jalickee:1969:FPS

[1316] John B. Jalickee, Arnold J. F. Siegert, and David J. Vezzetti. On fluids of particles with short-range repulsion and weak long-range attractive interaction. Journal of Mathematical Physics, 10(8):1442–1454, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1442_s1.

Krieger:1969:EQC

[1317] J. B. Krieger. Exact quantization conditions. II. Journal of Mathematical Physics, 10(8):1455–1458, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i8/p1455_s1.

Lewis:1969:EQT

[1318] H. R. Lewis, Jr. and W. B. Riesenfeld. An exact quantum theory of the time-dependent harmonic oscillator and of a charged particle in a time-dependent electromagnetic field. Journal of Mathematical Physics, 10(8):1458–1473, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1458_s1.

Barakat:1969:TPM

[1319] Richard Barakat and Elizabeth Baumann. Mth power of an N × N matrix and its connection with the generalized Lucas polynomials. Journal of Mathematical Physics, 10(8):1474–1476, August 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1474_s1.

Klink:1969:CGC

[1320] William H. Klink. Clebsch–Gordan coefficients and functions over cosets. Journal of Mathematical Physics, 10(8):1477–1490, August 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1477_
s1.

Kim:1969:SFE

[1321] Y. E. Kim. Solution of the Faddeev equation for local potentials by approximate product-integration. Journal of Mathematical Physics, 10 (8):1491–1494, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1491_s1.

Baus:1969:PIC

[1322] M. Baus. Possible instability of the cut-contributions to the ring equations. Journal of Mathematical Physics, 10(8):1494–1497, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/
p1494_s1.

Anderson:1969:BAI

[1323] N. Anderson, A. M. Arthurs, and P. D. Robinson. Bounds associated with integral equations of Kirkwood–Riseman type. Journal of Mathematical Physics, 10(8):1498–1500, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v10/i8/p1498_s1.

Bel:1969:SS

[1324] Luis Bel. Schwarzschild singularity. Journal of Mathematical Physics, 10(8):1501–1503, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1501_s1.

Belasco:1969:ICG

[1325] Elliot P. Belasco and Hans C. Ohanian. Initial conditions in general relativity: Lapse and shift formulation. Journal of Mathematical Physics, 10(8):1503–1507, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1503_s1.

Kobe:1969:SPC

[1326] Donald H. Kobe. Single-particle condensate and pair-correlation theory of inhomogeneous boson systems. Journal of Mathematical Physics, 10(8):1507–1517, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i8/p1507_s1.

Berzi:1969:RPL

[1327] Vittorio Berzi and Vittorio Gorini. Reciprocity principle and the Lorentz transformations. Journal of Mathematical Physics, 10(8):1518–1524, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1518_s1.

Otterson:1969:ETP

[1328] Paul Otterson. Equal-time pseudolimits of current commutators and Schwinger terms. Journal of Mathematical Physics, 10(8):1525–1530, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i8/p1525_s1.

Mills:1969:GTI

[1329] R. G. J. Mills and C. A. Hurst. Generalized triangular Ising lattice. Journal of Mathematical Physics, 10(8):1531–1540, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i8/p1531_s1.

Strecker:1969:ODP

[1330] Joseph L. Strecker. One-dimensional phase transition. Journal of Mathematical Physics, 10(8):1541–1554, August 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i8/p1541_s1.

Hayes:1969:QGH

[1331] C. F. Hayes. Quantization of the generalized Hamiltonian. Journal of Mathematical Physics, 10(9):1555–1558, September 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1555_s1.

Griffiths:1969:RRI

[1332] Robert B. Griffiths. Rigorous results for Ising ferromagnets of arbitrary spin. Journal of Mathematical Physics, 10(9):1559–1565, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i9/p1559_s1.

Exton:1969:CQE

[1333] Albert R. Exton, Ezra T. Newman, and Roger Penrose. Conserved quantities in the Einstein–Maxwell theory. Journal of Mathematical Physics, 10(9):1566–1570, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i9/p1566_s1.

Hecht:1969:WSS

[1334] K. T. Hecht and Sing Chin Pang. On the Wigner supermultiplet scheme. Journal of Mathematical Physics, 10(9):1571–1616, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1571_
s1.

Amit:1969:RSF

[1335] Daniel J. Amit. On the renormalization of the susceptibility of a Fermi liquid. Journal of Mathematical Physics, 10(9):1617–1620, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i9/p1617_s1.

delaPena-Auerbach:1969:NFS

[1336] L. de la Peña-Auerbach. New formulation of stochastic theory and quantum mechanics. Journal of Mathematical Physics, 10(9):1620–1630, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1620_s1.

Whitten:1969:PTB

[1337] R. C. Whitten. Potentials for three-body systems. Journal of Mathematical Physics, 10(9):1631–1634, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i9/p1631_s1.

Ciftan:1969:CSSb

[1338] Mikael Ciftan. On the combinatorial structure of state vectors in U(n). II. The generalization of hypergeometric functions on U(n) states. Journal of Mathematical Physics, 10(9):1635–1646, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1635_s1.

Baker:1969:CPS

[1339] George A. Baker, Jr. and Joseph Kahane. Critical-point singularities of the perturbation series for the ground state of a many-Fermion system. Journal of Mathematical Physics, 10(9):1647–1660, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1647_
s1.

Klauder:1969:PQC

[1340] John R. Klauder and L. Streit. Properties of “quadratic” canonical commutation relation representations. Journal of Mathematical Physics, 10 (9):1661–1669, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i9/p1661_s1.

Cook:1969:LPC

[1341] F. L. Cook and J. W. Jayne. Linear physical chains with Sturm–Liouville characteristic polynomials. Journal of Mathematical Physics, 10(9):1670– 1680, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1670_s1.

Hegerfeldt:1969:KIF

[1342] Gerhard C. Hegerfeldt. Kernel integral formulas for the canonical commutation relations of quantum fields. I. Representations with cyclic field. Journal of Mathematical Physics, 10(9):1681–1698, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1681_
s1.

Carmeli:1969:GAM

[1343] Moshe Carmeli. Group analysis of Maxwell’s equations. Journal of Mathematical Physics, 10(9):1699–1703, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i9/p1699_s1.

Santhanam:1969:GFC

[1344] T. S. Santhanam. Generating functions of classical groups and evaluation of partition functions. Journal of Mathematical Physics, 10(9):1704– 1710, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1704_s1.

Holman:1969:RT

[1345] Wayne J. Holman. Representation theory of SP(4) and SO(5). Journal of Mathematical Physics, 10(9):1710–1717, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1710_s1.

Pellicer:1969:NEG

[1346] R. Pellicer and R. J. Torrence. Nonlinear electrodynamics and general relativity. Journal of Mathematical Physics, 10(9):1718–1723, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i9/p1718_s1.

Ryan:1969:GFE

[1347] Michael P. Ryan, Jr. General form of the Einstein equations for a Bianchi type IX universe. Journal of Mathematical Physics, 10(9):1724–1728, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1724_s1.

Boyer:1969:CZS

[1348] Timothy H. Boyer. Concerning the zeros of some functions related to Bessel functions. Journal of Mathematical Physics, 10(9):1729–1744, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1729_s1.

Robinson:1969:CQN

[1349] David C. Robinson. Conserved quantities of Newman and Penrose. Journal of Mathematical Physics, 10(9):1745–1753, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1745_s1.

Bell:1969:ODB

[1350] G. M. Bell. One-dimensional bonded fluids. Journal of Mathematical Physics, 10(9):1753–1760, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i9/p1753_s1.

Rosina:1969:RDM

[1351] Mitja Rosina, Jerome K. Percus, Louis J. Kijewski, and Claude Garrod. Reduced density matrices of energy eigenstates. Journal of Mathematical Physics, 10(9):1761–1763, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i9/p1761_s1.

Ramanathan:1969:MDK

[1352] G. V. Ramanathan and G. Sandri. Model for the derivation of kinetic theory. Journal of Mathematical Physics, 10(9):1763–1773, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i9/p1763_s1.

Rowe:1969:PON

[1353] D. J. Rowe. Properties of overcomplete and nonorthogonal basis vectors. Journal of Mathematical Physics, 10(9):1774–1777, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1774_

s1. See erratum [1433].

Pajas:1969:DRS

[1354] P. Pajas. Degenerate representations of the sympletic groups. II. The noncompact group Sp(p, q). Journal of Mathematical Physics, 10(9): 1777–1788, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i9/p1777_s1.

Gabriel:1969:NMRa

[1355] John R. Gabriel. New methods for reduction of group representations. III. Journal of Mathematical Physics, 10(9):1789–1795, September 1969. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i9/p1789_
s1.

Harper:1969:DPA

[1356] Edward Y. Harper. Diffraction of plane acoustic waves and pulses as a singular perturbation problem. Journal of Mathematical Physics, 10 (9):1795–1803, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i9/p1795_s1.

Revzen:1969:SCA

[1357] M. Revzen and L. E. H. Trainor. Self-consistent approximations in many-body systems. Journal of Mathematical Physics, 10(9):1804–1808, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1804_s1.

Bassetto:1969:WHM

[1358] A. Bassetto and F. Paccanoni. Wiener–Hopf method and generalized integral transforms. Journal of Mathematical Physics, 10(9):1809–1812, September 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i9/p1809_s1.

Lee:1969:CEB

[1359] M. Howard Lee. Cluster expansion in a Bose system. Journal of Mathematical Physics, 10(10):1813–1818, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i10/p1813_s1.

Prosser:1969:FSI

[1360] Reese T. Prosser. Formal solutions of inverse scattering problems. Journal of Mathematical Physics, 10(10):1819–1822, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1819_s1.

Charap:1969:TVE

[1361] John M. Charap and B. M. Minton. Two-variable expansion of the scattering amplitude: An application of Appell’s generalized hypergeometric functions. Journal of Mathematical Physics, 10(10):1823–1833, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i10/p1823_s1.

Tokizawa:1969:CEN

[1362] Masamichi Tokizawa. On continuous eigenvalues in neutron thermalization. Journal of Mathematical Physics, 10(10):1834–1841, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i10/p1834_s1.

Debney:1969:SEE

[1363] G. C. Debney, R. P. Kerr, and A. Schild. Solutions of the Einstein and Einstein–Maxwell equations. Journal of Mathematical Physics, 10 (10):1842–1854, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i10/p1842_s1.

Garrod:1969:PHM

[1364] Claude Garrod and Mitja Rosina. Particle-hole matrix: Its connection with the symmetries and collective features of the ground state. Journal of Mathematical Physics, 10(10):1855–1861, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1855_s1.

Canosa:1969:DNM

[1365] José Canosa. Diffusion in nonlinear multiplicative media. Journal of Mathematical Physics, 10(10):1862–1868, October 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1862_s1.

deAzcarraga:1969:OMR

[1366] J. A. de Azcárraga and L. Oliver. Observables for massive relativistic particles of arbitrary spin. Journal of Mathematical Physics, 10(10):1869– 1874, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i10/p1869_s1.

Common:1969:PGP

[1367] A. K. Common. Properties of Generalizations to Padé approximants. Journal of Mathematical Physics, 10(10):1875–1880, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/
p1875_s1.

Percus:1969:OMT

[1368] Jerome K. Percus. One more technique for the dimer problem. Journal of Mathematical Physics, 10(10):1881–1884, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1881_s1.

Uzes:1969:PSG

[1369] C. A. Uzes. Plane-symmetric gauge field. Journal of Mathematical Physics, 10(10):1885–1888, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i10/p1885_s1.

Holman:1969:RTE

[1370] Wayne J. Holman III. Representation theory of SO(4, 1) and E(3, 1): an explicit spinor calculus. Journal of Mathematical Physics, 10(10):1888– 1897, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i10/p1888_s1.

Humi:1969:RLAa

[1371] Mayer Humi. Remarks on Lie-algebra-invariant equations. I. Journal of Mathematical Physics, 10(10):1897–1899, October 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1897_s1.

Humi:1969:RLAb

[1372] Mayer Humi. Remarks on Lie-algebra-invariant equations. II. Journal of Mathematical Physics, 10(10):1900–1902, October 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1900_s1.

Holland:1969:FTB

[1373] Douglas Francis Holland. Finite transformations and basis states of SU(n). Journal of Mathematical Physics, 10(10):1903–1905, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i10/p1903_s1.

Bailey:1969:DSS

[1374] Paul B. Bailey and George E. Barr. Diffraction by a slit or strip. Journal of Mathematical Physics, 10(10):1906–1913, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1906_s1.

Girardeau:1969:SMC

[1375] M. D. Girardeau. Soluble model of condensation. Journal of Mathematical Physics, 10(10):1914–1920, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i10/p1914_s1.

Kiang:1969:RTR

[1376] Hsiao S. Kiang. N-representability theorem for reduced density matrices. Journal of Mathematical Physics, 10(10):1920–1921, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/
p1920_s1.

Solimeno:1969:QHO

[1377] S. Solimeno, P. Di Porto, and B. Crosignani. Quantum harmonic oscillator with time-dependent frequency. Journal of Mathematical Physics, 10(10):1922–1928, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i10/p1922_s1.

Katz:1969:ADC

[1378] Amnon Katz. Alternative dynamics for classical relativistic particles. Journal of Mathematical Physics, 10(10):1929–1931, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/
p1929_s1.

Gabriel:1969:NMRb

[1379] J. R. Gabriel. New methods for reduction of group representations. IV. Journal of Mathematical Physics, 10(10):1932–1934, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i10/
p1932_s1.

Abhyankar:1969:RBE

[1380] K. D. Abhyankar and A. L. Fymat. Relations between the elements of the phase matrix for scattering. Journal of Mathematical Physics, 10(10):1935–1938, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i10/p1935_s1.

Bali:1969:CIS

[1381] Naren F. Bali, Darryl D. Coon, and Amnon Katz. Conformal-invariant scattering amplitude for spinless particles. Journal of Mathematical Physics, 10(10):1939–1952, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i10/p1939_s1.

Kermode:1969:ERF

[1382] Mark W. Kermode. Effective-range formula for two-body scattering coupled to a closed secondary channel. Journal of Mathematical Physics, 10(10):1953–1956, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i10/p1953_s1.

Tan:1969:UWH

[1383] H. S. Tan and D. K. Cheng. On the uncoupling of Wiener–Hopf equations for half-plane diffraction problems in a gyrotropic medium. Journal of Mathematical Physics, 10(10):1957–1963, October 1969. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v10/i10/p1957_s1.

Kschwendt:1969:NSI

[1384] H. Kschwendt. Numerical solution of integral equations using Legendre polynomials. Journal of Mathematical Physics, 10(10):1964–1968, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i10/p1964_s1.

Mickelsson:1969:RF

[1385] Jouko Mickelsson. Representation functions for SU(3). Journal of Mathematical Physics, 10(10):1968–1970, October 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i10/p1968_s1.

Leaf:1969:CTS

[1386] Boris Leaf. Canonical transformations and spectra of quantum operators. Journal of Mathematical Physics, 10(11):1971–1979, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i11/p1971_s1.

Leaf:1969:COS

[1387] Boris Leaf. Canonical operators for the simple harmonic oscillator. Journal of Mathematical Physics, 10(11):1980–1987, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i11/
p1980_s1.

Miller:1969:BEE

[1388] Melvin N. Miller. Bounds for effective electrical, thermal, and magnetic properties of heterogeneous materials. Journal of Mathematical Physics, 10(11):1988–2004, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i11/p1988_s1.

Miller:1969:BEB

[1389] Melvin N. Miller. Bounds for effective bulk modulus of heterogeneous materials. Journal of Mathematical Physics, 10(11):2005–2013, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v10/i11/p2005_s1. See erratum [1434].

Todorov:1969:SRC

[1390] I. T. Todorov and R. P. Zaikov. Spectral representation of the covariant two-point function and infinite-component fields with arbitrary mass spectrum. Journal of Mathematical Physics, 10(11):2014–2019, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i11/p2014_s1.

Asano:1969:PMN

[1391] Naruyoshi Asano, Tosiya Taniuti, and Nobuo Yajima. Perturbation method for a nonlinear wave modulation. II. Journal of Mathematical Physics, 10(11):2020–2024, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i11/p2020_s1.

Bracken:1969:CTD

[1392] A. J. Bracken and H. A. Cohen. On canonical SO(4, 1) transformations of the Dirac equation. Journal of Mathematical Physics, 10(11):2024– 2032, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i11/p2024_s1.

Sharp:1969:ILP

[1393] R. T. Sharp and C. S. Lam. Internal-labeling problem. Journal of Mathematical Physics, 10(11):2033–2038, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i11/p2033_s1.

Leplae:1969:BFS

[1394] L. Leplae and H. Umezawa. Boson formalism in superconductivity. Journal of Mathematical Physics, 10(11):2038–2046, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i11/
p2038_s1.

Finley:1969:SRI

[1395] J. D. Finley. Short-range interactions and analyticity in momentum transfer. Journal of Mathematical Physics, 10(11):2047–2059, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i11/p2047_s1.

Finley:1969:PCS

[1396] J. D. Finley. Possible characterization of short-range interactions. Journal of Mathematical Physics, 10(11):2060–2066, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i11/
p2060_s1.

Hammer:1969:PDL

[1397] C. L. Hammer and T. A. Weber. V -particle decay in the Lee model. Journal of Mathematical Physics, 10(11):2067–2077, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i11/
p2067_s1.

Mack:1969:ILR

[1398] G. Mack and I. Todorov. Irreducibility of the ladder representations of U(2, 2) when restricted to the Poincaré subgroup. Journal of Mathematical Physics, 10(11):2078–2085, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i11/p2078_s1.

Mukunda:1969:MFLa

[1399] N. Mukunda. Matrices of finite Lorentz transformations in a noncompact basis. I. Discrete series of O(2, 1). Journal of Mathematical Physics, 10 (11):2086–2092, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i11/p2086_s1.

Mukunda:1969:MFLb

[1400] N. Mukunda. Matrices of finite Lorentz transformations in a noncompact basis. II. Continuous series of O(2, 1). Journal of Mathematical Physics, 10(11):2092–2098, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i11/p2092_s1.

Chiang:1969:FNT

[1401] David Chiang. Further note on two binomial coefficient identities of Rosenbaum. Journal of Mathematical Physics, 10(11):2098, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v10/i11/p2098_s1. See comment [1435].

Ullah:1969:GDP

[1402] Nazakat Ullah. On a generalized distribution of the poles of the unitary collision matrix. Journal of Mathematical Physics, 10(11):2099–2103, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i11/p2099_s1.

Basu:1969:ADB

[1403] Debabrata Basu and S. N. Biswas. Accidental degeneracy in the Bethe– Salpeter equation. Journal of Mathematical Physics, 10(11):2104–2113, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i11/p2104_s1.

Deliyannis:1969:TO

[1404] Platon C. Deliyannis. Theory of observables. Journal of Mathematical Physics, 10(11):2114–2127, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i11/p2114_s1.

Frank:1969:ESC

[1405] William M. Frank. Erratum: “Strong-Coupling Limit in Potential Theory. II”. Journal of Mathematical Physics, 10(11):2128, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i11/

p2128_s1. See [1102].

Radhakrishnan:1969:IMS

[1406] D. Radhakrishnan. Internal multiplicity structure for the chain SU(n) ⊃ SU(n−1) ⊃ … ⊃ SU(2). Journal of Mathematical Physics, 10(12):2129– 2131, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2129_s1.

Bacry:1969:WHS

[1407] Henri Bacry and Arne Kihlberg. Wavefunctions on homogeneous spaces. Journal of Mathematical Physics, 10(12):2132–2141, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/
p2132_s1.

Choi:1969:TDS

[1408] Jong H. Choi. Treatment of degeneracies in the Schrödinger perturbation theory by partitioning technique. Journal of Mathematical Physics, 10 (12):2142–2148, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2142_s1.

Cunningham:1969:QFA

[1409] M. J. Cunningham and B. G. Wybourne. Quasiparticle formalism and atomic shell theory. Journal of Mathematical Physics, 10(12):2149–2155, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2149_s1.

Blaha:1969:CA

[1410] Stephen Blaha. Character analysis of U(N) and SU(N). Journal of Mathematical Physics, 10(12):2156–2168, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v10/i12/p2156_s1.

Zimmerman:1969:FIG

[1411] Robert L. Zimmerman. Functional integration and the generalized Matthews–Salam equations. Journal of Mathematical Physics, 10(12): 2169–2178, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2169_s1.

Rousseau:1969:APT

[1412] C. C. Rousseau and D. P. Saylor. Application of perturbation theory to many-body systems with localized particles. Journal of Mathematical Physics, 10(12):2179–2184, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i12/p2179_s1.

McNeil:1969:PRS

[1413] Richard P. McNeil and Richard A. Morrow. Physical regions of sixparticle processes. Journal of Mathematical Physics, 10(12):2185–2190, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2185_s1.

Calogero:1969:STB

[1414] F. Calogero. Solution of a three-body problem in one dimension. Journal of Mathematical Physics, 10(12):2191–2196, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/p2191_s1.

Calogero:1969:GSO

[1415] F. Calogero. Ground state of a one-dimensional N-body system. Journal of Mathematical Physics, 10(12):2197–2200, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/p2197_s1.

Shlidor:1969:SCA

[1416] O. Shlidor and M. Revzen. Self-consistent approximations in many-body systems. II. Journal of Mathematical Physics, 10(12):2201–2204, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i12/p2201_s1.

McQuistan:1969:VAO

[1417] R. B. McQuistan. Vacancy annihilation for one-dimensional dumbbell kinetics. Journal of Mathematical Physics, 10(12):2205–2207, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i12/p2205_s1.

Kvarda:1969:CSQ

[1418] Robert E. Kvarda. Canonical states in quantum-statistical mechanics. Journal of Mathematical Physics, 10(12):2208–2213, December 1969.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/
p2208_s1.

Glimm:1969:IRH

[1419] James Glimm and Arthur Jaffe. Infinite renormalization of the Hamiltonian is necessary. Journal of Mathematical Physics, 10(12):2213–2214, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2213_s1.

Katz:1969:RES

[1420] Amnon Katz. Relativistic effects of strong binding on slow particles. Journal of Mathematical Physics, 10(12):2215–2219, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/
p2215_s1.

Pahor:1969:HSM

[1421] S. Pahor and J. K. Shultis. Half-space multigroup transport theory. Journal of Mathematical Physics, 10(12):2220–2226, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/p2220_s1.

Alisauskas:1969:SGS

[1422] S. J. Alisauskas and A. P. Jucys. Substitution group and the stretched isoscalar factors for the group R_{5}. Journal of Mathematical Physics, 10 (12):2227–2233, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2227_s1.

Prosser:1969:SAC

[1423] Reese T. Prosser. Spectral analysis of classical central force motion. Journal of Mathematical Physics, 10(12):2233–2239, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v10/i12/p2233_s1.

Gelman:1969:FPI

[1424] Donald Gelman and Larry Spruch. Feynman path integrals and scattering theory. Journal of Mathematical Physics, 10(12):2240–2254, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v10/i12/p2240_s1.

Herbert:1969:KIH

[1425] D. C. Herbert. Kinematic interaction for the Heisenberg antiferromagnet at low temperature. Journal of Mathematical Physics, 10(12):2255– 2263, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2255_s1.

Sah:1969:SAP

[1426] Priyamvada Sah. Statistical average of product of phase sums arising in the study of disordered lattices. I. Journal of Mathematical Physics, 10 (12):2263–2266, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2263_s1.

Aslaksen:1969:CRT

[1427] Erik W. Aslaksen and John R. Klauder. Continuous representation theory using the affine group. Journal of Mathematical Physics, 10 (12):2267–2275, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2267_s1.

Antoine:1969:DFSb

[1428] J.-P. Antoine. Dirac formalism and symmetry problems in quantum mechanics. II. Symmetry problems. Journal of Mathematical Physics, 10 (12):2276–2290, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i12/p2276_s1.

Lewis:1969:UAT

[1429] Robert M. Lewis and Johannes Boersma. Uniform asymptotic theory of edge diffraction. Journal of Mathematical Physics, 10(12):2291–2305, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2291_s1.

Gutkowski:1969:APC

[1430] D. Gutkowski and A. Scalia. Analytic properties of a class of nonlocal interactions. II. Journal of Mathematical Physics, 10(12):2306–2318, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i12/p2306_s1.

Anonymous:1969:AIV

[1431] Anonymous. Author index to Volume 10. Journal of Mathematical Physics, 10(12):2319–2329, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Anonymous:1969:ASI

[1432] Anonymous. Analytic subject index to Volume 10. Journal of Mathematical Physics, 10(12):2330–2343, December 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rowe:1971:EPO

[1433] D. J. Rowe. Erratum: “Properties of Overcomplete and Nonorthogonal Basis Vectors”. Journal of Mathematical Physics, 12(4):742, April 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i4/p742_s1. See [1353].

Miller:1971:EBE

[1434] Melvin N. Miller. Erratum: “Bounds for Effective Bulk Modulus of Heterogeneous Materials”. Journal of Mathematical Physics, 12(6):1057, June 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v12/i6/p1057_s2. See [1389].

Gould:1971:CPC

[1435] H. W. Gould. Comment on a paper of Chiang. Journal of Mathematical Physics, 12(8):1743, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1401].

Calogero:1974:CPL

[1436] F. Calogero and C. Marchioro. A comment on the paper ‘Lower bounds on the energy eigenvalues of systems containing identical particles’. Journal of Mathematical Physics, 15(1):139–140, January 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1204].

Newton:1975:ADA

[1437] Roger G. Newton. Addendum: “Determination of the amplitude from the differential cross section by unitarity”, [J. Math. Phys. 9, 2050 (1968)]. Journal of Mathematical Physics, 16(6):1358, June 1975. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-REFERENCES 255

2427. URL http://jmp.aip.org/resource/1/jmapaq/v16/i6/p1358_

s1. See [1122].